### v1

# Weak Existence and Uniqueness for Super-Brownian Motion with Irregular Drift

Abstract
We establish weak existence and uniqueness for random field solutions of the one-dimensional SPDE
$d_{t}X_{t}=\frac{1}{2}\Delta X_{t}+h(X_{t})+\sqrt{X_{t}}\dot{W},\quad t\geq 0,$
where $\dot{W}$ is space–time white noise and $h$ is a bounded drift with $h(0)\geq 0$. The proof relies on an extension of the duality relation of the super-Brownian motion, which allows us to treat a broad class of admissible drifts, including functions that are non-Lipschitz or discontinuous at zero.
1 Introduction
We consider one-dimensional SPDEs of the form
$d_{t}X_{t}=\frac{1}{2}\Delta X_{t}+h(X_{t})+\sqrt{X_{t}}\dot{W},\quad X_{0}=f,$ (1.1)
where $f$ is a non-negative, continuous, bounded function, $\dot{W}$ is space-time white noise and
$h(x)=h_{1}(x)+h_{\infty}(x),\quad x\in\mathbb{R},$
with
$h(0)\geq 0.$ (1.2)
Hereby,
$h_{1}(x)=\int_{0}^{\infty}(\operatorname{e}^{-\lambda x})(\nu^{2}-\nu^{1})(d\lambda)\quad\forall x\in\mathbb{R},$ (1.3)
and
$h_{\infty}(x)=b_{0}\mathbbm{1}_{x=0}+b_{1}\mathbbm{1}_{x>0},\quad x\in\mathbb{R},$ (1.4)
for $b_{0}\geq\langle\nu_{1}-\nu_{2},1\rangle$, $b_{1}\in\mathbb{R}$ and finite measures $\nu_{1},\nu_{2}$ on $(0,\infty)$.
In this article, we establish weak well-posedness for solutions of (1.1) using duality methods. Hereby, we consider solutions in the space $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$, which is the space of continuous functions taking values in
$\mathcal{C}_{tem}^{+}=\{f\in\mathcal{C}(\mathbb{R})\colon f\geq 0,|f|_{(\lambda)}<\infty\quad\forall\lambda<0\},$
where
$|f|_{(\lambda)}=\sup_{x\in\mathbb{R}}|\operatorname{e}^{\lambda|x|}f(x)|,\quad f\in\mathcal{C}(\mathbb{R}).$
We also allow random initial values that are distributed according to a measure in $\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ (see Section 2.1). That means, we prove the following theorem.
Theorem 1.1.
Let $\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ with $p>2$ and assume $X_{0}\sim\eta$ is such that
$\sup_{x\in\mathbb{R}}\mathbb{E}[X_{0}(x)]<\infty.$ (1.5)
Then there exists a unique weak solution to (1.1) in $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$ in the sense of Definition 4.1.
The main tool for the proof of Theorem 1.1 is the derivation of a duality relation (see Proposition 4.3). When $h(0)=0$, this relation also allows us to prove properties of the cozero set of the solution, even when $h$ has a discontinuity at zero.
Our assumptions on the drift function $h$ permit significant irregularities, including Hölder continuity with exponent less than $\frac{1}{2}$ and even discontinuities at zero. Several examples are provided in Section 5. Irregularities at zero are of particular interest, as this is precisely where the noise coefficient becomes non-Lipschitz. As a result, standard techniques such as the Dawson–Girsanov theorem can no longer be directly applied to establish weak uniqueness when the drift is less than $\frac{1}{2}$-Hölder continuous or exhibits discontinuities at zero. To the best of our knowledge, when the drift is discontinuous, neither weak uniqueness nor weak existence is known It turns out, that we can establish weak well-posedness of (1.1) for drift terms $h$ that lead to non-unique solutions or non-existence of solutions in the case of the one-dimensional SDE
$dx_{t}=h(x_{t})dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=y,\quad t\geq 0,$
where $y\in\mathbb{R}$ and $(B_{t})_{t\geq 0}$ is a standard Brownian motion.
A similar behavior was also observed in [3] in the case of the Wright-Fisher SPDE.
Equation (1.1) with $h=0$ is the density of the super-Brownian motion in one dimension, which was extensively studied in the literature (see for instance [26, 15, 8, 10, 18, 23]). When the process is subject to some nonzero space dependent immigration, that is in the case where $h$ is continuous and does not depend on $X$ but instead on the spatial component $x$, pathwise non-uniqueness was established in [7].
When $h$ is continuous, (1.1) also fits in the setting of [25], where the existence of solutions is proven, but no uniqueness result was obtained for non-Lipschitz coefficients.
Results on weak well-posedness for SPDEs with non-Lipschitz coefficients remain scarce. In the setting of Super-Brownian Motion and other measure valued branching processes a key method for establishing weak uniqueness is the duality approach. We will also employ this approach in our analysis. Since this method relies on guessing the correct duality functions it is sensitive to changes in the coefficients of the SPDE and thus one usually needs to assume a certain structure of the coefficients. The duality method was for instance used in [1, 21]. Hereby, the authors of [1] are using moment duality and in [21] the Laplace transform is used. In this article we will also use the Laplace transform as a starting point for our duality relation. Our approach is essentially an extension of the well known duality relation between the Super Brownian Motion and the log Laplace equation. More precisely, if $X$ denotes the solution to (1.1) with $h=0$ and $V(\mu)$ is the solution to
$\begin{cases}&d_{t}V_{t}(\mu)=\frac{1}{2}\Delta V_{t}(\mu)-\frac{1}{2}V_{t}(\mu)^{2},\quad(t,x)\in(0,\infty)\times\mathbb{R}\\ &V_{t}(\mu)\Rightarrow\mu,\text{ as }t\rightarrow 0,\end{cases}$ (1.6)
for a finite measure $\mu$, then it is well-known (see for instance [8]) that
$\mathbb{E}[\exp(-\langle X_{t},\mu\rangle)]=\mathbb{E}[\exp(-\langle X_{0},V_{t}(\mu)\rangle)],\quad\forall t\geq 0,$
for all finite measures $\mu$. Furthermore, if $h\equiv a>0$, we get the following duality relation
$\mathbb{E}[\exp(-\langle X_{t},\mu\rangle)]=\mathbb{E}\Big{[}\exp\Big{(}-\langle X_{0},V_{t}(\mu)\rangle-a\int_{0}^{t}\langle V_{s}(\mu),1\rangle ds\Big{)}\Big{]},\quad\forall t\geq 0,$ (1.7)
which was for example demonstrated in [19]. In this article, inspired by [21, 22] as well as [1], we are able to cover more complex immigration mechanisms $h$ by introducing a space-time jump process to (1.6). We outline this procedure in more detail in Section 4.
More recently, the authors of [3] extended the results of [1] for the one-dimensional Wright–Fisher SPDE by allowing certain irregular drift terms that were previously not covered. A key element in their analysis was the phenomenon of coming down from infinity in the dual branching–coalescing particle system, which is discussed in detail in [2]. Similarly, in this article, we employ the theory of initial trace solutions to the log-Laplace equation (see [6, 17, 20]) to extend the class of admissible drift functions $h$, allowing in particular for certain discontinuities at zero. The framework of initial trace solutions enables us to introduce jumps of infinite height in our dual process. In Section 4, we give some intuition on how this allows us to obtain a duality relation when $h$ has certain discontinuities at zero. We note that the theory of initial trace solutions is closely related to the phenomenon of coming down from infinity used in [3], both conceptually and technically.
In comparison to [3], we are able to control exponential moments of the number of jumps of the dual process in sufficiently small time intervals. This allows us to extend the class of admissible drift terms $h$ to include functions with $b_{0}>0$ and $b_{1}\leq 0$.
For this purpose, we introduce an artificial branching process that allows us to keep track of the number of jumps of the dual process. Again, this approach was inspired by [3].
Organization of the Article
First, in Section 2, we introduce several function spaces and notation that is used throughout this article. In Section 3, we give a brief summary of results on certain non-linear parabolic equations, in particular the theory of initial traces and very singular solutions. Next, In Section 4, we outline the proof of Theorem 1.1. We then proceed to discuss several examples of admissible drift terms. In Section 6, we prove the duality relation (4.14). To this end, we introduce a branching process that allows us to control the moments of the number of jumps of the dual process in Section 6.1. We furthermore approximate the dual process by truncating the size of the jumps and prove the convergence of this approximation in Section 6.2. In Sections 7 and 8, we prove the weak existence and uniqueness of solutions by using the duality relation from Section 6. Finally, Section 9 is dedicated to the proof of Theorem 4.5, where we use the duality relation to prove properties of the cozero set of solutions to (1.1) under certain additional conditions.
Convention for Constants
In this article, constants are denoted by $C$ and $c$, with their precise values considered irrelevant. We reserve the freedom to adjust these values without altering the notation for readability. When necessary, dependence on additional parameters is expressed as $C=C(\lambda)$.
2 Preliminaries
Before we state our main results in the next section, we introduce some notation.
2.1 Notation
We use the following notation for common function spaces.
  1. 2.1.
    $\mathcal{C}(\mathbb{R})$ denotes the space of real-valued, continuous functions on $\mathbb{R}$ and $\mathcal{C}^{k}(\mathbb{R})$, $k\in N$ denotes the space of $k$-times continuously differentiable, real-valued functions on $\mathbb{R}$.
  2. $2.$
    $\mathcal{C}_{c}(\mathbb{R})$ denotes the space of compactly supported functions in $\mathcal{C}(\mathbb{R})$.
  3. 2.3.
    $\mathcal{C}_{tem}=\{f\in\mathcal{C}(\mathbb{R})\colon|f|_{(\lambda)}<\infty\quad\forall\lambda<0\}$, where $|f|_{(\lambda)}=\sup_{x\in\mathbb{R}}|\operatorname{e}^{\lambda|x|}f(x)|$ for $f\in\mathcal{C}(\mathbb{R})$.
  4. 2.4.
    $\mathcal{P}(\mathcal{C}_{tem}^{+})$ denotes the set of all probability measures on $\mathcal{C}_{tem}$
  5. 2.5.
    $\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})=\{\mu\in\mathcal{P}(\mathcal{C}_{tem}^{+})\colon\sup_{x\in\mathbb{R}}\operatorname{e}^{\lambda|x|}\int_{\mathcal{C}_{tem}^{+}}|f(x)|^{p}\mu(df)<\infty,\quad\forall\lambda<0\}.$
  6. 2.6.
    $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$ is the space of functions $f\colon\mathbb{R}_{+}\rightarrow\mathcal{C}_{tem}^{+}$, which are continuous with respect to the topology induced by $\mathcal{C}_{tem}$.
  7. 2.7.
    $L^{p}(A,m)$ for $p\in[1,\infty]$ denotes the usual $L^{p}$-space on the measure space $(A,\mathcal{A},m)$. If there is no ambiguity we write $L^{p}$ instead of $L^{p}(A,m)$ for $p\in[1,\infty]$.
  8. 2.8.
    For $p\in[1,\infty)$ we denote by $L^{1}_{loc}((0,\infty)\times\mathbb{R})$ the set of measurable functions $f\colon[0,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ such that $\int_{0}^{T}\int_{\mathbb{R}}|f(t,x)|^{p}dxdt<\infty$ for all $T>0$.
  9. 2.9.
    $\mathcal{M}_{F}^{+}$ denotes the set of finite, non-negative measures on $\mathbb{R}$.
  10. 2.10.
    $\mathcal{M}_{reg}^{+}$ denotes the set of non-negative outer regular measures on $\mathbb{R}$.
  11. 2.11.
    $\mathcal{B}(\mathbb{R})$ denotes the Borel $\sigma$-algebra.
If $A$ is a sub-space $\mathcal{C}(\mathbb{R})$, we denote by $A^{+}$ the subset of non-negative functions in $A$. Furthermore, we denote by $A_{c}$ the subspace of compactly supported functions in $A$. We also introduce the following notation
$\langle\mu,f\rangle=\int_{\mathbb{R}}f(x)\mu(dx),$
for a measure $\mu\in\mathcal{M}_{F}^{+}$ and a function $f\in\mathcal{C}(\mathbb{R})$. Furthermore, if $g\in L^{1}(\mathbb{R})$ we also write
$\langle f,g\rangle=\int_{\mathbb{R}}f(x)g(x)dx.$
Furthermore, for a measurable subset $D\subset\mathbb{R}$ we denote by $\mathrm{Leb}(D)$ the Lebesgue measure of $D$.
Finally, if $I$ is a set or a tuple, |I|=#I|I|=\#I denotes the number of elements in the set or the length of the tuple, respectively.
3 Non-linear Parabolic Equations and Initial Trace
In this section we discuss the nonlinear parabolic equation
$d_{t}V_{t}=\frac{1}{2}\Delta V_{t}-\frac{1}{2}V_{t}^{2},\quad t\geq 0$ (3.1)
on $\mathbb{R}$. This partial differential equation plays a central role in our proofs due to its close relation with the dual process. It is possible to derive well posedness results for (3.1) for highly irregular initial data. For a more through discussion we refer the reader to [20] and [17] . Here we shall only give a brief overview. More precisely, we discuss some basic properties of solutions to
$\begin{cases}&d_{t}V_{t}(\mu)=\frac{1}{2}\Delta V_{t}(\mu)-\frac{1}{2}V_{t}(\mu)^{2},\quad(t,x)\in(0,\infty)\times\mathbb{R}\\ &V_{t}(\mu)\Rightarrow\mu,\text{ as }t\rightarrow 0,\end{cases}$ (3.2)
where $\mu$ is a finite measure as well as solutions to the initial trace problem
$\begin{cases}&d_{t}\mathcal{W}_{t}(A,\nu)=\frac{1}{2}\Delta\mathcal{W}_{t}(A,\nu)-\frac{1}{2}\mathcal{W}_{t}(A,\nu)^{2},\quad(t,x)\in(0,\infty)\times\mathbb{R}\\ &\{y\in\mathbb{R}\colon\lim_{t\rightarrow 0}\int_{y-r}^{y+r}\mathcal{W}_{t}(A,\nu)(x)dx=\infty,\forall r>0\}=A\\ &\lim_{t\rightarrow 0}\int\phi\mathcal{W}_{t}(A,\nu)(x)dx=\int\phi\nu(dx),\forall\phi\in\mathcal{C}_{c}(A^{c}),\end{cases}$ (3.3)
where $A\subset\mathbb{R}$ is a closed set and $\nu$ is a non-negative Radon measure on $A^{c}$. If $A=\{x\}$ for some $x\in\mathbb{R}$ we sometimes write
$\mathcal{W}_{t}(x,\nu)=V_{t}(\nu+\infty\delta_{x}),\quad t>0,x\in\mathbb{R}.$ (3.4)
Furthermore, for $t<0$ we use the convention that $\mathcal{W}_{t}(A,\nu)=V_{t}(\mu)=0$ and $V_{0}(\mu)=\mu$.
3.1 Finite Measures
In [5] and [12, 22] the authors prove existence and uniqueness of solutions to (3.2) as well as several estimates. We summarized their results in this section. First, we get get existence and uniqueness by [22, Lemma 2.1], which we restated in the following proposition
Proposition 3.1.
Let $\mu$ be a finite measure. Then there exists a unique solution $V(\mu)$ (3.2) in $L^{1}_{loc}((0,\infty)\times\mathbb{R})$ satisfying
$V_{t}(\mu)(x)=(S_{t}\mu)(x)-\int_{0}^{t}\frac{1}{2}\Big{(}S_{t-s}V_{s}(\mu)^{2}\Big{)}(x)ds,\quad t>0,x\in\mathbb{R},$ (3.5)
where $(S_{t})_{t\geq 0}$ is the heat semi-group with transition density $(p_{t})_{t\geq 0}$ given by (4.1).
The following lemma is a well known fact about (3.2) (see for instance [22, Lemma 2.1]).
Lemma 3.2.
Let $\mu$ be a finite measure. Then it holds that
$V_{t}(\mu)(x)\leq S_{t}\mu(x),$
for all $t\geq 0$ and $x\in\mathbb{R}$.
The following lemma is [22, Lemma 2.6] and provides estimates that we shall use frequently.
Lemma 3.3.
Let $\mu$ and $\eta$ be two finite measures. Then the following holds:
  1. (i)
    For all $t\geq 0$ it holds that
    $V_{t}(\mu+\eta)(x)\leq V_{t}(\mu)(x)+V_{t}(\eta)(x),\quad x\in\mathbb{R}.$
  2. (ii)
    If $\mu\leq\eta$, then
    $V_{t}(\mu)(x)\leq V_{t}(\eta)(x),\quad x\in\mathbb{R},$
    for all $t\geq 0$.
Finally, using [14, Lemma 2.3] we can infer the following lemma.
Lemma 3.4.
Let $\mu$ be a finite, non-zero measure. Then it follows that
$\int_{0}^{\infty}\langle V_{t}(\mu),1\rangle dt=\infty.$
3.2 Initial Trace
In this section we shall give a brief overview of results about solutions to (3.3) based on [20], [17] and [3]. We denote by $\mathcal{T}$ the set of pairs $(A,\nu)$, where $A$ is a closed subset of $\mathbb{R}$ and $\nu$ is a Radon measure supported on $A^{c}$. We introduce the partial order $\preceq$ on $\mathcal{T}$, where
$(A,\nu)\preceq(\hat{A},\hat{\nu}),\text{ if }A\subset\hat{A}\text{ and }\nu\leq\hat{\nu}$.
We furthermore denote by $\mathcal{M}_{reg}^{+}$ the set of outer regular Borel measures on $\mathbb{R}$ and define the map
$\eta\colon\mathcal{T}\rightarrow\mathcal{M}_{reg}^{+},(A,\nu)\mapsto\eta^{(A,\nu)},$
where
$\eta^{A,\nu}(B)=\begin{cases}\infty,B\cap A\neq\emptyset,\\ \nu(B),B\cap A=\emptyset,\end{cases}$
for all Borel-measurable sets $B\subset\mathbb{R}$. Since $\nu\in\mathcal{M}_{reg}^{+}$ is not necessarily locally bounded, we introduce the set of singular sets $\mathcal{S}_{\eta}$ associated with $\eta$ given by
$\mathcal{S}_{\eta}=\{x\in\mathbb{R}\colon\eta(U)=\infty\text{ for every open neighborhood }U\text{ of }x\},$
as well as the set of regular points $R_{\eta}$ given by $R_{\eta}=\mathcal{S}_{\eta}^{c}$. In order to formulae the results of this section, we furthermore need the notion of $m$-weak convergence.
Definition 3.5.
A sequence $(\eta_{n})_{n\geq 1}\subset\mathcal{M}_{reg}^{+}$ is said to converge $m$-weakly to $\eta\in\mathcal{M}_{reg}^{+}$ if the following is satisfied:
  1. (i)
    For all open sets $U\subset\mathbb{R}$ such that $\eta(U)=\infty$ it follows that $\lim_{n\rightarrow\infty}\eta_{n}(U)=\infty$.
  2. (ii)
    For all compact sets $K\subset\mathcal{R}_{\eta}$ there exists $M>0$ and $N\in\mathbb{N}$ such that $\eta_{n}(K)<M$ for all $n\geq N$.
  3. (iii)
    For all $\phi\in\mathcal{C}_{c}(\mathcal{R}_{\eta})$ it holds that
    $\lim_{n\rightarrow\infty}\int\phi d\eta_{n}=\int\phi d\eta.$
In the next sections we shall rely heavily on the following result, which is taken from [20, Proposition 3.10] [17, Theorem 4] and [3, Section 2.1].
Proposition 3.6.
Let $(A,\nu)\in\mathcal{T}$. Then there exists a unique solution to (3.3). Furthermore, the following holds:
  1. 3.1.
[…]
    $\mathcal{W}_{t}(A,\nu)(x)\leq\mathcal{W}_{t}(\hat{A},\hat{\eta})(x),\quad\forall t>0,x\in\mathbb{R}.$
  2. 3.2.
    Let $((A_{n},\nu_{n}))_{n\geq 1}\subset\mathcal{T}$ such that $\eta^{A_{n},\nu_{n}}$ converges $m$-weakly to $\eta^{(A,\nu)}$, then it follows that $\mathcal{W}(A_{n},\nu_{n})$ converges to $\mathcal{W}(A,\nu)$ uniformly on compacts.
In particular, by Lemma 3.3 and Proposition 3.6 it follows that $V_{t}(n\delta_{x})y$ converges monotonically to $\mathcal{W}_{t}(x,0)(y)$ for all $t>0$ and $x,y\in\mathbb{R}$. Therefore,
$V_{t}(n\delta_{x})(y)\leq\mathcal{W}_{t}(x,0)(y),\forall t>0,x,y\in\mathbb{R},\quad n\geq 1$ (3.6)
which we shall frequently use in subsequent sections. The following proposition is a version of Theorem 1 from [6].
Proposition 3.7.
It holds that
$\mathcal{W}_{t}(\{x_{0}\},0)(x)=t^{-1}f(t^{-1/2}|x-x_{0}|),$
where $f\colon[0,\infty)\rightarrow[0,\infty)$ is a smooth function satisfying
$f(x)\asymp C\operatorname{e}^{-\frac{x^{2}}{2}}x\Big{(}1+o(x^{-2})\Big{)},$
as $x\rightarrow\infty$.
4 Proof Outline
In this section, we outline the proof of Theorem 1.1 and, in particular, motivate the duality relation, which is the main tool in our analysis. Before we outline our arguments in more detail, we define what we understand under a weak solution to (1.1).
Definition 4.1 (Weak Solution).
We say that (1.1) has a weak solution $(X_{t})_{t\geq 0}$, if there exists a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ such that $(X_{t})_{t\geq 0}$ is an adapted $\mathcal{C}_{tem}^{+}$ - valued continuous process with $X_{0}=\mu$ a.s. and such that $\dot{W}$ is a space-time white noise and for every $(t,x)\in(0,\infty)\times\mathbb{R}$ almost surely
$\begin{split}&X_{t}(x)=\int_{\mathbb{R}}p_{t}(x-y)X_{0}(y)dy+\int_{\mathbb{R}}\int_{0}^{t}p_{t-s}(x-y)h(X_{s}(y))dsdy\\ &\hskip 56.9055pt+\int_{\mathbb{R}}\int_{0}^{t}p_{t-s}(x-y)\sqrt{X_{s}(y)}\dot{W}(dsdy).\end{split}$
Hereby,
$p_{t}(x):=\frac{1}{\sqrt{2\pi t}}\operatorname{e}^{-x^{2}/(2t)}.$ (4.1)
Both, the proof of weak uniqueness and existence in Theorem 1.1 are based on a duality approach.
We first outline this approach for a simplified setting and then step by step add additional modifications to cover more general drift terms. Let us first consider the case where
$h_{\infty}\equiv\langle\nu_{1},1\rangle\text{ and }\nu_{2}\equiv 0,$ (4.2)
which means that
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda).$
If we apply Itô’s formula to $\exp(-\langle\phi,X_{t}\rangle)$, for $\phi\in\mathcal{C}^{2}(\mathbb{R})$ with $\phi\geq 0$. We obtain that
$\exp(-\langle\phi,X_{t}\rangle)-\int_{0}^{t}\exp(-\langle\phi,X_{s}\rangle)\Big{(}-\langle\frac{1}{2}\Delta\phi,X_{s}\rangle+\frac{1}{2}\langle\phi^{2},X_{s}\rangle-\langle\phi,h(X_{s})\rangle\Big{)}ds$ (4.3)
is an $\mathcal{F}_{t}^{X}$-martingale. Suppose we can show the existence of a measure-valued process $Y$ such that
$\exp(-\langle Y_{t},\psi\rangle)-\int_{0}^{t}\exp(-\langle Y_{s},\psi\rangle)\Big{(}-\langle\frac{1}{2}\Delta Y_{s},\psi\rangle+\frac{1}{2}\langle Y_{s}^{2},\psi\rangle-\langle Y_{s},h(\psi)\rangle\Big{)}ds,$ (4.4)
is an $\mathcal{F}^{Y}_{t}$-martingale for $\psi\in\mathcal{C}_{tem}^{+}$. Then, if $X,Y$ are independent, under certain technical conditions it follows that
$\mathbb{E}[\exp(-\langle X_{t},Y_{0}\rangle)]=\mathbb{E}[\exp(-\langle X_{0},Y_{t}\rangle)],\quad\forall t\geq 0,$
by [11, Theorem 4.4.11]. This would allow us to infer that the one-dimensional laws are unique and thus by standard arguments the finite dimensional laws of $X$ are unique as well. Consider the SPDE
$d_{t}Y_{t}=\frac{1}{2}\Delta Y_{t}-\frac{1}{2}Y_{t}^{2}+\int_{0}^{\infty}\int_{0}^{\infty}\int_{0}^{\infty}\lambda\mathbbm{1}_{r\leq Y_{t-}(x)}\mathcal{N}(dr,dx,d\lambda,dt),$ (4.5)
where $\mathcal{N}$ is a Poisson random measure with intensity
$drdx\nu^{1}(d\lambda)dt.$
A formal application of Ito’s formula shows that a solution to (4.5) solves (4.4) and thus the solution to (4.5) is our candidate for the dual process.
Now, let us consider the case where the $\nu^{1},\nu^{2}\neq 0$ and $h_{\infty}\equiv\langle\nu^{1}+\nu^{2},1\rangle$. That is, we now consider the case where
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})d\nu^{1}(d\lambda)+\int_{0}^{\infty}(1+\operatorname{e}^{-\lambda x})d\nu^{2}(d\lambda)$ (4.6)
We will see that the duality function changes in this setting. In order to define the modified duality function needed for this setting we also introduce an independent sequence of i.i.d. random variables $(M_{k})_{k\geq 1}$ taking values in $\{1,2\}$ such that
$\mathbb{P}(M_{k}=i)=\frac{\nu^{i}((0,\infty))}{\nu^{1}((0,\infty))+\nu^{2}((0,\infty))},\quad i\in\{1,2\},$ (4.7)
We furthermore define
$J_{t}=\#\{k\in\mathbb{N}\colon\bar{T}_{k}\leq t,M_{k}=2\},$ (4.8)
where $\bar{T}_{k}$ denotes the $k$-th jump of the process $(\langle Y_{t},1\rangle)_{t\geq 0}$. Hereby, $Y$ is constructed as before, but with $\nu^{1}$ replaced by $\nu=\nu^{1}+\nu^{2}$. Then, we will see in Section 6 that the duality relation is given by
$\mathbb{E}[\exp(-\langle X_{t},Y_{0}\rangle)]=\mathbb{E}[(-1)^{J_{t}}\exp(-\langle X_{0},Y_{t}\rangle)],\quad\forall t\geq 0.$ (4.9)
Now, we turn to the discontinuous part $h_{\infty}$. If we formally consider $\nu^{1}=d_{1}\delta_{\infty}$ and $\nu^{2}=d_{2}\delta_{\infty}$ with $d_{1},d_{2}\geq 0$ and plug it into
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+\int_{0}^{\infty}(1+\operatorname{e}^{-\lambda x})\nu^{2}(d\lambda)$ (4.10)
we arrive at the function
$h_{\infty}(x)=d_{1}\mathbbm{1}_{x>0}+2d_{2}\mathbbm{1}_{x=0}+d_{2}\mathbbm{1}_{x>0}=\lim_{n\rightarrow\infty}d_{1}(1-\operatorname{e}^{-nx})+\lim_{n\rightarrow\infty}d_{2}(1+\operatorname{e}^{-nx}).$
Thus, for each fixed $n$, if we set
$\nu_{n}^{1}=\nu^{1}_{|(0,n]}+d_{1}\delta_{n},\quad\nu_{n}^{2}=\nu^{2}_{|(0,n]}+d_{2}\delta_{n}$
we are again in the setting of (4.6) and (4.10) with $\nu^{i}$ replaced by $\nu^{i}_{n}$, $i=1,2$. Hereby, $\nu^{i}_{|(0,n])}$, $i=1,2$ denotes the restriction of $\nu^{i}$ to the interval $(0,n]$. Inspired by (4.9), this will allow us to establish approximate duality. That is, if
$\mathbb{E}[\exp(-\langle X_{t},Y_{0}\rangle)]=\lim_{n\rightarrow\infty}\mathbb{E}[(-1)^{J_{t}^{n}}\exp(-\langle X_{0},Y_{t}^{n}\rangle)],\quad\forall t\geq 0,$ (4.11)
for each $n\geq 1$, $Y^{n}$ solves (4.5) with $\nu^{1}$ replaced by $\nu^{1}_{n}+\nu^{2}_{n}$ and $J^{n}$ is defined accordingly as in (4.8). Then, we will be able to construct processes $Y,J$, that are limits of $Y^{n}$ and $J^{n}$, respectively, such that (4.11) holds without limit. Loosely speaking, such a $Y$ is a solution to (4.5) with $\nu^{1}$ formally replaced by $\nu^{1}+\nu^{2}+d_{1}\delta_{\infty}+d_{2}\delta_{\infty}$ and $J$ is constructed accordingly (see Section 6).
With the duality function as described above, we cannot cover the full range of parameters $b_{0}\geq 0$ and $b_{1}\in\mathbb{R}$ that are admissible for $h_{\infty}$. In fact, with (4.9) we are only able to prove weak well-posedness for drifts of the form
$\begin{split}h_{1}(x)&=\int_{0}^{\infty}\operatorname{e}^{-\lambda x}(\nu^{2}-\nu^{1})(d\lambda),\\ h_{\infty}(x)&=2d_{2}\mathbbm{1}_{x=0}+(d_{1}+d_{2})\mathbbm{1}_{x>0}+\langle\nu^{1}+\nu^{2},1\rangle.\end{split}$
Since $d_{1},d_{2}\geq 0$ this does in particular not cover the regime $b_{0}\geq 0$ and $b_{1}\leq 0$. Introducing an additional parameter $a\in\mathbb{R}$ and setting
$h_{\infty}(x)=2d_{2}\mathbbm{1}_{x=0}+(d_{1}+d_{2})\mathbbm{1}_{x>0}+\langle\nu^{1}+\nu^{2},1\rangle+a$ (4.12)
allows us to cover the complete regime of parameters $b_{0}\geq\langle\nu^{1}-\nu^{2},1\rangle$ and $b_{1}\in\mathbb{R}$ for $h_{\infty}$ given by (1.4). More precisely, we set
[…]
It is easily verified, that in this case $d_{1},d_{2}\geq 0$ and by plugging these definitions into (4.12) we recover (1.4).
Remark 4.2.
There is not necessarily a unique choice for $\nu^{1},\nu^{2}$ and $d_{1}$, $d_{2}$ and $a$ that represent a given function $h$.
In order to handle this class of drift function we need to modify the duality function again. In this case we will see in Section 6 that the approximate duality relation is given by
$\begin{split}&\mathbb{E}\Big{[}\exp\Big{(}-\langle X_{t},Y_{0}\rangle\Big{)}\Big{]}\\ &\quad=\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{n}}\exp\Big{(}-\langle X_{0},Y_{t}^{n}\rangle\Big{)}\exp\Big{(}-a\int_{0}^{t}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]},\quad\forall t\geq 0,\end{split}$ (4.13)
where $Y^{n}$ and $J^{n}$ are defined as in (4.11) with $d_{1},d_{2}$ as in (A). The relation (4.13) can again be easily motivated by a formal application of Ito’s formula. Hereby, we want to point out that $a$ may be negative. This means, one of our main tasks is to show that the expectation on right hand side of (4.13) is well defined. As before, we are able to construct the limits of $Y^{n}$ and $J^{n}$ and thus do away with the limit in (4.13). These considerations lead us to the following proposition, which we prove in Section 6
Proposition 4.3.
Assume that $(X_{t})_{t\geq 0}$ is a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{rap}^{+})$, $p\geq 2$ and such that (1.5) holds. Let $Y_{0}=\mu$ for $\mu\in\mathcal{M}_{F}^{+}$. Then, there exists $T_{0}>0$ such that there exist processes $(J_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$, taking values in $\mathbb{N}$ and $\mathcal{M}_{reg}^{+}$, respectively, that are independent of $X$ and such that
$\mathbb{E}\Big{[}\exp\Big{(}-\langle X_{t},Y_{0}\rangle\Big{)}\Big{]}=\mathbb{E}\Big{[}(-1)^{J_{t}}\exp\Big{(}-\langle X_{0},Y_{t}\rangle\Big{)}\exp\Big{(}-a\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{)}\Big{]},$ (4.14)
for all $t\in[0,T_{0}]$. Hereby, $a$ is given by (A) and $T_{0}$ depends on $a$ and $\langle 1,\mu\rangle$.
Remark 4.4.
Note that in the case $a>0$, we can prove (4.14) for all $t>0$.
Note that $(Y,J)$ are the limits of $(J^{n},Y^{n})$ in (4.13) that we discussed above with $d_{1},d_{2}$ as in (A).
Finally, when $\nu_{2}\equiv 0$ and $h_{\infty}\equiv\langle\nu_{1},1\rangle$ we can use Proposition 4.3 to show that the cozero set of solutions to (1.1) has finite Lebesgue measure. That is, we obtain the following result.
Theorem 4.5.
Suppose that almost surely, $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$. And let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $\nu_{n}^{2}\equiv 0$ and $b_{0}=0$. Then it follows that there exists $T_{0}>0$ such that for all $t\in[0,T_{0}]$ we have
$\mathbb{P}(\mathrm{Leb}(\{x\in\mathbb{R}\colon X_{t}(x)>0\})<\infty)=1.$
5 Examples
As already mentioned above, Theorem 1.1 establishes weak existence and uniqueness of solutions to (1.1) for discontinuous and non-Lipschitz drifts. Below we give several examples. We are particularly interested in examples that allow us to investigate the behavior of solutions to (1.1) at zero and compare our findings to the corresponding one dimensional stochastic differential equations.
5.0.1 Continuous Non-Lipschitz Drift
Set $\nu^{2}\equiv 0$, $\nu^{1}(d\lambda)=\lambda^{-1-\alpha}\mathbbm{1}_{\lambda\geq 1}$ with $\alpha\in(0,1)$ and
$b_{0},b_{1}=\langle\nu^{1},1\rangle.$
Then it follows that
$h(x)=\int_{1}^{\infty}(1-\operatorname{e}^{-\lambda x})\lambda^{-1-\alpha}d\lambda=|x|^{\alpha}\int_{x}^{\infty}(1-\operatorname{e}^{-\lambda})\lambda^{-1-\alpha}d\lambda.$
Clearly,
$C_{1}(\alpha)|x|^{\alpha}\leq|h(0)-h(x)|=h(x)\leq C_{2}(\alpha)|x|^{\alpha},\quad x\in[0,1],$
is not Lipschitz continuous at 0, but only $\alpha$-Hölder continuous.
5.0.2 Completely Monotone Functions
Let us now consider the case, where $b_{0}=b_{1}=0$. That means, we have
$h(x)=h_{1}(x)=\int_{0}^{\infty}\operatorname{e}^{-\lambda x}(\nu_{2}-\nu_{1})(d\lambda)$ (5.1)
where $\langle\nu_{2}-\nu_{1},1\rangle\geq 0$. By Bernsteins theorem on completely monotone functions (see [4]) all functions in the span of completely monotone functions that are non-negative at zero take the form (5.1). Recall, that a completely monotone function $f$ on $0,\infty)$ is continuous everywhere and in $\mathcal{C}^{\infty}((0,\infty))$ and satisfies
$(-1)^{n}f^{(n)}(x)\geq 0,\quad x>0,$
where $f^{n}$ denotes the $n$-th derivative of $f$.
5.0.3 Discontinuous Drift
In this section we discuss the discontinuous part of the drift. That is we set
$h(x)=h_{b_{0},b_{1}}(x)=b_{1}\mathbbm{1}_{x>0}+b_{0}\mathbbm{1}_{x=0},\quad x\geq 0,$
for $b_{0}\geq 0$ and $b_{1}\in\mathbb{R}$. This example allows us to examine some differences that occur in stochastic partial differential equations of the form (1.1) compared to their scalar counterparts of the form
$dx_{t}=h_{b_{0},b_{1}}(x_{t})dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=y,\quad t\geq 0,$ (5.2)
where $(B_{t})_{t\geq 0}$ is a standard Brownian motion. It turns out that if we choose $b_{0}=0$ and $b_{1}=1$, weak uniqueness fails for (5.2) when $x_{0}=0$. However, (1.1) has a unique weak solution for $h\equiv h_{0,1}$. Furthermore, if $b_{0}=c\in(0,\infty)$ and $b_{1}=1$, then the laws of the solutions to (5.2) are identical for each choice of $c$. The solutions to (1.1) with $h\equiv h_{c,1}$ on the other hand, have distinct laws for distinct values of $c$. Finally, if we consider the drift $h_{1,0}$, then (5.2) has no solution, while (1.1) is weakly well-posed. These observations are the content of the following two lemmas.
Lemma 5.1.
  1. (i)
    Weak uniqueness does not hold for
    $dx_{t}=\frac{1}{2}\mathbbm{1}_{x>0}dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=0,\quad t\geq 0.$ (5.3)
  2. (ii)
    The solutions to
    $dx_{t}=h_{c,\frac{1}{2}}(x_{t})dt+\sqrt{x_{t}}dB_{t},\quad t\geq 0,$ (5.4)
    have the same law for each $c\in(0,1/2)$.
  3. (iii)
    There exists no solution to the equation
    $dx_{t}=\mathbbm{1}_{x_{t}=0}dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=0,\quad t\geq 0.$ (5.5)
Proof.
First we prove $(i)$. Clearly, $x_{t}=0$ for all $t\geq 0$ is a solution to (5.3). Denote by $(y_{t})_{t\geq 0}$ the unique strong solution to (5.2) with $b_{0}=1/2$, $b_{1}=0$ and $y_{0}=0$ given by $Y_{t}=1/2(B_{t})^{2}$. We now show, that $(y_{t})_{t\geq 0}$ also solves (5.3). To this end it is enough to prove that
$\int_{0}^{t}\mathbbm{1}_{y_{s}=0}ds=0.$
However, this can be proven along the same lines as the proof of [3, Lemma 1.3]. That is,by [24, Theorem VI.1.7] we get that
$L_{t}^{0}(y)=2\int_{0}^{t}\mathbbm{1}_{y_{s}=0}ds,$
where $L_{t}^{0}(y)$ is the local time of $y$ at 0. By the same arguments as in the proof of [24, Proposition XI.1.5] we get that $L_{t}^{0}(y)\equiv 0$. This finishes the proof of $(i)$.
Next we prove $(ii)$. In order to prove the claim, we show that any solution to (5.4) with $c\in(0,1/2)$ solves
$dy_{t}=\frac{1}{2}dt+\sqrt{y_{t}}dB_{t},\quad t\geq 0.$ (5.6)
Note that (5.6) has pathwise unique solutions and thus also weakly unique solutions. In order to show that any solution $x$ to (5.4) also solves (5.6) it is enough to prove
$\int_{0}^{t}\mathbbm{1}_{x_{s}=0}ds=0.$
This can once again be achieved in a similar manner to the above.
Finally, we prove $(iii)$. Assume that there exists a weak solution $x$ to (5.5) and note that this solution has to be non-negative by a slight modification of [16, Theorem 1.4 ]. Then, similar to above we get that
$\int_{0}^{t}\mathbbm{1}_{x_{s}=0}ds=0.$
However, this means that $(x_{t})_{t\geq 0}$ is also a solution to
$dx_{t}=\sqrt{x_{t}}dB_{t},\quad x_{0}=0.$
However, this equation has a pathwise unique solution by the Yamada Watanabe Theorem, which is given by $x_{t}\equiv 0$. This would imply that
$0=\int_{0}^{t}\mathbbm{1}_{x_{s}=0}ds=t,$
which is a contradiction. Thus, there exists no solution to this equation. ∎
Lemma 5.2.
The laws of solutions to
$dX_{t}=\frac{1}{2}\Delta X_{t}+h_{c,1}(X_{t})dt+\sqrt{X_{t}}\dot{W},\quad x_{0}=f\in\mathcal{C}_{b},\quad t\geq 0,$ (5.7)
are different for different values of $c\in(0,1)$.
Proof.
The proof is similar to the proof of [3, Lemma 1.2]. Fix $c_{1},c_{2}\in(0,1)$ with $c_{1}\neq c_{2}$ and denote by $X^{1},X^{2}$ the solution to (5.7) with drift $h_{c_{1},1}$ and $h_{c_{2},1}$, respectively. We prove the claim by contradiction. To this end assume that $X^{1}\overset{d}{=}X^{2}$. Then it follows along the same lines as in the prove of [3, Lemma 1.2] that
$\begin{split}c_{1}\mathbb{E}\Big{[}\int_{\mathbb{R}}\int_{0}^{t}\phi(x)\mathbbm{1}_{X_{s}^{1}(x)=0}dsdx\Big{]}=c_{2}\mathbb{E}\Big{[}\int_{\mathbb{R}}\int_{0}^{t}\phi(x)\mathbbm{1}_{X_{s}^{2}(x)=0}dsdx\Big{]}\end{split}$
for all non-negative $\phi\in\mathcal{C}_{c}^{\infty}$ and all $t\geq 0$. In order to finish the prove it is enough to show that
$\mathbb{E}\Big{[}\int_{\mathbb{R}}\int_{0}^{t}\phi(x)\mathbbm{1}_{X_{s}^{1}(x)=0}dsdx\Big{]}>0,$
since we assumed that $X^{1}$ and $X^{2}$ have the same law. To this end we prove
$\mathbb{E}[\mathbbm{1}_{X_{t}^{i}(x)=0}]>0,\quad\forall t>0,x\in\mathbb{R},i=1,2,$
by a comparison argument, which follows along similar lines to the proof of [3, Lemma 1.2]. Let $X$ be the unique weak solution to
$dX_{t}=\frac{1}{2}\Delta X_{t}+1dt+\sqrt{X_{t}}\dot{W},\quad x_{0}=f,\quad t\geq 0,$
which is the density of the Super-Brownian motion with constant immigration. Furthermore, recall that $h_{1,1}\equiv 1$ and note that
$h_{c_{i},1}(x)\leq h_{1,1}(x),\quad\forall x\in\mathbb{R}$
And that $X^{i}$ as well as $X^{1}$ can be constructed as weak limits of an SPDE with Lipschitz continuous coefficients. Thus, by [25, Corollary 2.4] and weak uniqueness we get that $X$ stochastically dominates $X^{i}$, which implies
$\mathbb{E}[\mathbbm{1}_{X_{t}^{i}(x)=0}]\geq\mathbb{E}[\mathbbm{1}_{X_{t}(x)=0}],\quad\forall t>0,x\in\mathbb{R},i=1,2.$
However, using the duality formula for the Super Brownian motion with immigration given by (1.7) we get for $t>0$ that
$\begin{split}&\mathbb{E}[\mathbbm{1}_{X_{t}(x)=0}]=\lim_{m\rightarrow\infty}\mathbb{E}[\exp(-mX_{t}(x))]\\ &=\lim_{m\rightarrow\infty}\exp\Big{(}-\langle V_{t}(m\delta_{x}),X_{0}\rangle-\int_{0}^{t}\langle V_{s}(m\delta_{x}),1\rangle ds\Big{)}.\end{split}$
Using Proposition 3.6 we get that
$\begin{split}&\lim_{m\rightarrow\infty}\exp\Big{(}-\langle V_{t}(m\delta_{x}),X_{0}\rangle-\int_{0}^{t}\langle V_{s}(m\delta_{x}),1\rangle ds\Big{)}\\ &=\exp\Big{(}-\langle\mathcal{W}_{t}(x,0),X_{0}\rangle-\int_{0}^{t}\langle\mathcal{W}_{s}(x,0),1\rangle ds\Big{)}>0,\end{split}$
where $\mathcal{W}(x,0)$ is defined as in (3.4). This completes the proof. ∎
6 Duality
In this section we construct the dual process $Y$. The construction is in many ways inspired by ideas from [21] and [22]. Since the jumps of $Y$ take infinite values, we consider measures on $\overline{\mathbb{R}}_{+}=\mathbb{R}_{+}\cup\{\infty\}$, which is endowed with the metric
$d(x,y)=|\arctan(x)-\arctan(y)|,$
with the convention that $\arctan(\infty)=\pi/2$. We denote
$\nu=\nu^{1}+\nu^{2}+d_{1}\delta_{\infty}+d_{2}\delta_{\infty}$
where $d_{1},d_{2}$ are defined as in (A) and
$\nu([b,\infty])=(\nu^{1}+\nu^{2})([b,\infty))+d_{1}+d_{2},\quad b\in\overline{\mathbb{R}}_{+}.$
Furthermore, assume that $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu$ is a non-negative finite measure, $m\in\mathbb{R}_{+}\cup\{\infty\}$, and $x_{0}\in\mathbb{R}$. We construct the process iteratively. For this purpose, let $S_{1}$ be an $\exp(1)$ distributed random variable and denote
$T_{1}=\inf\{t>0|\langle\nu((0,\infty])\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds>S_{1}\}.$
Furthermore, let $M_{1}$ be an independent random variable with values in $\{1,2\}$ given by
$\mathbb{P}(M_{k}=i)=\frac{\nu^{i}((0,\infty))+d_{i}}{\nu((0,\infty])},\quad i\in\{1,2\},$ (6.1)
and let $Z_{1}$ be a random variable such that
$\mathbb{P}(Z_{1}\geq b|M_{1})=\frac{\nu^{M_{1}}([b,\infty))\mathbbm{1}_{b<\infty}+d_{M_{1}}}{\nu^{M_{1}}((0,\infty))+d_{M_{1}}},\quad b\in\overline{\mathbb{R}}_{+}.$
Finally we also define
$\mathbb{P}(U_{1}\in A|Y_{T_{1}-})=\frac{\int_{A}Y_{T_{1}-}(x)dx}{\int_{\mathbb{R}}Y_{T_{1}-}(x)dx},$
for $A\in\mathcal{B}(A)$. Then we set
$\begin{cases}Y_{t}=V_{t}(Y_{0}),\quad 0<t<T_{1},\\ Y_{T_{1}}=V_{T_{1}-}(Y_{0})+Z_{1}\delta_{U_{1}},\quad\text{if }T_{1}<\infty.\end{cases}$
Note, the the process takes values in $\mathcal{M}_{reg}^{+}$ at $Y_{T_{1}}$. We then continue this construction, by letting $(M_{k})_{k=2}^{\infty}$ be i.i.d. copies of $M_{1}$, letting $(S_{k})_{k=1}^{\infty}$ i.i.d $\exp(1)$ distributed random variables, independent of $(M_{k})_{k=1}^{\infty}$. Now we define
$T_{k}=\inf\{t>0|\nu((0,\infty])>\int_{0}^{t}\int_{\mathbb{R}}V_{s}(Y_{T_{k-1}})dxds>S_{k}\},\quad\bar{T}_{k}=\sum_{i=1}^{k}T_{i}$ (6.2)
for $k\geq 2$ and set $\bar{T}_{0}=T_{0}=0$. Moreover, we define random variables $(U_{k})_{k=1}^{\infty}$ as
$\mathbb{P}(U_{k}\in A|Y_{\bar{T}_{k}-})=\frac{\int_{A}Y_{\bar{T}_{k}-}(x)dx}{\int_{\mathbb{R}}Y_{\bar{T}_{k}-}(x)dx}$ (6.3)
for $A\in\mathcal{B}(\mathbb{R})$ and
$\mathbb{P}(Z_{k}\geq b|M_{k})=\frac{\nu^{M_{k}}([b,\infty))\mathbbm{1}_{b<\infty}+d_{M_{k}}}{\nu^{M_{k}}((0,\infty))+d_{M_{k}}},\quad b\in\overline{\mathbb{R}}^{+}.$ (6.4)
Hereby, $(Z_{k},M_{k})$ and $(T_{k},U_{k})$ are independent for each $k\geq 1$. Finally, we set
$\mathcal{F}_{t}=\cap_{\epsilon>0}\sigma(Y_{s}(x),s\leq t+\epsilon,x\in\mathbb{R})$
and we choose $(Z_{k})_{k\geq 1}$ and $(S_{k})_{k\geq 1}$and $(M_{k})_{k\geq 1}$ in such a way that $(Z_{\ell},S_{\ell},M_{\ell})_{\ell\geq k}$ is independent of $\mathcal{F}_{\bar{T}_{k-1}}$ for all $k\geq 1$.
We iteratively set
$\begin{cases}Y_{t}=V_{t-\bar{T}_{k-1}}(Y_{T_{k-1}}),\quad\bar{T}_{k-1}<t<\bar{T}_{k},\\ Y_{\bar{T}_{k}}=V_{T_{k}}(Y_{\bar{T}_{k-1}})+Z_{k}\delta_{U_{k}},\quad\text{ if }\bar{T}_{k}<\infty\end{cases}$ (6.5)
and continue this construction for all $k\geq 0$. Again, note that this process takes values in $\mathcal{M}_{reg}^{+}$. The next section demonstrates that $\bar{T}_{k}\rightarrow\infty$ as $k\rightarrow\infty$.
6.1 An Artificial Branching Process
The aim of this section is to derive bounds for $Y$, which only depend on $Y_{0}$ and $\nu((0,\infty])$. These bounds are crucial in order to prove that $Y$ only has finitely many jumps in each finite time interval and in order to show that all the terms in (4.14) have finite moments. To this end, we introduce an artificial branching mechanism that will allow us to keep track of the number of jumps of $Y$. This approach is inspired by the ideas from [3].
Let
$I_{t}(Y_{0})\subset\mathcal{U}=\cup_{k=1}^{\infty}\mathbbm{N}^{k}$
be the collections of labels of particles that are alive at time $t$. And denote by
$|I_{t}(Y_{0})|,$
the number of particles in $I_{t}(Y_{0})$ for all $t\geq 0$. The $i$-th child of the particle with label $\alpha=(\alpha_{1},\ldots\alpha_{m})$ is labeled $\beta=(\alpha_{1},\ldots\alpha_{m},i)$. Below we give a description of the branching mechanism.
  1. 6.1.
    We denote the first particle by $\mathcal{Z}^{(1)}=(Y^{(1)},\mathfrak{M}^{(1)})$, where $Y^{(1)}_{t}=V_{t}(Y_{0})$ for $t\geq 0$ and we let $\mathfrak{M}^{(1)}(t)$ for $t\geq 0$ be a Poisson point process on $(0,\infty]\times(0,\infty)^{2}\times\mathbb{R}$ with intensity
    $\nu(d\lambda)dtdrdx.$ (6.6)
    We furthermore denote by $\mathfrak{N}^{(1)}$ the random measure defined by
    $\mathfrak{N}^{(1)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{Y_{t}^{(1)}(x)\leq r}dxdrdt,$
    for $G_{1}\subseteq(0,\infty],G_{2},G_{3}\subseteq(0,\infty)$and $G_{4}\subseteq\mathbb{R}$ are measurable sets.
  2. 6.2.
    To every particle $\mathcal{Z}^{(\alpha)}$ that was born at time $\tau_{\alpha}$ we associate a tuple $(Y^{\alpha},\mathfrak{M}^{\alpha})$, where $Y^{(\alpha)}_{t}=\mathcal{W}_{t-\tau_{\alpha}-}(0,0)\mathbbm{1}_{t>\tau_{\alpha}}$. Furthermore, $\mathfrak{M}^{(\alpha)}$ is a Poisson point process on $(0,\infty]\times(0,\infty)^{2}\times\mathbb{R}$ with intensity given by (6.6). Hereby, the processes $\mathfrak{M}^{(\alpha)}$ are independent. Moreover, we define the random measure $\mathfrak{N}^{(\alpha)}$ given by
    $\mathfrak{N}^{(\alpha)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{Y_{t}^{(\alpha)}(x)\leq r}dxdrdt,$
    for $G_{1}\subseteq(0,\infty],G_{2},G_{3}\subseteq(0,\infty)$and $G_{4}\subseteq\mathbb{R}$ are measurable sets.
  3. 6.3.
    A particle $\mathcal{Z}^{(\alpha)}$, $\alpha\in\mathcal{U}$ gives birth to a new particles at time $\tau$, which is an arrival time of $\mathfrak{N}^{(\alpha)}$. If $\tau$ is the time of the $i$-th branching of $\mathcal{Z}^{(\alpha)}$ then the new particle will be labeled $\beta=(\alpha,i)$ and the parent keeps the label $\alpha$.
Recall that we denote by $I_{t}(Y_{0})$ the set of particles that are alive at time $t$. That is we have
$I_{t}(Y_{0})=\{\alpha\in\mathcal{U}\colon\langle Y_{t}^{(\alpha)},1\rangle>0\}.$
The next lemmas allow us to control the moments of the number of jumps of the dual process.
Lemma 6.1.
For all $T>0$ there exists $C=C(T)>0$ such that
$\mathbb{E}[|I_{t}(Y_{0})|]<C,$
for all $t\leq T$.
Proof.
First we prove the statement for $T_{0}>0$ sufficiently small. First, we define
$I_{t}^{0}=\{(1)\},\quad I_{t}^{i}=\{\alpha\in\mathcal{U}||\alpha|=i+1,\langle Y_{t}^{(\alpha)},1\rangle>0\},\quad i\geq 1$
where $|\alpha|$ denotes the length of the vector $\alpha$. Let $T_{0}>0$ be chosen sufficiently small and $t\leq T_{0}$. The precise value of $T_{0}$ will be determined later. We will show that
$\mathbb{E}{|I_{t}^{i}|}\leq 2^{-i},\quad 0\leq t\leq T_{0}.$ (6.7)
First, we prove (6.7) for the case $i=1$. The expected number of children of the particle $Z^{(1)}$ until time $t\in[0,T_{0}]$ is given by the expectation of $\mathfrak{N}^{(1)}_{t}=\mathfrak{N}^{(1)}((0,\infty],(0,t],(0,\infty),\mathbb{R})$, that is
$\begin{split}&\mathbb{E}{|I_{t}^{1}|}=\mathbb{E}[\mathfrak{N}^{(1)}_{t}]\\ &=\nu((0,\infty])\mathbb{E}\Big{[}\int_{0}^{t}\int_{\mathbb{R}}Y_{s}^{(1)}(x)dxds\Big{]}=\nu((0,\infty])\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds.\end{split}$
Since $(t,x)\mapsto V_{t}(Y_{0})(x)$ is integrable on $0,T]\times\mathbb{R}$ for all $T>0$, we can find $T_{0}>0$ sufficiently small so that
$\mathbb{E}{|I_{t}^{1}|}\leq 2^{-1}.$
Now suppose that (6.7) holds for $i\geq 1$. Denote by $(\tilde{\mathfrak{N}}^{(\alpha)})_{\alpha\in\mathcal{U}}$ a family of independent Poisson processes on $(0,\infty)$ with intensity given by
$\nu((0,\infty])\int_{\mathbb{R}}\mathcal{W}_{t}(0,0)(x)dxdt.$
Then it follows that
$\mathbb{E}[|I_{t}^{{i+1}}|]=\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}^{i}}\mathfrak{N}^{(\alpha)}_{t}\Big{]}\leq\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}^{i}}\tilde{\mathfrak{N}}^{(\alpha)}_{t}\Big{]}\leq\mathbb{E}[|I_{t}^{i}|]\nu((0,\infty])\int_{0}^{t}\int_{\mathbb{R}}\mathcal{W}_{s}(0,0)(x)dxds\leq 2^{-i}2^{-1},$
if we choose $T_{0}$ sufficiently small, so that
$\nu((0,\infty])\int_{0}^{T_{0}}\int_{\mathbb{R}}\mathcal{W}_{s}(0,0)(x)dxds\leq 2^{-1}.$
Hereby, we used that the intensity of $\tilde{\mathfrak{N}}^{(\alpha)}$ is larger than the intensity of $\mathfrak{N}^{(\alpha)}((0,\infty],\cdot,(0,\infty),\mathbb{R})$ for each $\alpha$. Thus, we arrive at
$\mathbb{E}[|I_{t}|]\leq\sum_{i=0}^{\infty}\mathbb{E}[|I_{t}^{i}|]\leq 1+\sum_{i=1}^{\infty}2^{-i},$
for all $0\leq t\leq T_{0}$.
Now, we need to prove the claim for general $T>0$. For every $T>0$ we can find $M\in\mathbb{N}$ such that $T\leq MT_{0}$. Then, using the fact that no particle dies, we can estimate
$\mathbb{E}[|I_{T}(Y_{0})|]\leq\mathbb{E}[|I_{MT_{0}}(Y_{0})|]\leq\mathbb{E}[\sum_{\alpha\in I_{T_{0}}}|\tilde{I}_{T_{0},MT_{0}}^{\alpha}|],$
where
[…]
Note that for $\alpha\neq(1)$ it holds that
$\mathbb{E}[|I^{\alpha}_{T_{0},MT_{0}}|]\leq\mathbb{E}[|I_{(M-1)T_{0}}(\mathcal{W})|],$
where $I_{t}(\mathcal{W})$ denotes the particles that are alive at time $t$ in the branching process described in 6.1.-6.3. where $Y^{(1)}$ is replaced by $\mathcal{W}(0,0)$. This simply follows from the fact that
$\int_{0}^{t}\langle\mathcal{W}_{s+r}(0,0),1\rangle dr\leq\int_{0}^{t}\langle\mathcal{W}_{r}(0,0),1\rangle dr,\quad\forall s,t\geq 0,$
which follows from Proposition 3.7. Thus, we iteratively get that
$\begin{split}&\mathbb{E}[|I_{T}(Y_{0})|]\leq\mathbb{E}[|I_{MT_{0}}(Y_{0})|]\\ &\leq\mathbb{E}[|I_{T_{0}}(Y_{0})|-1]\mathbb{E}[|I_{(M-1)T_{0}}(\mathcal{W})|]+\mathbb{E}[|I_{(M-1)T_{0}}(V_{T_{0}}(Y_{0}))|]\leq\ldots\leq C(T)\end{split}$
This finishes the proof. ∎
Lemma 6.2.
Let $Y_{0}=\mu+m\delta_{x_{0}}$ with $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then for every $\gamma>0$, there exists $T_{0}>0$ such that
$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}(Y_{0})||}]<\infty,$ (6.8)
for all $t\leq T_{0}$ . Hereby, $T_{0}>0$ only depends on $\nu((0,\infty])$, $\langle\mu,1\rangle$ and $\gamma$.
Proof.
Fix $t\leq T_{0}$, where $T_{0}$ is sufficiently small, the precise value will be chosen later. Define $\eta_{t}^{i,k}$ for $i,k\in\mathbb{N}$ to be i.i.d. Poisson distributed random variables with parameter
$\lambda=\nu((0,\infty])\Big{(}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds\Big{)}.$
Furthermore, set $Z_{0}=1$ and
$Z_{n}^{t}=\sum_{k=1}^{Z_{n-1}^{t}}\eta_{t}^{n,k},\quad n\geq 1.$
Then it easily follows that $|I_{t}^{i}|$ is stochastically dominated by $Z_{i}^{t}$ for each $i\geq 1$. Thus, we can also infer that
$|I_{t}|=\sum_{i=0}^{\infty}|I_{t}^{i}|$
is stochastically dominated by
$Z^{t}=\sum_{i=0}^{\infty}Z_{i}^{t}.$
The random variable $Z^{t}$ is nothing else, but the total progeny of $(Z^{t}_{n})_{n\geq 0}$ and its law can be made explicit by using the Otter-Dwass formula (see [9]), whenever $\lambda\leq 1$. Thus $T_{0}$ must be small enough such that $\lambda\leq 1$. In our setting, we get
$\mathbb{P}(Z^{t}=k)=\operatorname{e}^{-\lambda k}\frac{(\lambda k)^{k-1}}{k!},\quad k\geq 1.$
which is the Borel-Tanner distribution. Using Stirling’s approximation, we thus get that
$\mathbb{P}(Z^{t}=k)\asymp\sqrt{2\pi}\operatorname{e}^{(1-\lambda)k}\lambda^{k-1}k^{-3/2}\asymp\sqrt{2\pi}\operatorname{e}^{(1-\lambda+\log(\lambda))k-\log(\lambda)}k^{-3/2}$
Thus,
$\sqrt{2\pi}\operatorname{e}^{-\log(\lambda)}\sum_{k=1}^{\infty}k^{-3/2}\operatorname{e}^{(\gamma+1-\lambda+\log(\lambda))k}<\infty\quad\Rightarrow\quad\sum_{k=1}^{\infty}e^{\gamma k}\mathbb{P}(Z^{t}=k)<\infty$
and
$\sum_{k=1}^{\infty}k^{-3/2}\operatorname{e}^{(\gamma+1-\lambda+\log(\lambda))k}<\infty,$
whenever
$\gamma\leq\lambda-1-\log(\lambda).$ (6.9)
When $\lambda\rightarrow 0$, we get that $\lambda-1-\log(\lambda)\rightarrow\infty$.
If we choose $T_{0}>0$ small enough, we thus get that (6.9) is satisfied and we obtain that
$\mathbb{E}[\operatorname{e}^{\gamma Z^{t}}]<\infty.$
Note that the function $x\mapsto\operatorname{e}^{\gamma x}$ is non-decreasing. Thus, the fact that $|I_{t}|$ is stochastically dominated by $Z^{t}$ implies that
$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}|}]\leq\mathbb{E}[\operatorname{e}^{\gamma Z^{t}}]<\infty.$
This completes the proof. ∎
Now, we need to relate the branching process to the process $Y$. Since the birth times of new particles happen with higher frequency than the jumps of $t\mapsto\langle Y_{t},1\rangle$, we aim to control its time integral, which is sufficient for our arguments. Recall that $(S_{i})_{i=1}^{\infty}$ is the sequence of independent $\exp(1)$-distributed random variables from the construction of $Y$ as defined in (6.2). We now define random variables $(\bar{R}_{i})_{i=1}^{\infty}$ as follows
$\begin{split}&\bar{R}_{0}=0,\\ &\bar{R}_{i}=\inf\{t>\bar{R}_{i-1}\colon\sum_{\alpha\in I_{\bar{R}_{i-1}}}\int_{\bar{R}_{i-1}}^{t}\langle Y^{(\alpha)}_{s},1\rangle ds\nu((0,\infty])>S_{i}\},\quad i\geq 1.\end{split}$
and set
$R_{i}=\bar{R}_{i}-\bar{R}_{i-1},\quad i\geq 1.$
Furthermore set
$\hat{Y}_{t}=\sum_{\alpha\in I_{t}}Y_{t}^{(\alpha)},\quad t\geq 0.$
Then, by construction we get the following identity
$\hat{Y}_{t}=V_{t}(Y_{0})+\sum_{k=1}^{|I_{t}|}\mathcal{W}_{t-\bar{R}_{k}}(0,0).$ (6.10)
Using this construction we can prove our first estimates.
Lemma 6.3.
For all $i\geq 1$ the following inequalities hold.
  1. (i)
    It holds almost surely that
    $R_{i}\leq T_{i}$ (6.11)
  2. (ii)
    It holds almost surely that
    $\langle Y_{\bar{T}_{i-1}+s},1\rangle\leq\langle\hat{Y}_{\bar{R}_{i-1}+s},1\rangle,\quad\forall 0<s\leq T_{i}$ (6.12)
  3. (iii)
    It holds almost surely that
    $\int_{0}^{t}\langle Y_{s},1\rangle ds\leq\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds,\quad\forall t\in(\bar{T}_{i-1},\bar{T}_{i}).$ (6.13)
Proof.
We prove the claim by induction over $i\geq 1$. It is easy to see that $R_{1}=T_{1}$ and then (6.12) and (6.13) are immediate for $i=1$. Suppose that (6.11), (6.12) and (6.13) hold for all $j\leq i$ for some $i\geq 1$. We now show that this implies the claim for $i+1$. First, we prove (6.12). We apply Lemma 3.3 repeatedly to arrive at
$\begin{split}\langle Y_{\bar{T}_{i+s}},1\rangle&\leq\langle V_{s}(Y_{\bar{T}_{i}-}),1\rangle+\langle\mathcal{W}_{s}(0,0),1\rangle\\ &\leq\langle V_{s+T_{i}}(Y_{\bar{T}_{i-1}-}),1\rangle+\langle\mathcal{W}_{s+\bar{T}_{i}}(0,0),1\rangle+\langle\mathcal{W}_{s}(0,0),1\rangle\\ &\leq\langle V_{s+\bar{T}_{i}}(Y_{0}),1\rangle+\sum_{k=1}^{i}\langle\mathcal{W}_{s+\bar{T}_{i}-\bar{T}_{k}}(0,0),1\rangle,\quad a.s..\end{split}$
By the induction hypothesis, we know that $R_{j}\leq T_{j}$ for all $j\leq i$. Thus, we have
$\bar{T}_{i}-\bar{T}_{k}=\sum_{\ell=k+1}^{i}T_{\ell}\geq\bar{R}_{i}-\bar{R}_{k},$
for all $0\leq k\leq i$. By the monotonicity relations
$\langle\mathcal{W}_{s}(0,0),1\rangle\leq\langle\mathcal{W}_{t}(0,0),1\rangle\quad 0<t\leq s,$ (6.14)
$\langle V_{s}(Y_{0}),1\rangle\leq\langle V_{t}(Y_{0}),1\rangle\quad 0<t\leq s,$ (6.15)
we therefore arrive at
$\begin{split}&\langle Y_{\bar{T}_{i}+s},1\rangle\leq\langle V_{s+\bar{R}_{i}}(Y_{0}),1\rangle+\sum_{k=1}^{i}\langle\mathcal{W}_{s+\bar{R}_{i}-\bar{R}_{k}}(0,0),1\rangle\\ &\leq\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle\quad a.s.,\end{split}$
for all $0<s<\bar{T}_{i+1}$. Hereby, (6.14) follows from Proposition 3.7 and (6.15) follows from from integrating (3.5).
Now, we show
$R_{i+1}\leq T_{i+1}$ (6.16)
Recall that
$T_{i+1}=\inf\{t\geq 0\colon\nu((0,\infty])\int_{0}^{t}\langle Y_{\bar{T}_{i}+s},1\rangle ds>S_{i+1}\}$
and
$R_{i+1}=\inf\{t\geq 0\colon\nu((0,\infty])\int_{0}^{t}\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle ds>S_{i+1}\}$
Therefore, (6.16) is immediate by (6.12) for $i+1$.
Finally, we need to verify
$\int_{0}^{t}\langle Y_{s},1\rangle ds\leq\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds,\quad\forall t\in(\bar{T}_{i},\bar{T}_{i+1}).$
Let $t\in(\bar{T}_{i},\bar{T}_{i+1})$ be fixed but arbitrary and define
$j^{\ast}=\sup\{k\geq 1\colon\bar{R}_{k}\leq t\}.$
Since $t\in(\bar{T}_{i},\bar{T}_{i+1})$ and $\bar{T}_{k}\geq\bar{R}_{k}$ for all $k\leq i+1$ by (6.16) we get $j^{\ast}\geq i$. Note that
$\begin{split}&\nu((0,\infty])\int_{0}^{t}\langle Y_{s},1\rangle ds\\ &=\nu((0,\infty])\Big{(}\int_{0}^{\bar{T}_{i}}\langle Y_{s},1\rangle ds+\int_{\bar{T}_{i}}^{t}\langle Y_{s},1\rangle ds\Big{)}=\sum_{k=1}^{i}S_{k}+\nu((0,\infty])\int_{\bar{T}_{i}}^{t}\langle Y_{s},1\rangle ds,\end{split}$
and
$\begin{split}&\nu((0,\infty])\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds\\ &=\nu((0,\infty])\Big{(}\int_{0}^{\bar{R}_{j^{\ast}}}\langle\hat{Y}_{s},1\rangle ds+\int_{\bar{R}_{j^{\ast}}}^{t}\langle\hat{Y}_{s},1\rangle ds\Big{)}=\sum_{k=1}^{j\ast}S_{k}+\nu((0,\infty])\int_{\bar{R}_{j^{\ast}}}^{t}\langle\hat{Y}_{s},1\rangle ds.\end{split}$
Thus, when $j^{\ast}>i$, the proof is finished. In the case when $j^{\ast}=i$ it is enough to verify that
$\int_{\bar{T}_{i}}^{t}\langle Y_{s},1\rangle ds\leq\int_{\bar{R}_{i}}^{t}\langle\hat{Y}_{s},1\rangle ds.$
Using (6.12) for $i+1$ and using that $\bar{R}_{i}\leq\bar{T}_{i}$ we get
$\begin{split}&\int_{\bar{T}_{i}}^{t}\langle Y_{s},1\rangle ds=\int_{0}^{t-\bar{T}_{i}}\langle Y_{\bar{T}_{i}+s},1\rangle ds\leq\int_{0}^{t-\bar{T}_{i}}\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle ds\\ &\quad\leq\int_{0}^{t-\bar{R}_{i}}\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle ds=\int_{\bar{R}_{i}}^{t}\langle\hat{Y}_{s},1\rangle ds.\end{split}$
This finishes the proof. ∎
With these estimates at hand, we can prove the following proposition.
Proposition 6.4.
Let $Y$ be the process constructed in (6.5) with $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then it follows that
$\sup_{n}\mathbb{E}\Big{[}\int_{0}^{T}\langle Y_{t},1\rangle dt\Big{]}\leq C(T),$ (6.17)
where $C(T)$ only depends on $\nu((0,\infty])$ and $\langle\mu,1\rangle$.
Proof.
Since $\bar{R}_{i}\rightarrow\infty$ as $i\rightarrow\infty$ we get by the monotone convergence theorem and using (6.13) that
$\begin{split}&\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{]}=\lim_{i\rightarrow\infty}\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\mathbbm{1}_{t\leq\bar{T}_{i}}\Big{]}\\ &\leq\mathbb{E}[\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds].\end{split}.$
By using Lemma 6.1, this immediately yields the following inequality
$\begin{split}&\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{]}\leq\mathbb{E}\Big{[}\int_{0}^{t}\sum_{\alpha\in I_{s}}\langle Y_{s}^{(\alpha)},1\rangle ds\Big{]}\\ &\leq\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s}^{(1)},1\rangle ds\Big{]}+\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds\Big{]}\\ &\leq\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds+C\mathbb{E}[I_{t}]\leq C(t),\end{split}$
where $C(t)>0$ only depends on $\nu((0,\infty])$ and $\langle\mu,1\rangle$. ∎
Lemma 6.5.
Let $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then, for every $\gamma\in\mathbb{R}$, there exists $T_{0}>0$ such that for all $t\leq T_{0}$ we have
$\mathbb{E}\Big{[}\exp\Big{(}-\gamma\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{)}\Big{]}<C,$
Hereby, $T_{0}>0$ and $C>0$ only depend on $\langle\mu,1\rangle$ and $\nu((0,\infty])$.
Proof.
If $\gamma\geq 0$, the statement follows immediately. Thus, we only need to consider the case $\gamma<0$. By Lemma 6.2 and (6.13) and using that
$\int_{0}^{t}\langle Y^{\alpha}_{s},1\rangle ds\leq\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds,\quad\forall\alpha\in I_{t},t\geq 0,$
we get that there exists $T_{0}>0$ such that
$\mathbb{E}\Big{[}\exp\Big{(}-\gamma\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{)}\Big{]}\leq\mathbb{E}\Big{[}\exp\Big{(}-\gamma|I_{t}|\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds-\gamma\int_{0}^{t}\langle V_{s}(\mu),1\rangle ds\Big{)}\Big{]}<C,$
for all $t\leq T_{0}$. This finishes the proof. ∎
6.2 Convergence of the Dual Process
Denote by $(Y_{t}^{n,m}(Y_{0}^{m}))_{t\geq 0}$ the process as constructed in (6.5) with initial condition $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$ and $n\in\mathbb{N}\cup\{\infty\}$, $m\in\mathbb{R}_{+}\cup\{\infty\}$. We assume that $Y_{0}^{n,m}=Y_{0}^{m}$, $\langle Y_{0}^{m},1\rangle>0$ and that, for $n\in\mathbb{N}$, $\nu$ is replaced by
$\nu_{n}=\nu^{1}_{|[0,n]}+\nu^{2}_{|[0,n]}+d_{1}\delta_{n}+d_{2}\delta_{n}=\nu^{1}_{n}+\nu^{2}_{n}+d_{1}\delta_{n}+d_{2}\delta_{n}.$
Here, $d_{1}$, $d_{2}$ and $a$ are defined as in (A). For $i\geq 1$, we denote by $\bar{T}_{n,i}^{m}$ the jump times, by $T_{n,i}=\bar{T}_{n,i}^{m}-\bar{T}_{n,i-1}^{m}$ the time in between jumps, by $Z_{n,i}^{m}$ the jump sizes and by $U_{n,i}^{m}$ the jump positions defined as in (6.2)-(6.4). Moreover, we denote by $M_{n,i}^{m}$ the random variables defined as in (6.1). Finally, we shall denote by $(S_{n,k}^{m})_{k\geq 1}$ the sequence independent of $\exp(1)$-distributed random variables from the construction of the process $Y^{n,m}$ in the beginning of Section 6.
In the remainder of the section we prove that
$\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{m,m}}\exp\Big{(}-\langle Y_{t}^{m,m},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{m,m},1\rangle ds\Big{)}\Big{]}\\ &\quad=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y_{t}^{\infty,\infty},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,\infty},1\rangle ds\Big{)}\Big{]},\end{split}$
and
$\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{n,m}}\exp\Big{(}-\langle Y_{t}^{n,m},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{n,m},1\rangle ds\Big{)}\Big{]}\\ &\quad=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,m}}\exp\Big{(}-\langle Y_{t}^{\infty,m},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,m},1\rangle ds\Big{)}\Big{]}.\end{split}$
and
$\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,m}}\exp\Big{(}-\langle Y_{t}^{\infty,m},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,m},1\rangle ds\Big{)}\Big{]}\\ &\quad=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y_{t}^{\infty,\infty},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,\infty},1\rangle ds\Big{)}\Big{]},\end{split}$
for $t\in[0,T_{0}]$, where $T_{0}$ is chosen so that (6.8) holds with $\gamma=2a$ and where $X_{0}$ is a bounded and continuous function on $\mathbb{R}$.
Note, that $T_{0}$ can be chosen independent of $n,m$ since
$\sup_{n}\nu_{n}((0,\infty))\leq\nu((0,\infty]).$
Before we begin the proof, we introduce the following notation. For $\boldsymbol{x}=(x_{1},\ldots x_{k})$ for some $k\geq 2$ we denote $\overset{\leftarrow}{\boldsymbol{x}}=(x_{1},\ldots x_{k-1})$. Furthermore, for fixed $\mu\in\mathcal{M}_{F}^{+}$ define
$\mathfrak{V}^{1}(u,z,t)=V_{t}(\mu+z\delta_{u}),\quad(u,z,t)\in\mathbb{R}\times\overline{\mathbb{R}}_{+}\times(0,\infty)$
and for $k>1$ we define recursively
$\mathfrak{V}^{k}(\boldsymbol{u},\boldsymbol{z},\boldsymbol{t})=V_{t_{k}}(\mathfrak{V}^{k-1}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\overset{\leftarrow}{\boldsymbol{t}})+z_{k}\delta_{u_{k}}),\quad(\boldsymbol{u},\boldsymbol{z},\boldsymbol{t})\in\mathbb{R}^{k}\times\overline{\mathbb{R}}_{+}^{k}\times(0,\infty)^{k}.$
Recall, that $\overline{\mathbb{R}}_{+}=\mathbb{R}_{+}\cup\{\infty\}$ is endowed with the metric
$d(x,y)=|\arctan(x)-\arctan(y)|,$
with the convention that $\arctan(\infty)=\pi/2$. We will use $\mathfrak{V}^{k}$ later as follows. For all $k>1$, when $\boldsymbol{u}=(x_{0},U_{n,1}^{m},\ldots U_{n,k-1}^{m})$, $\boldsymbol{z}=(m,Z_{n,1}^{m},\ldots Z_{n,k-1}^{m})$ and $\boldsymbol{t}=(T_{n,1}^{m},\ldots T_{n,k-1}^{m},t)$, then
$\mathfrak{V}^{k}(\boldsymbol{u},\boldsymbol{z},\boldsymbol{t})(x)=Y^{n,m}_{\bar{T}_{n,k-1}^{m}+t}(x),\quad x\in\mathbb{R},$ (6.18)
and for all $t\in(0,T_{n,k}^{m})$.
Lemma 6.6.
For every $k\in\mathbb{N}$ the map
$\mathbb{R}^{k}\times\overline{\mathbb{R}}_{+}^{k}\times(0,\infty)^{k}\ni(\boldsymbol{u},\boldsymbol{z},\boldsymbol{t})\mapsto\mathfrak{V}^{k}(\boldsymbol{u},\boldsymbol{z},\boldsymbol{t})$
is continuous with respect to the topology induced by uniform convergence on compacts.
Proof.
First, let us demonstrate that $\mu_{n}=\psi_{n}+z_{n}\delta_{u_{n}}$, $n\geq 1$ converges $m$-weakly to $\mu\in\mathcal{M}_{reg}^{+}$, if $(\psi_{n})_{n\geq 1}\subset\mathcal{C}(\mathbb{R})$ converges uniformly on compacts to $\psi\in\mathcal{C}(\mathbb{R})$, $(z_{n})_{n\geq 1}\subset\bar{R}_{+}$ converges to $z$ with respect to $d$ and $(u_{n})_{n\geq 1}\subset\mathbb{R}$ converges to $u$. Clearly, $S_{\mu}$ is empty when $z\neq\infty$ and when $z=\infty$ we have $S_{\mu}=\{u\}$. It is easy to verify that in the latter case $\eta^{(\emptyset,\mu_{n})}$ converges $m$-weakly to $\eta^{u,\psi}$. By Proposition 3.6, this demonstrates that
$V(\psi_{n}+z_{n}\delta_{u_{n}})\rightarrow V(\psi+z\delta_{u})\$
uniformly on compacts of $(0,\infty)\times\mathbb{R}$. Furthermore, let $\{t_{n}\}_{n\geq 1}\subset(0,\infty)$ be a sequence that converges to $t\in(0,\infty)$. Then it follows that
$V_{t_{n}}(\psi_{n}+z_{n}\delta_{u_{n}})\rightarrow V_{t}(\psi+z\delta_{u}),\quad\text{as }n\rightarrow\infty,$
uniformly on compacts of $\mathbb{R}$. Now it is straightforward to see that the continuity of $\mathfrak{V}^{k}$ can be proven inductively. ∎
We furthermore introduce, for $k\geq 1$, the map
$\Theta^{k}\colon\mathbb{R}^{k+1}\times(0,\infty)^{k}\times\overline{\mathbb{R}}_{+}^{k+1}\times(0,\infty)\rightarrow\mathbb{R}_{+},$
given by
$(\boldsymbol{u},\boldsymbol{t},\boldsymbol{z},s)\mapsto\inf\{r\geq 0\colon\int_{0}^{r}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})\rangle dt>s\}.$
Note that when $\boldsymbol{u}=(x_{0},U_{n,1}^{m},\ldots U_{n,k}^{m})$, $\boldsymbol{t}=(T_{n,1}^{m},\ldots T_{n,k}^{m})$ and $\boldsymbol{z}=(m,Z_{n,1}^{m},\ldots Z_{n,k}^{m})$ we have
$\Theta^{k}(\boldsymbol{u},\boldsymbol{t},\boldsymbol{z},S_{n,k+1}^{m}/\nu_{n}((0,\infty]))=T^{m}_{n,k+1}.$ (6.19)
Lemma 6.7.
For every $k\geq 1$ the map $\Theta^{k}$ is continuous on $\mathbb{R}^{k+1}\times(0,\infty)^{k}\times\overline{\mathbb{R}}_{+}^{k+1}\times(0,\infty)$.
Proof.
Let $(\boldsymbol{u},\boldsymbol{t},\boldsymbol{z},s)\in\mathbb{R}^{k+1}\times(0,\infty)^{k}\times\overline{\mathbb{R}}_{+}^{k+1}\times(0,\infty)$, denote $\boldsymbol{x}=(\boldsymbol{u},\boldsymbol{t},\boldsymbol{z},s)$ and let $B_{r}(\boldsymbol{x})$ be defined as
$B_{r}(\boldsymbol{x})=\{\boldsymbol{y}\in\mathbb{R}^{k+1}\times(0,\infty)^{k}\times\overline{\mathbb{R}}_{+}^{k+1}\times(0,\infty)\colon d_{k+1}(\boldsymbol{x},\boldsymbol{y})\leq r\},$
where $d_{k+1}$ is the canonical metric on $\mathbb{R}^{k+1}\times(0,\infty)^{k}\times\overline{\mathbb{R}}_{+}^{k+1}\times(0,\infty)$. First, we prove that
$T:=\sup_{\boldsymbol{y}\in B_{1}(\boldsymbol{x})}\Theta^{k}(\boldsymbol{y})<\infty.$ (6.20)
This claim can easily be verified by denoting $\boldsymbol{y}=(\boldsymbol{v},\boldsymbol{\tau},\boldsymbol{w},b)\in\mathbb{R}^{k+1}\times(0,\infty)^{k}\times\overline{\mathbb{R}}_{+}^{k+1}\times(0,\infty)$ and noting that
$\begin{split}&\sup_{\boldsymbol{y}\in B_{1}(\boldsymbol{x})}\Theta^{k}(\boldsymbol{y})=\sup_{\boldsymbol{y}\in B_{1}(\boldsymbol{x})}\inf\{r\geq 0\colon\int_{0}^{r}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt>b\}\\ &\leq\sup_{\boldsymbol{y}\in B_{1}(\boldsymbol{x})}\inf\{r\geq 0\colon\int_{0}^{r}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt>s+1\}\\ &\leq\sup_{\boldsymbol{y}\in B_{1}(\boldsymbol{x})}\inf\{r\geq 0\colon\int_{0}^{r}\langle V_{t+\bar{t}}(\phi),1\rangle dt>s+1\}<\infty,\end{split}$ (6.21)
where
$\bar{t}=\sum_{i=1}^{k}t_{i}+1.$ (6.22)
Hereby, we used Lemma 3.3 and the fact that by Lemma 3.4 we have
$\langle V_{t}(\mu),1\rangle\leq\langle V_{s}(\mu),1\rangle,\quad\forall 0\leq s\leq t\text{ and }\int_{0}^{\infty}\langle V_{t}(\mu),1\rangle ds=\infty$ (6.23)
for all finite, non-zero measures $\mu$. Now, let $\epsilon>0$ be arbitrary, $\delta>0$ sufficiently small and $\boldsymbol{y}=(\boldsymbol{v},\boldsymbol{\tau},\boldsymbol{w},b)\in B_{\delta}(\boldsymbol{x})$. Then we get
$\begin{split}&\Theta^{k}(\boldsymbol{u},\boldsymbol{t},\boldsymbol{z},s)\\ &=\inf\Big{\{}r\geq 0\colon\int_{0}^{r}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt>s+\boldsymbol{\Delta}\theta_{r}^{k}(\overset{\leftarrow}{\boldsymbol{y}},\overset{\leftarrow}{\boldsymbol{x}})\Big{\}},\end{split}$
where
$\overset{\leftarrow}{\boldsymbol{x}}=(\boldsymbol{u},\boldsymbol{t},\boldsymbol{z}),\quad\overset{\leftarrow}{\boldsymbol{y}}=(\boldsymbol{v},\boldsymbol{\tau},\boldsymbol{w}).$
and
$\boldsymbol{\Delta}\theta_{r}^{k}(\overset{\leftarrow}{\boldsymbol{y}},\overset{\leftarrow}{\boldsymbol{x}})=\int_{0}^{r}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}}),1\rangle dt.$
Now, we denote
$\eta(\delta)=\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\sup_{r\in[0,T]}|\boldsymbol{\Delta}\theta_{r}^{k}(\overset{\leftarrow}{\boldsymbol{y}})|.$
It follows that
$\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s-\eta)\leq\Theta^{k}(\overset{\leftarrow}{\boldsymbol{x}},s)\leq\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s+\eta),$
and thus
$|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{x}},s)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|\leq|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s+\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|+|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s-\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|$
It thus remains to prove that $\eta(\delta)\rightarrow 0$ as $\delta\rightarrow 0$ and
$\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s+\tilde{\eta})-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|\rightarrow 0,$ (6.24)
as $\tilde{\eta}\rightarrow 0$. First, denote
$D_{R}=\bigcup_{i=1}^{k+1}B_{R}(u_{i}).$
We get for $\bar{t}\in(0,T)$ and $R>0$ that
$\begin{split}&\eta(\delta)\leq\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,1\rangle dt\\ &\leq 2\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt\\ &\qquad+\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{\bar{t}}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,\mathbbm{1}_{D_{R}}\rangle dt\\ &\qquad+2k\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,\infty),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt.\end{split}$ (6.25)
The second term on the right hand side of (6.25) converges to zero as $\delta\rightarrow 0$ by Proposition 3.6 and for $\tilde{\epsilon}>0$ arbitrary we can choose $\bar{t}\in(0,T)$ small enough and $R>0$ large enough such that
$\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt+k\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,0),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt\quad<\tilde{\epsilon}$.
Thus, it remains to prove (6.24). To this end it is enough to prove
[…]
converges to zero as $\tilde{\eta}\rightarrow 0$. Note that by using Lemma 3.3 and (6.23) we have
$\begin{split}\int_{r_{1}}^{r_{2}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt\geq\int_{r_{1}}^{r_{2}}\langle V_{t+\bar{t}}(\phi),1\rangle dt\geq|r_{2}-r_{1}|\langle V_{T+\bar{t}}(\phi),1\rangle,\end{split}$
where $\bar{t}$ is given by (6.22) and $T$ is given by (6.20). Thus, $r_{\Delta}(\tilde{\eta})\rightarrow 0$ as $\tilde{\eta}\rightarrow 0$. This finishes the proof. ∎
In the following lemma we establish the weak convergence of the random vectors
$\boldsymbol{U}_{k}^{n,m}=(x_{0},U_{n,1}^{m},\ldots U_{n,k}^{m})$ (6.26)
$\boldsymbol{Z}_{k}^{n,m}=(m,Z_{n,1}^{m},\ldots Z_{n,k}^{m})$ (6.27)
$\boldsymbol{T}_{k}^{n,m}=(T_{n,1}^{m},\ldots T_{n,k}^{m})$ (6.28)
$\boldsymbol{M}_{k}^{n,m}=(M_{n,1}^{m},\ldots M_{n,k}^{m}),$ (6.29)
for $k\geq 1$, which will allow us to prove the convergence towards the dual process.
Lemma 6.8.
For every $k\in\mathbb{N}$ it follows that
$(\boldsymbol{U}_{k}^{n,m},\boldsymbol{Z}_{k}^{n,m},\boldsymbol{T}_{k}^{n,m},\boldsymbol{M}_{k}^{n,m})\Rightarrow(\boldsymbol{U}_{k}^{\infty,m},\boldsymbol{Z}_{k}^{\infty,m},\boldsymbol{T}_{k}^{\infty,m},\boldsymbol{M}_{k}^{\infty,m}),\quad\forall m\in\mathbb{R}_{+}$ (6.30)
as $n\rightarrow\infty$ and
$(\boldsymbol{U}_{k}^{m,m},\boldsymbol{Z}_{k}^{m,m},\boldsymbol{T}_{k}^{m,m},\boldsymbol{M}_{k}^{m,m})\Rightarrow(\boldsymbol{U}_{k}^{\infty,\infty},\boldsymbol{Z}_{k}^{\infty,\infty},\boldsymbol{T}_{k}^{\infty,\infty},\boldsymbol{M}_{k}^{\infty,\infty}),$ (6.31)
as $m\rightarrow\infty$ along integers and
$(\boldsymbol{U}_{k}^{\infty,m},\boldsymbol{Z}_{k}^{\infty,m},\boldsymbol{T}_{k}^{\infty,m},\boldsymbol{M}_{k}^{\infty,m})\Rightarrow(\boldsymbol{U}_{k}^{\infty,\infty},\boldsymbol{Z}_{k}^{\infty,\infty},\boldsymbol{T}_{k}^{\infty,\infty},\boldsymbol{M}_{k}^{\infty,\infty})$ (6.32)
as $m\rightarrow\infty$ along integers, with respect to the canonical metric of $\mathbb{R}^{k+1}\times\overline{\mathbb{R}}_{+}^{k+1}\times\mathbb{R}_{+}^{k}\times\{1,2\}^{k}$.
Proof.
We only prove (6.31). The other convergences follow in a similar manner. Furthermore, the first components of $\boldsymbol{U}_{k}^{m,m}$ and $\boldsymbol{Z}_{k}^{m,m}$ are convergent and are deterministic.
We denote
$(P_{m,k}^{1},P_{m,k}^{2},P_{m,k}^{3},P_{m,k}^{4})=(Z_{m,k}^{m},M_{m,k}^{m},U_{m,k}^{m},T_{m,k}^{m})$
and
$D_{1}\times D_{2}\times D_{3}\times D_{4}=\overline{\mathbb{R}}_{+}\times\{1,2\}\times\mathbb{R}\times\mathbb{R}_{+}.$
It is enough to prove
$\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}\prod_{\ell=1}^{k}\prod_{i=1}^{4}f^{i}_{\ell}(P_{m,\ell}^{i})\Big{]}=\mathbb{E}\Big{[}\prod_{\ell=1}^{k}\prod_{i=1}^{4}f_{\ell}^{i}(P_{\infty,\ell}^{i})\Big{]}.$ (6.33)
for all continuous, bounded functions $f_{k}^{i}$ on $D_{i}$ for $i\in\{1,\ldots,4\}$ and $k\geq 1$. We prove the claim via induction. First set $k=1$ and let $f^{i}$ for $i\in\{1,\ldots 4\}$ be arbitrary, continuous, bounded function on $D_{i}$ for $i\in\{1,\ldots k\}$. We need to prove
$\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}\prod_{i=1}^{4}f^{i}(P_{m,1}^{i})\Big{]}=\mathbb{E}\Big{[}\prod_{i=1}^{4}f^{i}(P_{\infty,1}^{i})\Big{]}.$ (6.34)
We get
$\mathbb{E}\Big{[}\prod_{i=1}^{4}f^{i}(P_{m,1}^{i})\Big{]}=\mathbb{E}\Big{[}\mathbb{E}\Big{[}\prod_{i=1}^{4}f^{i}(P_{m,1}^{i})|T_{m,1}^{m}\Big{]}\Big{]}=\mathbb{E}\Big{[}\mathbb{E}\Big{[}\prod_{i=1}^{3}f^{i}(P_{m,1}^{i})|T_{m,1}^{m}\Big{]}f^{4}(T_{m,1}^{m})\Big{]}.$ (6.35)
Now, by using that $(Z_{m,1}^{m},M_{m,1}^{m})$ and $U_{m,1}^{m}$ are conditionally independent given $T_{m,1}^{m}$, we arrive at
$\mathbb{E}\Big{[}\prod_{i=1}^{3}f^{i}(P_{m,1}^{i})|T_{m,1}^{m}\Big{]}=\mathbb{E}[f^{1}(Z_{m,1}^{m})f^{2}(M_{m,1}^{m})|T_{m,1}^{m}]\mathbb{E}[f^{3}(U_{m,1}^{m})|T_{m,1}^{m}].$ (6.36)
Furthermore, note that $(Z_{m,1}^{m},M_{m,1}^{m})$ is independent of $T_{m,1}^{m}$ and thus
$\begin{split}&\mathbb{E}[f^{1}(Z_{m,1}^{m})f^{2}(M_{m,1}^{m})|T_{m,1}^{m}]=\mathbb{E}[f^{1}(Z_{m,1}^{m})f^{2}(M_{m,1}^{m})]\\ &\mathbb{E}[\mathbb{E}[f^{1}(Z_{m,1}^{m})|M_{m,1}^{1}]f^{2}(M_{m,1}^{m})]]\\ &=\mathbb{E}[f^{1}(\bar{Z}_{m}(1))]\mathbb{E}[\mathbbm{1}_{M_{m,1}^{m}=1}f^{2}(1)]+\mathbb{E}[f^{1}(\bar{Z}_{m}(2))]\mathbb{E}[\mathbbm{1}_{M_{m,1}^{m}=2}f^{2}(2)],\end{split}$ (6.37)
where $\bar{Z}_{m}(j)$ is a random variable distributed according to
$\mathbb{P}(\bar{Z}_{m}(j)\geq b)=\frac{\nu^{j}_{m}([b,\infty))}{\nu^{j}_{m}((0,\infty))},$
for $j\in\{1,2\}$ and $b>0$. Furthermore, we get
$\mathbb{E}[f^{3}(U_{m,1}^{m})|T_{m,1}^{m}]=\frac{1}{\int_{\mathbb{R}}\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})(x)dx}\int_{\mathbb{R}}f^{3}(x)\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})(x)dx.$ (6.38)
Together, (6.34) , (6.36), (6.37) and (6.38) yield
$\begin{split}&\mathbb{E}\Big{[}\prod_{i=1}^{4}f^{i}(P_{m,1}^{i})\Big{]}\\ &=\sum_{j=1}^{2}\mathbb{E}[f^{1}(\bar{Z}_{m}(j))]\mathbb{E}[\mathbbm{1}_{M_{m,1}^{m}=j}f^{2}(j)]\\ &\qquad\times\mathbb{E}\Big{[}\frac{1}{\int_{\mathbb{R}}\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})(x)dx}\int_{\mathbb{R}}f^{3}(x)\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})(x)dxf^{4}(T_{m,1}^{m})\Big{]}.\end{split}$ (6.39)
Clearly $\bar{Z}_{m}(j)$ and $M_{m,1}^{m}$ converge in distribution to $\bar{Z}_{\infty}(j)$ and $M_{\infty,1}^{\infty}$, respectively. Hereby, $\bar{Z}_{\infty}(j)$ is distributed according to
$\mathbb{P}(\bar{Z}_{\infty}(j)\in(0,b])=\frac{\nu^{j}((0,b])+d_{j}\mathbbm{1}_{b=\infty}}{\nu^{j}((0,\infty))+d_{j}},\quad j=1,2,b\in\bar{\mathbb{R}}^{+}.$
Thus, it only remains to show the convergence of
$\frac{1}{\int_{\mathbb{R}}\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})dx}\int_{\mathbb{R}}f^{3}(x)\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})dx.$
To this end note, that $S_{m,1}^{m}$ has the same distribution for all $m$ and that $Z_{m,0}=m$ converges weakly to $Z_{\infty,0}=\infty$ with respect to the metric $d$ as $m\rightarrow\infty$. Along similar lines as in the proof of Lemma 6.7 we get that the map
$(z,s)\mapsto\inf\{r\geq 0\colon\int_{0}^{r}\langle V_{t}(\phi+z\delta_{x}),1\rangle dt>s\}$
is continuous and bounded with respect to the canonical metric on $\overline{\mathbb{R}}_{+}\times\mathbb{R}_{+}$. Note that
$T^{m}_{m,1}=\inf\{r\geq 0\colon\int_{0}^{r}\langle V_{t}(\phi+m\delta_{x}),1\rangle dt>S_{m,1}^{m}/\nu^{m}((0,\infty])\}.$
Thus, we get that $T_{m,1}^{m}$ converges weakly to $T_{\infty,1}^{\infty}$. Furthermore, the map
$(t,z)\mapsto\frac{1}{\int_{\mathbb{R}}\mathfrak{V}^{1}(x_{0},z,t)(x)dx}\int_{\mathbb{R}}f^{3}(x)\mathfrak{V}^{1}(x_{0},z,t)(x)dx$
is continuous and bounded for all $(t,z)\in(0,\infty)\times\overline{\mathbb{R}}_{+}$ . Since $\mathbb{P}(T_{m,1}^{m}=0)=0$ the convergence
$\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}\frac{1}{\int_{\mathbb{R}}\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})(x)dx}\int_{\mathbb{R}}f^{3}(x)\mathfrak{V}^{1}(x_{0},m,T_{m,1}^{m})(x)dxf^{4}(T_{m,1}^{m})\Big{]}\\ &=\mathbb{E}\Big{[}\frac{1}{\int_{\mathbb{R}}\mathfrak{V}^{1}(x_{0},\infty,T_{\infty,1}^{\infty})(x)dx}\int_{\mathbb{R}}f^{3}(x)\mathfrak{V}^{1}(x_{0},\infty,T_{\infty,1}^{\infty})(x)dxf^{4}(T_{\infty,1}^{\infty})\Big{]},\end{split}$
follows. Along similar lines to the above we get that
$\begin{split}&\mathbb{E}\Big{[}\prod_{i=1}^{4}f^{i}(P_{\infty,1}^{i})\Big{]}\\ &=\sum_{j=1}^{2}\mathbb{E}[f^{1}(\bar{Z}_{\infty}(1))]\mathbb{E}[\mathbbm{1}_{M_{\infty,1}^{\infty}=j}f^{2}(M_{\infty,1}^{\infty})]\\ &\qquad\times\mathbb{E}\Big{[}\frac{1}{\int_{\mathbb{R}}\mathfrak{V}^{1}(x_{0},\infty,T_{\infty,1}^{\infty})(x)dx}\int_{\mathbb{R}}f^{3}(x)\mathfrak{V}^{1}(x_{0},\infty,T_{\infty,1}^{\infty})(x)dxf^{4}(T_{\infty,1}^{\infty})\Big{]}\end{split}$
and thus (6.34) holds.
Now we assume that (6.33) holds for some $k\geq 1$ and prove that it also holds for $k+1$. To this end let $f^{i}_{\ell}$ be bounded, continuous functions on $D^{i}$, for $i\in\{1,\ldots 4\}$ and $\ell\in\{1,\ldots k+1\}$. Note that $(Z_{m,k+1}^{m},M_{m,k+1}^{m})$ is independent of $(\boldsymbol{U}_{k+1}^{m,m},\boldsymbol{Z}_{k}^{m,m},\boldsymbol{T}_{k+1}^{m,m},\boldsymbol{M}_{k}^{m,m})$ . Thus it is enough to show
$\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}\prod_{\ell=1}^{k}\prod_{i=1}^{4}f^{i}_{\ell}(P_{m,\ell}^{i})f_{k+1}^{3}(U_{k+1,m}^{m})f_{k+1}^{4}(T_{m,k+1}^{m})\Big{]}\\ &\qquad=\mathbb{E}\Big{[}\prod_{\ell=1}^{k}\prod_{i=1}^{4}f_{\ell}^{i}(P_{\infty,\ell}^{i})f_{k+1}^{3}(U_{\infty,k+1}^{\infty})f_{k+1}^{4}(T_{\infty,k+1}^{\infty})\Big{]}.\end{split}$
First, consider
$\mathbb{E}\Big{[}\prod_{\ell=1}^{k}\prod_{i=1}^{3}f^{i}_{\ell}(P_{m,\ell}^{i})f_{k+1}^{4}(T_{m,k+1}^{m})\Big{]},$
and note that by (6.19) we have
$f_{k+1}^{4}(T_{m,k+1}^{m})=f_{k+1}^{4}(\Theta^{k}(\boldsymbol{U}_{k}^{m,m},\boldsymbol{T}_{k}^{m,m},\boldsymbol{Z}_{k}^{m,m},S_{m,k+1}^{m}/\nu_{m}((0,\infty]))).$
Thus, by Lemma 6.7 and the continuous mapping theorem, it follows that
$(\boldsymbol{U}_{k}^{m,m},\boldsymbol{T}_{k+1}^{m,m},\boldsymbol{Z}_{k+1}^{m,m},\boldsymbol{M}_{k+1}^{m,m})\Rightarrow(\boldsymbol{U}_{k}^{\infty,\infty},\boldsymbol{T}_{k+1}^{\infty,\infty},\boldsymbol{Z}_{k+1}^{\infty,\infty},\boldsymbol{M}_{k+1}^{\infty,\infty}),$
as $m\rightarrow\infty$. Again, since
$(\boldsymbol{u},\boldsymbol{t},\boldsymbol{z},t)\mapsto\frac{1}{\int_{\mathbb{R}}V_{t}(\mathfrak{V}^{k}(\boldsymbol{u},\boldsymbol{z},\boldsymbol{t}))(x)dx}\int_{\mathbb{R}}f^{3}_{k+1}(x)V_{t}(\mathfrak{V}^{k}(\boldsymbol{u},\boldsymbol{z},\boldsymbol{t}))(x)dx$
is bounded and continuous on $\mathbb{R}^{k}\times(0,\infty)^{k}\times\bar{\mathbb{R}}_{+}^{k}\times(0,\infty)$, the convergence
$(\boldsymbol{U}_{k+1}^{m,m},\boldsymbol{T}_{k+1}^{m,m},\boldsymbol{Z}_{k+1}^{m,m},\boldsymbol{M}_{k+1}^{m,m})\Rightarrow(\boldsymbol{U}_{k+1}^{\infty,\infty},\boldsymbol{T}_{k+1}^{\infty,\infty},\boldsymbol{Z}_{k+1}^{\infty,\infty},\boldsymbol{M}_{k+1}^{\infty,\infty})$
follows. This finishes the proof. ∎
Lemma 6.9.
Let $a\in\mathbb{R}$ and let $X_{0}$ be a bounded and continuous function. Then there exists $T_{0}>0$, independent of $X_{0}$, such that the following convergences hold:
  1. (i)
    $\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{m,m}}\exp\Big{(}-\langle Y^{m,m}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big{)}\Big{]},\end{split}$
  2. (ii)
    $\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{n,m}}\exp\Big{(}-\langle Y^{n,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{n,m}_{s},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,m}}\exp\Big{(}-\langle Y^{\infty,m}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,m}_{s},1\rangle ds\Big{)}\Big{]},\end{split}$
  3. (iii)
    $\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,m}}\exp\Big{(}-\langle Y^{\infty,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{\infty,m}_{s},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big{)}\Big{]}\end{split}$
    for all $t\in[0,T_{0}]$.
Proof.
We only show (i), since (ii) and (iii) follow along similar lines. Let $T_{0}>0$ be such that (6.8) is satisfied with $\gamma=2a$. Fix arbitrary $t>0$ and let $\epsilon>0$. By (6.11), there exists a sequence of random variables $(\bar{R}_{i})_{i\geq 1}$ such that
$\mathbb{P}(\bar{T}_{m,i}^{m}\leq t)\leq\mathbb{P}(\bar{R}_{i}\leq t),\quad\forall m,i\geq 1,\quad t\geq 0.$
Furthermore, by using Lemma 6.1, it is not hard to see that $\lim_{i\rightarrow\infty}\bar{R}_{i}=\infty$ almost surely. Thus, there exists $k\in\mathbb{N}$ such that
$\sup_{m\in\mathbb{N}}\mathbb{P}(\bar{T}_{m,k}^{m}\leq t)<\epsilon/2$ (6.40)
and
$\mathbb{P}(\bar{T}_{\infty,k}^{\infty}\leq t)<\epsilon/2.$
Thus we arrive at
$\begin{split}&\Big{|}\mathbb{E}\Big{[}(-1)^{J_{t}^{m,m}}\exp\Big{(}-\langle Y^{m,m}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\Big{)}\Big{]}\\ &\quad-\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big{)}\Big{]}\Big{|}\\ &\leq C\epsilon+\sum_{i=1}^{k}\Big{|}\mathbb{E}[(-1)^{J_{t}^{m,m}}\exp(-\langle Y^{m,m}_{t}(x),X_{0})-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\rangle)\mathbbm{1}_{t\in[\bar{T}_{m,i-1}^{m},\bar{T}_{m,i}^{m})}]\\ &\qquad-\mathbb{E}[(-1)^{J_{t}^{\infty,\infty}}\exp(-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds)\mathbbm{1}_{t\in[\bar{T}_{\infty,i-1}^{\infty},\bar{T}_{\infty,i}^{\infty})}]\Big{|},\end{split}$
where we used that
$\sup_{m}\mathbb{E}\Big{[}\exp\Big{(}-2a\int_{0}^{t}\langle Y_{s}^{m,m},1\rangle ds\Big{)}\Big{]}<\infty.$
Note that by (6.18) we get for $i\in\{2,\ldots,k\}$ that
$Y_{t}^{m,m}(x)=\mathfrak{V}^{i}(\boldsymbol{U}_{i-1}^{m,m},\boldsymbol{Z}_{i-1}^{m,m},(\boldsymbol{T}_{i-1}^{m,m},t-\bar{T}_{m,i-1}^{m}))(x),\quad x\in\mathbb{R}$ (6.41)
and for $t\in(\bar{T}_{m,i-1}^{m},\bar{T}_{m,i}^{m})$ and
$Y_{t}^{\infty,\infty}(x)=\mathfrak{V}^{i}(\boldsymbol{U}_{i-1}^{\infty,\infty},\boldsymbol{Z}_{i-1}^{\infty,\infty},(\boldsymbol{T}_{i-1}^{\infty,\infty},t-\bar{T}_{\infty,i-1}^{\infty}))(x),\quad x\in\mathbb{R}$ (6.42)
and for $t\in(\bar{T}_{\infty,i-1}^{\infty},\bar{T}_{\infty,i}^{\infty})$. Due to Lemma 6.1 and Lemma 6.3, we get that the number of jumps in $[0,t]$ is almost surely bounded and thus
$\sup_{m\in\mathbb{N}}J_{t}^{m,m}<\infty$
almost surely. Furthermore, note that
$J_{t}^{m,m}=\sum_{\ell=1}^{i-1}\mathbbm{1}_{M_{m,\ell}^{m}=2},\quad t\in[\bar{T}_{i-1,m}^{m},\bar{T}_{i,m}^{m}).$ (6.43)
Now, it is sufficient to show that
$\begin{split}&\Big{|}\mathbb{E}\Big{[}(-1)^{J_{t}^{m,m}}\exp\Big{(}-\langle Y^{m,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\rangle\Big{)}\mathbbm{1}_{t\in[\bar{T}_{m,i-1}^{m},\bar{T}_{m,i}^{m})}\Big{]}\\ &\qquad-\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big{)}\mathbbm{1}_{t\in[\bar{T}_{\infty,i-1}^{\infty},\bar{T}_{\infty,i}^{\infty})}\Big{]}\Big{|}\end{split}$ (6.44)
converges to zero as $m\rightarrow\infty$. Let $\epsilon>0$ be arbitrary and choose $K>0$ sufficiently large. We get that
$\begin{split}&\sup_{m}\mathbb{E}\Big{[}\Big{|}\exp\Big{(}-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\rangle\Big{)}-\exp\Big{(}-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\rangle\Big{)}\wedge K\Big{|}\Big{]}\\ &=\sup_{m}\mathbb{E}\Big{[}\exp\Big{(}-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\rangle\Big{)}\mathbbm{1}_{\exp\big{(}-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\rangle\big{)}>K}\Big{]}\\ &\leq\frac{C}{K}<\epsilon/2,\end{split}$ (6.45)
for $K>0$ sufficiently large by using Lemma 6.5. Similarly, we can ensure that
$\mathbb{E}\Big{[}\Big{|}\exp\Big{(}-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\rangle\Big{)}-\exp\Big{(}-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\rangle\Big{)}\wedge K\Big{|}\Big{]}<\epsilon/2.$ (6.46)
Due to Lemma 6.8 and (6.41), (6.42) and (6.43) we also get that
$\begin{split}&\Big{|}\mathbb{E}\Big{[}(-1)^{J_{t}^{m,m}}\Big{(}\exp\Big{(}-\langle Y^{m,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{m,m}_{s},1\rangle ds\rangle\Big{)}\wedge K\Big{)}\mathbbm{1}_{t\in[\bar{T}_{m,i-1}^{m},\bar{T}_{m,i}^{m})}\Big{]}\\ &\quad-\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\Big{(}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big{)}\wedge K\Big{)}\mathbbm{1}_{t\in[\bar{T}_{\infty,i-1}^{\infty},\bar{T}_{\infty,i}^{\infty})}\Big{]}\Big{|},\end{split}$
converges to zero as $m\rightarrow\infty$. This together with (6.45) and (6.46) demonstrates the convergence of (6.44) to zero as $m\rightarrow\infty$ and thus finishes the proof. ∎
Lemma 6.10.
Let $T_{0}>0$ be such that (6.8) holds for $\gamma=2a$ and let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with initial condition $X_{0}$ satisfying (1.5). Furthermore, let $(g_{n})_{n\geq 1}$ be a sequence of measurable bounded functions on $\mathbb{R}$, that converge boundedly pointwise to a bounded, measurable function $g$ on $\mathbb{R}$. Then it follows for all $T\leq T_{0}$ that
$\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n,m},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{2}\Big{]}=0,$
for all $m\in\mathbb{R}_{+}$.
Proof.
Fix $m\in\mathbb{N}$. To lighten notation, we suppress the subscript $m$ for the rest of the proof. Note that by Lemma 6.2 and (6.13) we have
$\sup_{n}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{4}\Big{]}\leq C\sup_{n}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},1\rangle ds\Big{)}^{4}\Big{]}\leq C\mathbb{E}[I_{T}^{4}]<\infty$
Furthermore, by Lemma 6.8 it follows that
$\begin{split}&(U_{n,1},U_{n,2},\ldots,T_{n,1},T_{n_{2}},\ldots,Z_{n,1},Z_{n,2},\ldots)\\ &\qquad\Rightarrow(U_{\infty,1},U_{\infty,2},\ldots,T_{\infty,1},T_{\infty,2},\ldots,Z_{\infty,1},Z_{\infty,2},\ldots)\end{split}$ (6.47)
and thus by the Skorohod representation theorem there exists a probability space such that the convergence (6.47) holds almost surely. We now show that almost surely
$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds=0.$
First, we can use a similar argument to the proof of Lemma 6.8 to see that for almost all $\omega$ there exists $M(\omega)\in\mathbb{N}$ such that
$\sup_{n}\bar{T}(\omega)_{n,M(\omega)+1}>T\quad,M(\omega)<\infty.$
Furthermore, note that there exists $D(\omega)>0$ such that almost surely
$\sup_{n}|U_{n,k}(\omega)|<D(\omega),\quad D(\omega)<\infty,$
for all $k\leq M$, since $(U_{n,k})_{n\geq 1}$ converges almost surely for each fixed $k$. Let $\epsilon>0$ be arbitrary. Then, by using Lemma 3.2 and Proposition 3.6 we get the following. There exists $\bar{D}(\omega)>D(\omega)$ such that
$\int_{0}^{T}\int_{B(0,\bar{D})^{c}}Y^{n}_{s-}(x)dxds\leq M\int_{0}^{T}\int_{B(0,\bar{D})^{c}}\big{(}\mathcal{W}(D,0)\big{)}(x)dxds\leq\epsilon,$
with $\bar{D}(\omega)<\infty$ almost surely. Furthermore, there exists $\delta_{1}(\omega)>0$ such that
$M\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds<\epsilon.$
Finally, once again by Proposition 3.7, there exists $K(\omega)>0$ such that
$M\sup_{s\in(\delta_{1},T),x\in B(0,\bar{D}+D)}\mathcal{W}_{s}(0,0)(x)\leq K<\infty,$
almost surely. Thus, we get
$\begin{split}&\int_{0}^{T}\langle Y_{s-}^{n},|h_{n}(X_{T-s})-h(X_{T-s})|\rangle ds\\ &\leq C\epsilon+\sum_{k=1}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\epsilon+\int_{0}^{T}\int_{B(0,\bar{D})}\big{(}V_{s}(\phi+m\delta_{x_{0}})\big{)}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\quad+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds,\end{split}$
where $C=\sup_{x\in\mathbb{R}}|g(x)|+\sup_{n\geq 1}\sup_{x\in\mathbb{R}}|g_{n}(x)|<\infty$ is independent of $n$. Furthermore, for $\delta_{1}>0$ sufficiently small we get
$\begin{split}&\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq CM\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n+\delta_{1}}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|h_{n}(X_{T-s}(x))-h(X_{T-s}(x))|dxds\\ &\leq C\epsilon+M\big{(}\sup_{s\in(\delta_{1},T),x\in B(0,\bar{D}+D)}\mathcal{W}_{s}(0,0)(x)\big{)}\int_{0}^{T}\int_{B(0,\bar{D})}|h_{n}(X_{T-s}(x))-h(X_{T-s}(x))|dxds\\ &\leq C\epsilon+K\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\end{split}$
By the dominated convergence theorem, we get that
$\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\rightarrow 0,$
as $n\rightarrow\infty$. This demonstrates that
$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\leq 2C\epsilon$
almost surely. Since $\epsilon>0$ was arbitrary, this proves that
$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds=0,$
almost surely. Thus, by uniform integrability, we get
$\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{2}]\\ &=\mathbb{E}\Big{[}\Big{(}\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{2}\Big{]}=0.\end{split}$
6.3 Proof of Approximate Duality
In this section, we assume that there exists a weak solution $(X_{t})_{t\geq 0}$ to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) is satisfied and prove an (approximate) duality relation. When $h$ is continuous, the existence of a solution was already resolved in [25, Theorem 1.1]. The existence of solutions when $h$ is discontinuous is proven in Section 7 and relies on the duality relation for the continuous case. In order to establish the duality relation, we consider approximation of the process $Y$ defined in (6.5). To this end, we denote by $Y^{n}$ the process constructed in (6.5) with $\nu$ replaced by
$\nu_{n}=\nu^{1}_{|[0,n]}+\nu^{2}_{|[0,n]}+d_{1}\delta_{n}+d_{2}\delta_{n}$
and $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}$. Accordingly, we define
$h_{n}(x)=\int_{0}^{n}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+\int_{0}^{n}(1+\operatorname{e}^{-\lambda x})\nu^{2}(d\lambda)+d_{1}(1-\operatorname{e}^{-nx})+d_{2}(1+\operatorname{e}^{-nx})+a,\quad x\in\mathbb{R},$ (6.48)
for $n\in\mathbb{N}$, which converges pointwise to $h$ given by
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+\int_{0}^{\infty}(1+\operatorname{e}^{-\lambda x})\nu^{2}(d\lambda)+d_{1}(\mathbbm{1}_{x>0})+d_{2}(1+\mathbbm{1}_{x=0})+a,\quad x\in\mathbb{R}.$ (6.49)
Lemma 6.11.
[…]
$\sup_{s\in[0,t]}\sup_{x\in\mathbb{R}}\mathbb{E}\big{[}X_{s}(x)\big{]}<\infty,$
for all $t>0$.
Proof.
Note that
$\begin{split}&\mathbb{E}[X_{t}(x)]=\mathbb{E}[\big{(}S_{t}X_{0}\big{)}(x)]+\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}\mathbb{E}[h(X_{s}(y))]dyds\\ &\leq\mathbb{E}[\big{(}S_{t}X_{0}\big{)}(x)]+C\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}(x-y)dyds,\end{split}$
for all $x\in\mathbb{R}$ and $t\geq 0$. This implies the result by standard estimates. ∎
Lemma 6.12.
[…]
$\begin{split}&\mathbb{E}_{X}\Big{[}\exp(-\langle V_{t}(\mu)),X_{T-t}\rangle+v_{t})\Big{]}=\mathbb{E}_{X}\Big{[}\exp(-\langle V_{0}(\mu),X_{T}\rangle)\Big{]}\\ &\qquad+\mathbb{E}_{X}\Big{[}\int_{0}^{t}\exp(-\langle V_{s}(\mu),X_{T-s}\rangle+v_{s})\big{(}\langle V_{s}(\mu),h(X_{T-s})\rangle+v_{s}^{\prime}\big{)}ds\Big{]},\quad t\in[0,T].\end{split}$
for all finite measures $\mu$.
Proof.
The proof follows along similar lines as the proof of [21, Lemma 2.6], but we include it for the sake of completeness. Let $(\mu_{k})_{k\geq 1}$ be a sequence in $\mathcal{C}_{c}^{\infty}$ that converges weakly to $\mu$. Then it follows that $V_{\cdot}(\mu_{k})\in\mathcal{C}^{1,2}(\mathbb{R}_{+}\times\mathbb{R})\cap L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$, where $\psi\in L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$ if
$\int_{0}^{t}\int_{\mathbb{R}}\psi_{s}^{2}(x)dxds<\infty,\quad\forall t>0.$ (6.50)
Now, let $T\geq 0$ and $t\in[0,T]$. Thus, using (6.50) we get by Itô’s formula
$\begin{split}&\mathbb{E}_{X}\Big{[}\exp(-\langle V_{0}(\mu_{k})),X_{T}\rangle)\Big{]}=\mathbb{E}_{X}\Big{[}\exp(-\langle V_{T-t}(\mu_{k}),X_{t}\rangle+v_{T-t})\Big{]}\\ &\qquad-\mathbb{E}_{X}\Big{[}\int_{t}^{T}\exp\big{(}-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s}\big{)}\big{(}\langle V_{T-s}(\mu_{k}),h(X_{s})\rangle+v_{T-s}^{\prime}\big{)}ds\Big{]},\end{split}$ (6.51)
since $V_{t}(\mu_{k})$ solves (3.2) and
$\mathbb{E}_{X}\Big{[}\int_{t}^{T}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-\cdots}(\mu_{k}))\Big{]}=0.$ (6.52)
Here, $M$ is an orthogonal martingale measure such that for all $\psi\in\mathcal{C}(\mathbb{R}_{+}\times\mathbb{R})\cap L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$, $M_{t}(\psi)$ is a square integrable martingale with quadratic variation given by
$\langle M(\psi)\rangle_{t}=\int_{0}^{t}\langle\psi_{s}^{2},X_{s}\rangle ds,\quad t\geq 0.$
Hereby, (6.52) follows due to the fact that
$M_{s}(V_{T-s}(\mu_{k}))$
is a square-integrable martingale since, by Lemma 6.11 and (6.50), we have for $t\in[0,T]$ that
$\begin{split}&\mathbb{E}[\langle M_{\cdot}(V_{T-\cdot})\rangle_{t}]=\mathbb{E}\Big{[}\int_{0}^{t}\int_{\mathbb{R}}V_{T-s}(\mu_{k})(x)^{2}X_{s}(x)dxds\Big{]}\\ &\leq\int_{0}^{t}\int_{\mathbb{R}}V_{T-s}(\mu_{k})(x)^{2}dxds\sup_{s\in[0,T]}\sup_{x\in\mathbb{R}}\mathbb{E}[X_{s}(x)]<\infty.\end{split}$
This and the inequality
$\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})\leq\exp(v_{T-s}),\quad\forall s\leq t,$ (6.53)
shows that $\int_{0}^{t}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-s}(\mu_{k}))$ is a martingale, which implies (6.52). It remains to show the convergence as $k\rightarrow\infty$. Using [22, Lemma 2.1], the dominated convergence theorem and setting $\tilde{t}=T-t$ in (6.51) yields
$\begin{split}&\mathbb{E}_{X}\Big{[}\exp(-\langle V_{t}(\mu)),X_{T-t}\rangle+v_{t})\Big{]}=\mathbb{E}_{X}\Big{[}\exp(-\langle V_{0}(\mu),X_{T}\rangle)\Big{]}\\ &\qquad+\mathbb{E}_{X}\Big{[}\int_{0}^{t}\exp(-\langle V_{s}(\mu),X_{T-s}\rangle+v_{s})\big{(}\langle V_{s}(\mu),h(X_{T-s})\rangle+v_{s}^{\prime}\big{)}ds\Big{]},\quad\forall t\in[0,T].\end{split}$
Fix arbitrary $m\in\mathbb{R}_{+}$, $x_{0}\in\mathbb{R}$ and $\mu\in\mathcal{M}_{F}^{+}$. Until the end of this section let $Y^{n}$ denote the process $Y^{n,m}$ constructed in Section 6.2 with $Y_{0}=\mu+m\delta_{x_{0}}$. We also suppress the upper index $m$ from $T_{n,k}^{m}$, $\bar{T}_{n,k}^{m}$, $U_{n,k}^{m}$, $Z_{n,k}^{m}$ and $M_{n,k}^{m}$. Furthermore, let the processes $Y^{n}$ be independent of $X$ for all $n\geq 1$. To this end define the point process
$p^{(n)}\colon D_{p^{(n)}}\rightarrow\{1,2\}\times\mathbb{R}_{+}\times\mathbb{R},$
with
$D_{p^{(n)}}=\{\bar{T}_{n,1},\bar{T}_{n,2},\ldots\},$
and
$p^{(n)}(\bar{T}_{n,k})=(M_{n,k},Z_{n,k},U_{n,k}),\forall k\geq 1.$
The associated counting measure is given by
$\mathfrak{P}^{(n)}(t,B)=\#\{s\in D_{p^{(n)}}|s\leq t,p^{(n)}(s)\in B)\},$
for measurable $B\subseteq\{1,2\}\times\mathbb{R}_{+}\times\mathbb{R}$.
Lemma 6.13.
The compensator of $\mathfrak{P}^{(n)}$ is given by
$\hat{\mathfrak{P}}^{(n)}(t,B_{1}\times B_{2}\times B_{3})=\int_{0}^{t}\sum_{i=1}^{2}\mathbbm{1}_{i\in B_{1}}\int_{B_{2}}\int_{B_{3}}Y^{n}_{s-}(x)dx\nu_{n}^{i}(d\lambda)ds,$ (6.54)
for measurable set $B_{1}\subseteq\{1,2\}$, $B_{2}\subseteq\mathbb{R}_{+}$ and $B_{3}\subseteq\mathbb{R}$.
Proof.
We have that
$\mathbb{E}[\mathfrak{P}^{(n)}(t,B)]\leq\mathbb{E}[\mathfrak{N}^{(n)}(t,\{1,2\}\times\mathbb{R}_{+}\times\mathbb{R})]\leq\nu_{n}((0,\infty))\mathbb{E}\Big{[}\int_{0}^{t}\int_{\mathbb{R}}Y_{s-}^{n}(x)dxds\Big{]}\leq C(T),$
where the last inequality follows by Proposition 6.4. That means, that $\mathfrak{P}^{(n)}$ is a monotone increasing, adapted, integrable process and thus, by the Doob-Meyer theorem the compensator exists. A calculation shows that it takes the form (6.54). ∎
Lemma 6.14.
Let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) holds and let $a$ be defined as in (A). Let $Y^{n}$ be as above, independent of $X$. Then there exists a time $T_{0}>0$ only depending on $\langle\mu,1\rangle$ and $a$ such that for all $T\leq T_{0}$ we have
$\begin{split}&\mathbb{E}\Big{[}(-1)^{J_{t}^{n}}\exp\Big{(}-\langle Y_{t}^{n},X_{T-t}\rangle-a\int_{0}^{t}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}=\mathbb{E}\Big{[}\exp(-\langle Y_{0},X_{T}\rangle)\Big{]}\\ &+\mathbb{E}\Big{[}\int_{0}^{t}(-1)^{J_{s-}^{n}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle-a\int_{0}^{s}\langle Y_{r}^{n},1\rangle dr\Big{)}\Big{(}\langle Y_{s-}^{n},h(X_{T-s})\rangle-\langle Y_{s-}^{n},h_{n}(X_{T-s})\rangle\Big{)}ds\Big{]}.\end{split}$
Proof.
First, choose $T_{0}>0$ small enough such that Lemma 6.5 is satisfied with $\gamma=2a$ and let $T\leq T_{0}$. Furthermore, set $v_{t}=-a\int_{0}^{t}\langle Y^{n}_{s},1\rangle ds$ for $t\geq 0$. By Lemma 6.12 we have
$\begin{split}&\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle-a\int_{0}^{\bar{T}_{n,k+1-}}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k}}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k}},X_{T-\bar{T}_{n,k}}\rangle-a\int_{0}^{\bar{T}_{n,k}}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}\\ &\quad+\mathbb{E}_{X}\Big{[}\int_{\bar{T}_{n,k}}^{\bar{T}_{n,k+1}}(-1)^{J^{n}_{s-}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle-a\int_{0}^{s}\langle Y_{r}^{n},1\rangle dr\Big{)}\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big{]},\end{split}$ (6.55)
since $t\mapsto X_{t}$ is continuous and $J_{\bar{T}_{n,k+1}-}=J_{s}$ for all $s\in(\bar{T}_{n,k},\bar{T}_{n,k+1})$. Let $\boldsymbol{\Delta}Y_{s}=Y_{s}-Y_{s-}$ for all $s\geq 0$, then we get
$\begin{split}&\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle-a\int_{0}^{\bar{T}_{n,k+1-}}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle+v_{\bar{T}_{n,k+1}}\Big{)}\Big{]}\\ &+\exp\big{(}v_{\bar{T}_{n,k+1}}\big{)}\mathbb{E}_{X}\Big{[}(-1)^{J_{\bar{T}_{n,k+1}}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\\ &\hskip 199.16928pt-(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\Big{]}\\ &=\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k}}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k}},X_{T-\bar{T}_{n,k}}\rangle+v_{\bar{T}_{n,k}}\Big{)}\Big{]}\\ &+\mathbb{E}_{X}\Big{[}\int_{\bar{T}_{n,k}}^{\bar{T}_{n,k+1}}(-1)^{J^{n}_{s-}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle+v_{s}\Big{)}\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big{]}\\ &-\exp\big{(}v_{\bar{T}_{n,k+1}}\big{)}\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\\ &\times\Big{(}(1-\exp(-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}))\mathbbm{1}_{M_{n,k}=1}+\big{(}1+\exp\big{(}-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\big{)}\mathbbm{1}_{M_{n,k}=2}\Big{)}\Big{]},\end{split}$
where in the second equality we used (6.55) and the definition of $J^{n}$. This leads to
$\begin{split}&\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{t}}\exp\big{(}-\langle Y_{t}^{n},X_{T-t}\rangle\big{)}\exp\Big{(}-a\int_{0}^{t}\langle Y^{n}_{s},1\rangle ds\Big{)}\Big{]}=\mathbb{E}\Big{[}\exp\big{(}-\langle Y_{0},X_{T}\rangle\big{)}\Big{]}\\ &+\mathbb{E}_{X}\Big{[}\int_{0}^{t}(-1)^{J^{n}_{s-}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle-a\int_{0}^{s}\langle Y_{r}^{n},1\rangle dr\Big{)}\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big{]}\\ &-\int_{0}^{t}\int_{\mathbb{R}}\int_{\mathbb{R}_{+}}\int_{\mathbb{R}}(-1)^{J^{n}_{s-}}\mathbb{E}_{X}\Big{[}\exp\Big{(}-\langle Y^{n}_{s-},X_{t-s}\rangle-a\int_{0}^{s}\langle Y_{r}^{n},1\rangle dr\Big{)}\\ &\qquad\times\Big{(}\big{(}1-\operatorname{e}^{-\lambda X_{T-s}(x)}\big{)}\mathbbm{1}_{r=1}+(1+\operatorname{e}^{-\lambda X_{T-s}(x)})\mathbbm{1}_{r=2}\Big{)}\Big{]}\mathfrak{P}^{(n)}(dr,d\lambda,dx,ds).\end{split}$ (6.56)
By Lemma 6.13 the compensator of $\mathfrak{P}^{(n)}$ is given by (6.54). Furthermore, by Lemma 6.5 it follows that
$\mathbb{E}\Big{[}\exp\Big{(}-a\int_{0}^{T_{0}}\langle Y_{r}^{n},1\rangle dr\Big{)}\Big{]}<\infty.$
Now take an expectation $\mathbb{E}_{Y}$ on both sides of (6.56) and use the definition of $h_{n}$ and $h$, given by (6.48) and (6.49), respectively, to get the desired result. ∎
Corollary 6.15.
[…]
$\lim_{n\rightarrow\infty}\mathbb{E}[(-1)^{J^{n}_{T}}\exp\Big{(}-\langle Y_{T}^{n},X_{0}\rangle-a\int_{0}^{T}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}=\mathbb{E}[\exp(-\langle Y_{0},X_{T}\rangle)],$
for all $T\leq T_{0}$.
Proof.
By Lemma 6.14 we have
$\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J^{n}_{T}}\exp\Big{(}-\langle Y_{T}^{n},X_{0}\rangle-a\int_{0}^{T}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}[\exp\big{(}-\langle Y_{0},X_{T}\rangle\big{)}]\\ &+\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\int_{0}^{t}(-1)^{J^{n}_{s-}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle-a\int_{0}^{s}\langle Y_{r-}^{n},1\rangle dr\Big{)}\Big{(}\langle Y_{s-}^{n},h(X_{T-s})\rangle-\langle Y_{s-}^{n},h_{n}(X_{T-s})\rangle\Big{)}ds\Big{]}.\end{split}$
Furthermore, by Lemma 6.10 and Lemma 6.5 we have
$\begin{split}&\lim_{n\rightarrow\infty}\Big{|}\mathbb{E}\Big{[}\int_{0}^{t}(-1)^{J^{n}_{s-}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle-a\int_{0}^{s}\langle Y_{r-}^{n},1\rangle dr\Big{)}\Big{(}\langle Y_{s-}^{n},h(X_{T-s})\rangle-\langle Y_{s-}^{n},h_{n}(X_{T-s})\rangle\Big{)}ds\Big{]}\Big{|}\\ &\leq\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\exp\Big{(}-a\int_{0}^{t}\langle Y_{s}^{n},1\rangle ds\Big{)}\int_{0}^{t}\langle Y_{s-}^{n},|h_{n}(X_{T-s})-h(X_{T-s})|\rangle ds\Big{]}\\ &\leq\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\exp\Big{(}-2a\int_{0}^{t}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}^{1/2}\mathbb{E}\Big{[}\Big{(}\int_{0}^{t}\langle Y_{s-}^{n},|h_{n}(X_{T-s})-h(X_{T-s})|\rangle ds\Big{)}^{2}\Big{]}^{1/2}=0.\end{split}$
This concludes the proof. ∎
Finally, we can prove Proposition 4.3.
Proof of Proposition 4.3.
The assertion follows immediately from Corollary 6.15 and Lemma 6.9. ∎
7 Proof of Weak Existence
In this section we prove the weak existence part of Theorem 1.1. Again, our arguments were inspired by [3].
Note, that we only need to proof weak existence in the case where $h$ is discontinuous due to [25, Theorem 1.1]. That means we need to consider the case where $b_{0}\neq b_{1}\neq 0$. For simplicity, we thus only consider the case where $\nu^{1}=\nu^{2}\equiv 0$. The prove relies on considering a sequence $(X^{n})_{n\geq 1}$ of solutions to (1.1), where $h$ is approximated by continuous functions $h_{n}$ given by
$h_{n}(x)=d_{1}(1-\operatorname{e}^{-\lambda nx})+d_{2}(1+\operatorname{e}^{-\lambda nx})+a,$
where $d_{1},d_{2}$ and $a$ are given according to (A). The sequence of functions $(h_{n})_{n\geq 1}$ converges boundedly pointwise to $h$. As mentioned above, we get the existence of a weak solution $X^{n}$ by applying [25, Theorem 1.1] for each $n\in N$. When $h$ is replaced by $h_{n}$, we also get the weak uniqueness of solutions by applying the results from Section 8. Furthermore, we can apply the duality relation from Corollary 6.15 for $X^{n}$, which is conditional on existence.
In a first step, we prove tightness of the sequence $(X^{n})_{n\geq 1}$, which can be achieved in a similar manner to the proves from [25]. The main difficulty is to show that the limit point $\bar{X}$ is a solution to (1.1). Similar to the approach from [3], we will use the duality relation from Section 6 for this purpose.
7.1 Tightness
Recall, that $X_{0}\in\mathcal{P}_{p}(\mathcal{C}_{tem})$ for $p>2$. We can apply standard methods to derive the following lemma.
Lemma 7.1.
We have for all $\lambda>0$, $0\leq q\leq p/2$ and $T>0$ that
$\sup_{n\geq 1}\sup_{0\leq t\leq T}\int_{\mathbb{R}}\operatorname{e}^{-\lambda|x|}\mathbb{E}[|X_{t}^{n}(x)|^{2q}]dx<\infty$ (7.1)
and there exist $\gamma>2$ and $C_{\lambda}>0$, independent of $m$, such that
$\mathbb{E}[|X_{t}^{n}(x)-X_{t^{\prime}}^{n}(x^{\prime})|^{2q}]\leq C_{\lambda}(|t-t^{\prime}|^{\gamma}+|x-x^{\prime}|^{\gamma})\operatorname{e}^{\lambda|x|},$ (7.2)
for $t\geq 0$ and $x,x^{\prime}\in\mathbb{R}$ such that $|x-x^{\prime}|\leq 1$.
Proof.
The proof follows from standard estimates as in the proof of [25, Theorem 2.6]. ∎
Lemma 7.1 demonstrates by [25, Lemma 6.3] that the sequence $(X^{n})_{n\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables is tight. Thus, by Prohorov’s theorem and Skorohod’s representation theorem, there exists a sub-sequence $(X^{n_{k}})_{k\geq 1}$ and another sequence $(\tilde{X}^{n_{k}})_{k\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables such that the following holds.
  1. 7.1.
    The sequence $(\tilde{X}^{n_{k}})_{k\geq 1}$ is defined on a common probability space;
  2. 7.2.
    For all $k\geq 1$ we have that $X^{n_{k}}\overset{law}{=}\tilde{X}^{n_{k}}$;
  3. 7.3.
    The sequence $(\tilde{X}^{n_{k}})_{k\geq 1}$ converges almost surely to a limit $X$ with respect to the topology of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$.
In the following we shall denote $\tilde{X}^{n_{k}}$ by $X^{n}$, with a slight abuse of notation, for the sake of readability.
7.2 Convergence of $h_{n}(X^{n})$
Lemma 7.2.
There exists $T_{0}>0$ such that for all $x\in\mathbb{R}$ and $0\leq t\leq T_{0}$ it holds that
$\exp(-nX^{n}_{t}(x))\rightarrow\mathbbm{1}_{X_{t}(x)=0}$
in $L^{q}$ for $q\in[1,\infty)$ as $m\rightarrow\infty$.
Proof.
The proof is similar to the proof of [3, Lemma 5.3]. First, we prove the $L^{1}$ convergence. Let $T_{0}>0$ be such that Proposition 4.3 holds for $X^{n}$, for all $n\geq 0$. Note that it is possible to choose $T_{0}>0$ uniformly in $n$, since it only depends on $\sup_{n}\nu_{n}^{1}(0,\infty)+\nu_{n}^{2}(0,\infty)$. For simplicity denote $X_{t}(x)$ by $Z$ and $X^{n}_{t}(x)$ by $Z_{n}$. Below we use the fact that almost surely $Z_{n},Z\geq 0$. Thus we get
$\begin{split}&\mathbb{E}[|\exp(-nZ_{n})-\mathbbm{1}_{Z=0}|]\\ &=\mathbb{E}[|\exp(-nZ_{n})-\mathbbm{1}_{Z=0}|\mathbbm{1}_{Z=0}]+\mathbb{E}[|\exp(-nZ_{n})-\mathbbm{1}_{Z=0}|\mathbbm{1}_{Z>0}]\\ &=\mathbb{E}[|\exp(-nZ_{n})-1|\mathbbm{1}_{Z=0}]+\mathbb{E}[|\exp(-nZ_{n})|\mathbbm{1}_{Z>0}]\\ &=\mathbb{E}[\mathbbm{1}_{Z=0}]-\mathbb{E}[\exp(-nZ_{n})]+2\mathbb{E}[|\exp(-nZ_{n})|\mathbbm{1}_{Z>0}].\end{split}$ (7.3)
Since $Z_{n}\rightarrow Z$ almost surely, we get that $2\mathbb{E}[|\exp(-nZ_{n})|\mathbbm{1}_{Z>0}]\rightarrow 0$. By Lemma 6.9 and by applying Corollary 6.15 with $X^{n}$, we furthermore get that
$\begin{split}&\mathbb{E}[\mathbbm{1}_{Z=0}]=\lim_{m\rightarrow\infty}\mathbb{E}[\exp(-mX_{t}(x))]\\ &=\lim_{m\rightarrow\infty}\lim_{n\rightarrow\infty}\mathbb{E}[\exp(-mX^{n}_{t}(x))]\\ &=\lim_{m\rightarrow\infty}\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{n,m}}\exp\Big{(}-\langle Y^{n,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{n,m},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,\infty},1\rangle ds\Big{)}\Big{]},\end{split}$
and
$\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}[\exp(-nZ_{n})]=\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{n,n}}\exp\Big{(}-\langle Y^{n,n}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{n,n},1\rangle ds\Big{)}\Big{]}\\ &\quad=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,\infty},1\rangle ds\Big{)}\Big{]}.\end{split}$
Thus it follows that
$\mathbb{E}[\mathbbm{1}_{Z=0}]-\mathbb{E}[\exp(-nZ_{n})]\rightarrow 0,$
as $n\rightarrow\infty$. By the above and (7.3) the $L^{1}$ convergence of $\exp(-nX_{t}^{n}(x))$ to $\mathbbm{1}_{X_{t}(x)=0}$ follows. From this it is standard to derive the convergence in $L^{q}$ for $q\in[1,\infty)$. ∎
Proof of Weak Existence.
First, we prove weak existence of solutions up to time $T_{0}>0$ such that Proposition 4.3 holds. Note that $T_{0}$ does not depend on the initial condition. Thus, we can extend our solution provided that $\mathbb{P}\circ X_{T_{0}}^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and satisfies (1.5). This, however, readily follows along similar lines to Lemma 7.1 and by Lemma 6.11. Fix arbitrary $\phi\in\mathcal{C}_{c}^{\infty}(\mathbb{R})$. Recall that according to our conventions $X^{n}\rightarrow X$ in $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$ almost surely as $n\rightarrow\infty$. Since $X^{n}$ solves (1.1) with drift $h_{n}$ for $n\geq 1$, we have that almost surely,
$\begin{split}&\int_{\mathbb{R}}\phi(x)X^{n}_{t}(x)dx-\int_{\mathbb{R}}\phi(x)X_{0}(x)dx\\ &\qquad=\frac{1}{2}\int_{\mathbb{R}}\int_{0}^{t}\phi^{\prime\prime}(x)X_{s}^{n}(x)dsdx+\int_{\mathbb{R}}\int_{0}^{t}\phi(x)h_{n}(X_{s}^{n}(x))dsdx+M_{t}^{n}(\phi),\end{split}$ (7.4)
where $(M_{t}^{n}(\phi))_{t\geq 0}$ is an $L^{2}$-martingale with quadratic variation
$\langle M_{\cdot}^{n}(\phi)\rangle_{t}=\int_{\mathbb{R}}\int_{0}^{t}\phi(x)^{2}X_{s}^{n}(x)dsdx.$
From Lemma 7.2 we know that the second term on the right hand side converges to
$\int_{\mathbb{R}}\int_{0}^{t}\phi(x)h(X_{s}(x))dsdx$
in $L^{2}$. Furthermore, due to Lemma 7.1 and since $p>2$ we have that $(X^{n}_{\cdot}-X\cdot)^{2}$ on the interval $[0,t]$ and restricted to the support of $\phi$ is uniformly integrable. For this reason,
$\frac{1}{2}\int_{\mathbb{R}}\int_{0}^{t}\phi^{\prime\prime}(x)X_{s}^{n}(x)dsdx\rightarrow\frac{1}{2}\int_{\mathbb{R}}\int_{0}^{t}\phi^{\prime\prime}(x)X_{s}(x)dsdx,$
in $L^{2}$ as $n\rightarrow\infty$ and
$\int_{\mathbb{R}}\phi(x)X^{n}_{t}(x)dx\rightarrow\int_{\mathbb{R}}\phi(x)X_{t}(x)dx$
in $L^{2}$ as $n\rightarrow\infty$. This implies that $M_{t}^{n}(\phi)$ converges in $L^{2}$ to a random variable $M_{t}(\phi)$ for all $t\in[0,T_{0}]$. Now, it is standard to deduce that $(M_{t}(\phi))_{t\in[0,T_{0}]}$ is a $L^{2}$-martingale with quadratic variation
$\langle M_{\cdot}(\phi)\rangle_{t}=\int_{\mathbb{R}}\int_{0}^{t}\phi(x)^{2}X_{s}(x)dsdx,\quad t\in[0,T_{0}].$
Thus, all the terms in (7.4) converge and we get
$\begin{split}&\int_{\mathbb{R}}\phi(x)X_{t}(x)dx-\int_{\mathbb{R}}\phi(x)X_{0}(x)dx\\ &\qquad=\frac{1}{2}\int_{\mathbb{R}}\int_{0}^{t}\phi^{\prime\prime}(x)X_{s}(x)dsdx+\int_{\mathbb{R}}\int_{0}^{t}\phi(x)h(X_{s}(x))dsdx+M_{t}(\phi),\end{split}$
for $t\in[0,T_{0}]$. ∎
8 Proof of Weak Uniqueness
Lemma 8.1.
Let $X^{1},X^{2}$ be two solutions to (1.1) with $X_{0}^{1},X_{0}^{2}\sim\mu\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) is satisfies. Then, it follows that
$X_{t}^{1}\overset{d}{=}X_{t}^{2}$
for all $t\in[0,T_{0}]$, where $T_{0}$ only depends on $a$.
Proof.
First, we choose $T_{0}$ as in Proposition 4.3 under the assumption that $\langle Y_{0},1\rangle=1$. Now, fix $t\in[0,T_{0}]$ and let $f$ be a fixed but arbitrary, non-negative simple function. Then it follows for $g=\frac{1}{\|f\|_{L^{1}}}f$ that
$\mathbb{E}[\exp(-\lambda\langle g,X_{t}^{1}\rangle)]=\mathbb{E}[\exp(-\lambda\langle g,X_{t}^{2}\rangle)]\quad\forall\lambda\in[0,1],$ (8.1)
by Proposition 4.3. Since the map $\lambda\mapsto\mathbb{E}[\exp(-\lambda\langle g,X_{t}^{i}\rangle)]$ is analytic on $(0,\infty)$ for $i\in\{1,2\}$, (8.1) extends to all $\lambda\geq 0$. In particular, we have that
$\mathbb{E}[\exp(-\lambda\langle f,X_{t}^{1}\rangle)]=\mathbb{E}[\exp(-\lambda\langle f,X_{t}^{2}\rangle)]\quad\forall\lambda\geq 0.$
Since $f$ was arbitrary it follows by [13, Corollary 2.3] that $X^{1}_{t}\overset{d}{=}X_{t}^{2}$. ∎
Theorem 8.2.
The solution to (1.1) is weakly unique in $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$.
Proof.
It is enough to prove weak uniqueness on the interval $0,T_{0}]$ and that $\mathbb{P}\circ(X_{T_{0}})^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and that $X_{T_{0}}$ satisfies (1.5), since $T_{0}$ does not depend on the initial value $X_{0}$. This follows along similar lines as Lemma 7.1 and by Lemma 6.11. Thus, it remains to prove uniqueness on $0,T_{0}]$. However, by using Lemma 8.1, this follows along the same lines as the proof of [11, Theorem 4.2]. It is only necessary to show that t $\mathbb{P}\circ(X_{t})^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and that $X_{t}$ satisfies (1.5) for all $t\in[0,T_{0}]$. However, this once again follows along similar lines as Lemma 7.1 and from Lemma 6.11. ∎
9 Proof of Theorem 4.5
Before we proof Theorem 4.5, we derive the following helpful lemma.
Lemma 9.1.
Let $s>0$, $a>0$ and $x,y\in(-\infty,a)$ such that $x\leq y$. Then
$\int_{a}^{\infty}\mathcal{W}_{s}(x,0)(z)dz\leq\int_{a}^{\infty}\mathcal{W}_{s}(y,0)(z)dz.$
Proof.
By Proposition 3.7 we have
$\begin{split}&\int_{a}^{\infty}\mathcal{W}_{s}(x,0)(z)dz=s^{-1}\int_{a}^{\infty}f(s^{-1/2}|z-x|)dz\\ &=s^{-1}\int_{a-x+y}^{\infty}f(s^{-1/2}|z^{\prime}-y|)dz^{\prime}\leq\int_{a}^{\infty}\mathcal{W}_{s}(y,0)(z^{\prime})dz^{\prime},\end{split}$
which proves the assertion. ∎
Proof of Theorem 4.5.
Let $T_{0}>0$ be small enough such that
$\mathbb{E}[\operatorname{e}^{|I_{T_{0}}(Y_{0})||}]<\infty$ (9.1)
with $Y_{0}=\infty\delta_{0}$. Such $T_{0}>0$ exists due to Lemma 6.2. For $s=0$ the assertion follows by assumption. Thus, let $s\in(0,T_{0}]$. Using Proposition 4.3 and Remark 4.4 it follows that
$\begin{split}&\mathbb{E}\Big{[}\int_{\mathbb{R}}\mathbbm{1}_{X_{s}(x)>0}dx\Big{]}=\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\int_{\mathbb{R}}\big{(}1-\operatorname{e}^{-nX_{s}(x)}\big{)}dx\Big{]}\\ &=\lim_{n\rightarrow\infty}\int_{\mathbb{R}}\big{(}1-\mathbb{E}[\operatorname{e}^{-nX_{s}(x)}]\big{)}dx\\ &=\lim_{n\rightarrow\infty}\int_{\mathbb{R}}\big{(}1-\mathbb{E}[\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})\rangle}]\big{)}dx.\end{split}$ (9.2)
Note, that it is enough to prove
$\mathbb{E}\Big{[}\int_{\mathbb{R}}\mathbbm{1}_{X_{s}(x)>0}dx\Big{]}<\infty.$
We denote
$R_{x}=(|x|-R)/2,$
where $R$ is such that $\operatorname{supp}(X_{0})\subseteq B(0,R)$ and we also assume $R>2$. Furthermore for $|x|>2R$, denote
$\tau_{R_{x},n}=\inf\{k\geq 0\colon|U_{n,k}^{x}|\leq R_{x}+R\},$
where $U_{n,k}^{x}$ is the location of the $k$-th jump of the process $Y(n\delta_{x})$ and $U_{n,0}^{x}=x$. Moreover, define
$K_{t}^{n}=\max\{k\in\mathbb{N}\colon\bar{T}_{n,k}^{x}\leq t\},\quad t\geq 0,$
where $\bar{T}_{n,k}^{x}$ is time of the $k$-th jump of the process $Y(n\delta_{x})$. That is, $K^{n}_{t}$ is the number of jumps of $Y(n\delta)$ on the time interval $(0,t]$. We furthermore denote $T_{n,k}^{x}=\bar{T}_{n,k}^{x}-\bar{T}_{n,k-1}^{x}$, for $k\geq 1$. We obtain the following decomposition
$\begin{split}&\mathbb{E}[(1-\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})})]\\ &=\mathbb{E}[(1-\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})})\mathbbm{1}_{K_{s}^{n}>|x|^{1/4}}]\\ &\quad+\mathbb{E}[(1-\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})})\mathbbm{1}_{K_{s}^{n}\leq|x|^{1/4}}\mathbbm{1}_{\tau_{R_{x},n}\leq K_{s}^{n}}]\\ &\quad+\mathbb{E}[(1-\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})})\mathbbm{1}_{K_{s}^{n}\leq|x|^{1/4}}\mathbbm{1}_{\tau_{R_{x},n}>K_{s}^{n}}]\\ &=I_{1}^{n}(s,x)+I_{2}^{n}(s,x)+I_{3}^{n}(s,x)\end{split}$ (9.3)
With out loss of generality we assume that $x>2R$. First, we use (9.1) and the fact that $|K_{s}^{n}|$ is stochastically dominated by $|I_{s}(Y_{0})|$ in order to estimate $I_{1}^{n}(s,x)$. Hereby, the stochastic domination follows from (6.11). We arrive at
$I_{1}^{n}(s,x)\leq\mathbb{P}(|I_{s}(Y_{0})|^{8}>x^{2})\leq\frac{\mathbb{E}[|I_{s}(Y_{0})|^{8}]}{x^{2}}\wedge 1.$ (9.4)
Note, that (9.1) implies that $\mathbb{E}[|I_{s}(Y_{0})|^{8}]<\infty$. Next, we estimate $I_{3}^{n}(s,x)$ for $x>2R$. By using Lemma 3.3 and Proposition 3.6 we arrive at
$I_{3}^{n}(s,x)\leq\mathbb{E}[\langle X_{0},Y_{s}(n\delta_{x})\rangle\mathbbm{1}_{K_{s}^{n}\leq|x|^{1/4}}\mathbbm{1}_{\tau_{R_{x},n}>K^{n}_{s}}]\leq C\sum_{k=0}^{\lceil|x|^{1/4}\rceil}\langle\mathbbm{1}_{(-\infty,R]},\mathcal{W}_{s-\bar{T}_{n,k}^{x}}(R_{x}+R,0)\rangle,$ (9.5)
where we used Lemma 9.1 and the fact that $U_{n,k}^{x}>R+R_{x}$ for all $0\leq k\leq|x|^{1/4}+1$ and that $\operatorname{supp}(X_{0})\subseteq B(0,R)$. Note, that by Proposition 3.7 we have that $\mathcal{W}_{t}(0,0)(y)=t^{-1}f(t^{-1/2}|y|)$, $t>0,y\in\mathbb{R}$ such that
$\sup_{y\in[s^{-1/2},\infty)}\frac{f(y)}{\operatorname{e}^{-1/2y^{2}}y}\leq c(s).$
Therefore, for all $r>0$ we have
$\begin{split}&\langle\mathbbm{1}_{(-\infty,R]},\mathcal{W}_{r}(R_{x}+R,0)\rangle\leq C\int_{-\infty}^{R}\mathcal{W}_{r}(R_{x}+R,0)(y)dy\\ &=\int_{-\infty}^{-R_{x}}\mathcal{W}_{r}(0,0)(y)dy\leq C\int^{\infty}_{R_{x}}r^{-3/2}|y|\operatorname{e}^{-\frac{y^{2}}{2r}}dy=Cr^{-1/2}\operatorname{e}^{-\frac{R_{x}^{2}}{2r}},\end{split}$
where we used that $R_{x}>1$ for $x>2R$. This leads to
$\langle\mathbbm{1}_{(-\infty,R]},\mathcal{W}_{r}(R_{x}+R,0)\rangle\leq Cr^{-1/2}(\operatorname{e}^{-\frac{(R_{x})^{2}}{2r}}).$ (9.6)
For fixed $y\in\mathbb{R}$ the map
$r\mapsto r^{-1/2}\operatorname{e}^{-\frac{y^{2}}{2r}},\quad r>0,$
is maximized for $r=y^{2}$ and is monotonously increasing for $r\leq y^{2}$. Recall that $r=s-\bar{T}_{n,k}^{x}\leq s$ and thus we get the estimate
$I_{3}^{n}(s,x)\leq C(|x|^{1/4}+1)\min(s,R_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{(R_{x})^{2}}{2\min(R_{x}^{2},s)}}.$ (9.7)
Finally, we estimate $I_{2}^{n}(s,x)$. To this end, we brake up the event into several parts that we can handle more easily.
$\begin{split}&I_{2}^{n}(s,x)\leq\mathbb{P}(\tau_{R_{x},n}\leq K^{n}_{s},K^{n}_{s}\leq|x|^{1/4})\\ &\leq\mathbb{P}(U_{k,x}^{n}\leq R_{x}+R\text{ for some }k\in\{1,\ldots,K^{n}_{s}\},K^{n}_{s}\leq|x|^{1/4})\\ &\leq\sum_{m=1}^{\lceil|x|^{1/4}\rceil}\sum_{k=1}^{m}\mathbb{P}(U_{n,j}^{x}>R_{x}+R,j\in\{1,\ldots k-1\},U_{n,k}^{x}\leq R_{x}+R,K_{s}^{n}=m).\end{split}$ (9.8)
Now, we denote $\eta_{x}=R_{x}/(|x|^{1/4}+1)$. Then, for $1\leq k\leq m$ we get
$\begin{split}&\mathbb{P}(U_{n,j}^{x}>R_{x}+R,j\in\{1,\ldots k-1\},U_{n,k}^{x}\leq R_{x}+R,K_{s}^{n}=m)\\ &\leq\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\\ &\qquad+\mathbb{P}(U_{n,j}^{x}\leq x-j\eta_{x}\text{ for some }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\end{split}$ (9.9)
We first estimate the first term on the right hand side of (9.9). For $1\leq k\leq m\leq|x|^{1/4}+1$ we get
$\begin{split}&\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots,k-1\},K_{s}^{n}=m)\\ &\leq\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots,k-1\},\bar{T}_{n,k}^{x}\leq s)\\ &=\mathbb{E}[\mathbbm{1}_{\bar{T}_{k,x}^{n}\leq s}\mathbbm{1}_{U_{n,j}^{x}>x-j\eta_{x},j\in\{1,\ldots,k-1\}}\mathbbm{1}_{U_{n,\ell}^{x}\leq R+R_{x}}]\\ &=\mathbb{E}[\mathbbm{1}_{\bar{T}_{n,k}^{x}\leq s}\mathbbm{1}_{U_{n,j}^{x}>x-j\eta_{x},j\in\{1,\ldots,k-1\}}\mathbb{E}[\mathbbm{1}_{U_{n,\ell}^{x}\leq R+R_{x}}|\mathcal{G}_{n,x,k}]],\end{split}$
where $\mathcal{G}_{n,x,k}$ is the $\sigma$-algebra generated by $(U_{n,1}^{x},\ldots U_{n,k-1}^{x},T_{n,1}^{x},\ldots T_{n,k}^{x})$ . For $y\in\mathbb{R}$ we get using Lemma 3.3, Proposition 3.6 and the construction of the process $Y(n\delta_{x})$ that
$\begin{split}&\mathbbm{1}_{\bar{T}_{n,k}^{x}\leq s}\mathbb{E}[\mathbbm{1}_{U_{n,k}^{x}\leq y}|\mathcal{G}_{n,x,k}]\\ &=\mathbbm{1}_{\bar{T}_{n,k}^{x}\leq s}\frac{\int_{-\infty}^{y}Y_{\bar{T}_{n,k}^{x}-}(n\delta_{x})(y^{\prime})dy^{\prime}}{\int_{\mathbb{R}}Y_{\bar{T}_{n,k}^{x}-}(n\delta_{x})(y^{\prime})dy^{\prime}}\leq C\sum_{j=0}^{k-1}\int_{-\infty}^{y}\mathcal{W}_{\bar{T}_{n,k}^{x}-\bar{T}_{n,j}^{x}}(U_{n,j}^{x},0)(y^{\prime})dy^{\prime}),\end{split}$ (9.10)
where we also used that
$\inf_{n\geq 1}\int_{\mathbb{R}}Y_{\bar{T}_{n,k}^{x}-}(n\delta_{x})(y^{\prime})dy^{\prime}\geq\int_{\mathbb{R}}V_{s}(1\delta_{x})(y^{\prime})dy^{\prime}>c>0,$ (9.11)
for $\bar{T}_{n,k}^{x}\leq s$. Note that we again used Lemma 3.3, Proposition 3.6 and the construction of $Y(n\delta_{x})$ as well as (6.15) in order to derive (9.11). Using (9.10) and Lemma 9.1, we thus get that
$\begin{split}&\mathbb{E}[\mathbbm{1}_{U_{n,k}^{x}\leq R+R_{x}}|\mathcal{G}_{n,x,k}]\leq C\sum_{j=0}^{k-1}\int_{-\infty}^{R+R_{x}}\mathcal{W}_{\bar{T}_{n,k}^{x}-\bar{T}_{n,j}^{x}}(U_{n,j}^{x},0)(y^{\prime})dy^{\prime}\\ &\leq C\sum_{j=0}^{k-1}\int_{-\infty}^{R+R_{x}}\mathcal{W}_{\bar{T}_{n,k}^{x}-\bar{T}_{j,x}^{n}}(x-(k-1)\eta_{x},0)(y^{\prime})dy^{\prime}\\ &\leq Ck\min(s,\eta_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{(\eta_{x})^{2}}{2\min(s,\eta_{x}^{2})}},\end{split}$
on the event $\{\bar{T}_{n,k}^{x}\leq s,U_{n,j}^{x}>x-j\eta_{x}j\in\{1,\ldots k-1\}\}$. Hereby, we again used Proposition 3.7 and the fact that $x-(k-1)\eta_{x}-R-R_{x}\geq\eta_{x}>c>0$ for all $k\leq\lceil|x|^{1/4}\rceil$ to estimate the integral in the last inequality in the same was as in (9.6) and (9.7). Thus we arrive at
$\begin{split}&\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots,k-1\},K_{s}^{n}=m)\\ &\leq Ck\min(s,\eta_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{(\eta_{x})^{2}}{2\min(s,\eta_{x}^{2})}}.\end{split}$ (9.12)
Now, we estimate the second term on the right hand side of (9.9). First we can estimate
$\begin{split}&\mathbb{P}(U_{n,j}^{x}\leq x-j\eta_{x}\text{ for some }j\in\{1,\ldots,k-1\},K_{s}^{n}=m)\\ &\leq\sum_{\ell=1}^{k-1}\mathbb{P}(U_{n,j}^{x}>x-j\eta_{x},j\in\{1,\ldots,\ell-1\},U_{n,\ell}^{x}\leq x-\ell\eta_{x},\bar{T}_{n,\ell}^{x}\leq s),\end{split}$ (9.13)
for $1\leq k\leq m\leq|x|^{1/4}+1$. Similarly to above we get for $1\leq\ell\leq|x|^{1/4}+1$ that
$\begin{split}&\mathbb{P}(U_{n,j}^{x}>x-j\eta_{x},j\in\{1,\ldots,\ell-1\},U_{n,\ell}^{x}\leq x-\ell\eta_{x},\bar{T}_{n,\ell}^{x}\leq s)\\ &=\mathbb{E}[\mathbbm{1}_{\bar{T}_{n,\ell}^{x}\leq s}\mathbbm{1}_{U_{n,j}^{x}>x-j\eta_{x},j\in\{1,\ldots,\ell-1\}}\mathbbm{1}_{U_{n,\ell}^{x}\leq x-\ell\eta_{x}}]\\ &=\mathbb{E}[\mathbbm{1}_{\bar{T}_{n,\ell}^{x}\leq s}\mathbbm{1}_{U_{n,j}^{x}>x-j\eta_{x},j\in\{1,\ldots,\ell-1\}}\mathbb{E}[\mathbbm{1}_{U_{n,\ell}^{x}\leq x-\ell\eta_{x}}|\mathcal{G}_{n,x,\ell}]]\\ &\leq C\ell\min(s,\eta_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{\eta_{x}^{2}}{\min(s,\eta_{x}^{2})}},\end{split}$
and thus
$\begin{split}&\mathbb{P}(U_{n,j}^{x}\leq x-j\eta_{x}\text{ for some }j\in\{1,\ldots,k-1\},K_{s}^{n}=m)\\ &\leq Ck^{2}\min(s,\eta_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{\eta_{x}^{2}}{\min(s,\eta_{x}^{2})}},\quad 1\leq k\leq m\leq|x|^{1/4}+1.\end{split}$ (9.14)
Together, (9.8), (9.9), (9.12) and (9.14) yield
$I_{2}(s,x)\leq C(|x|^{1/4}+1)^{4}\min(s,\eta_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{(\eta_{x})^{2}}{2\min(s,\eta_{x}^{2})}}.$ (9.15)
Plugging (9.4), (9.5),(9.15) and (9.3) into (9.2) yields
$\begin{split}&\lim_{n\rightarrow\infty}\int_{\mathbb{R}}\big{(}1-\mathbb{E}[\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})\rangle}]\big{)}dx\\ &\leq\lim_{n\rightarrow\infty}\int_{B(0,2R)^{c}}I^{n}_{1}(s,x)+I^{n}_{2}(s,x)+I^{n}_{3}(s,x)dx+4R\\ &\leq C\int_{B(0,2R)^{c}}x^{-2}\wedge 1dx+C\int_{B(0,2R)^{c}}\min\Big{(}1,(|x|^{1/4}+1)\min(s,R_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{(R_{x})^{2}}{2\min(s,R_{x}^{2})}}\Big{)}dx\\ &\quad+C\int_{B(0,2R)^{c}}\min\Big{(}1,(|x|^{1/4}+1)^{4}\min(s,\eta_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{(\eta_{x})^{2}}{2\min(s,\eta_{x}^{2})}}\Big{)}dx+4R.\end{split}$
Note that there exists $R^{\ast}\geq 2R$ such that
$R_{x}^{2}\geq s,\eta_{x}^{2}\geq s,\quad x\geq R^{\ast}.$
Thus, we get
$\begin{split}&\lim_{n\rightarrow\infty}\int_{\mathbb{R}}\big{(}1-\mathbb{E}[\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})\rangle}]\big{)}dx\\ &\leq C\Big{(}R^{\ast}+s^{-1/2}\int_{R^{\ast}}^{\infty}x\operatorname{e}^{-\frac{(x-R)^{2}}{8s|x|^{1/2}}}dx\Big{)},<\infty.\end{split}$
where we used that $\eta_{x}\leq R_{x}$ for $x\geq R^{\ast}$. This proves the assertion. ∎

### v2

@@
 $d_{t}X_{t}=\frac{1}{2}\Delta X_{t}+h(X_{t})+\sqrt{X_{t}}\dot{W},\quad t\geq 0,$
-where $\dot{W}$ is space–time white noise and $h$ is a bounded drift with $h(0)\geq 0$.
-The proof relies on an extension of the duality relation of the super-Brownian motion, which allows us to treat a broad class of admissible drifts, including functions that are non-Lipschitz or discontinuous at zero.
+where $\dot{W}$ denotes space–time white noise and $h$ is a bounded drift with $h(0)\geq 0$.
+The proof relies on an extension of the duality relation for super-Brownian motion, which allows us to treat a broad class of admissible drifts, including functions that are non-Lipschitz or discontinuous at zero.
+In particular, well-posedness is derived for certain drifts that are Hölder continuous at zero with exponent $\alpha\in(0,1)$.
+We also allow discontinuous drifts of the form $h(x)=b_{0}\mathbbm{1}_{x=0}+b_{1}\mathbbm{1}_{x>0},$ where $b_{0}\geq 0$, $b_{1}\in\mathbb{R}$.
+Additionally, if $h(0)=0$ and the initial condition is continuous and compactly supported, we show that the Lebesgue measure of the non-zero set of $X$ is finite.
+The proofs are based on duality.
+We use a log-Laplace equation which is perturbed by jump noise as the equation for the dual process and the jumps of the dual process are allowed to take infinite values.
+We believe that the results for the dual process are also of independent interest.
+Under suitable assumptions on $h$ we also prove survival of $X$ with positive probability, using rescaling and comparison to the KPP-equation with branching noise.
 1 Introduction
 We consider one-dimensional SPDEs of the form
-$d_{t}X_{t}=\frac{1}{2}\Delta X_{t}+h(X_{t})+\sqrt{X_{t}}\dot{W},\quad X_{0}=f,$ (1.1)
+$d_{t}X_{t}(x)=\frac{1}{2}\Delta X_{t}(x)+h(X_{t}(x))+\sqrt{X_{t}(x)}\dot{W}(t,x),\quad X_{0}=f,x\in\mathbb{R},t\geq 0,$ (1.1)
 where $f$ is a non-negative, continuous, bounded function, $\dot{W}$ is space-time white noise and
@@
 $h_{\infty}(x)=b_{0}\mathbbm{1}_{x=0}+b_{1}\mathbbm{1}_{x>0},\quad x\in\mathbb{R},$ (1.4)
-for $b_{0}\geq\langle\nu_{1}-\nu_{2},1\rangle$, $b_{1}\in\mathbb{R}$ and finite measures $\nu_{1},\nu_{2}$ on $(0,\infty)$.
-In this article, we establish weak well-posedness for solutions of (1.1) using duality methods.
-Hereby, we consider solutions in the space $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$, which is the space of continuous functions taking values in
+for $b_{0}\geq\langle\nu^{1}-\nu^{2},1\rangle$, $b_{1}\in\mathbb{R}$ and finite measures $\nu^{1},\nu^{2}$ on $(0,\infty)$.
+In this article, using duality methods, we establish weak well-posedness for solutions of (1.1) for drifts satisfying conditions (1.21.4).
+A solution to (1.1) with the drift $h(x)=cx$, $c\in\mathbb{R}$, describes the density of the classical super-Brownian motion (SBM) with drift in one dimension, which has been extensively studied in the literature (see, for instance, [36, 21, 12, 14, 25, 32]), and for which weak well-posedness is well known.
+Our goal is to establish weak well-posedness for the much broader class of drifts satisfying conditions (1.21.4), beyond the classical case of $h(x)=cx$.
+We will comment further on this below.
+We consider solutions in the space $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$, which is the space of continuous functions taking values in
 $\mathcal{C}_{tem}^{+}=\{f\in\mathcal{C}(\mathbb{R})\colon f\geq 0,|f|_{(\lambda)}<\infty\quad\forall\lambda<0\},$
@@
 $|f|_{(\lambda)}=\sup_{x\in\mathbb{R}}|\operatorname{e}^{\lambda|x|}f(x)|,\quad f\in\mathcal{C}(\mathbb{R}).$
-We also allow random initial values that are distributed according to a measure in $\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ (see Section 2.1). That means, we prove the following theorem.
+We also allow random initial conditions distributed according to a measure in $\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ (see Section 2.1). We now state the main results of the paper.
 Theorem 1.1.
@@
 Then there exists a unique weak solution to (1.1) in $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$ in the sense of Definition 4.1.
-The main tool for the proof of Theorem 1.1 is the derivation of a duality relation (see Proposition 4.3).
-When $h(0)=0$, this relation also allows us to prove properties of the cozero set of the solution, even when $h$ has a discontinuity at zero.
-Our assumptions on the drift function $h$ permit significant irregularities, including Hölder continuity with exponent less than $\frac{1}{2}$ and even discontinuities at zero.
-Several examples are provided in Section 5.
-Irregularities at zero are of particular interest, as this is precisely where the noise coefficient becomes non-Lipschitz.
-As a result, standard techniques such as the Dawson–Girsanov theorem can no longer be directly applied to establish weak uniqueness when the drift is less than $\frac{1}{2}$-Hölder continuous or exhibits discontinuities at zero.
-To the best of our knowledge, when the drift is discontinuous, neither weak uniqueness nor weak existence is known It turns out, that we can establish weak well-posedness of (1.1) for drift terms $h$ that lead to non-unique solutions or non-existence of solutions in the case of the one-dimensional SDE
-$dx_{t}=h(x_{t})dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=y,\quad t\geq 0,$
-where $y\in\mathbb{R}$ and $(B_{t})_{t\geq 0}$ is a standard Brownian motion.
-A similar behavior was also observed in [3] in the case of the Wright-Fisher SPDE.
-Equation (1.1) with $h=0$ is the density of the super-Brownian motion in one dimension, which was extensively studied in the literature (see for instance [26, 15, 8, 10, 18, 23]).
-When the process is subject to some nonzero space dependent immigration, that is in the case where $h$ is continuous and does not depend on $X$ but instead on the spatial component $x$, pathwise non-uniqueness was established in [7].
-When $h$ is continuous, (1.1) also fits in the setting of [25], where the existence of solutions is proven, but no uniqueness result was obtained for non-Lipschitz coefficients.
-Results on weak well-posedness for SPDEs with non-Lipschitz coefficients remain scarce.
-In the setting of Super-Brownian Motion and other measure valued branching processes a key method for establishing weak uniqueness is the duality approach.
+A natural question concerning the properties of the solution is the following: How large can the cozero set of the solution be in the case where there is no upward drift at the zero points of the solution?
+That is, in the case when $h(0)=0$.
+In this setting, we prove that the cozero set of the solution to (1.1) has finite Lebesgue measure.
+Theorem 1.2.
+Suppose that almost surely, $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$. Let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $h$ given by
+$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+b_{1}\mathbbm{1}_{x>0},\quad x\in\mathbb{R},$ (1.6)
+with $b_{1}\in\mathbb{R}$ such that $h(0)\geq 0$. Then there exists $T_{0}>0$ such that for all $t\in[0,T_{0}]$ we have
+$\mathbb{P}(\mathrm{Leb}(\{x\in\mathbb{R}\colon X_{t}(x)>0\})<\infty)=1.$ (1.7)
+Further fundamental questions concern the size of the closure of the cozero set, in particular whether the compact support property holds under the assumptions of Theorem 1.2.
+We conjecture the following.
+Conjecture 1.3.
+Suppose that almost surely, $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$.
+Let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $h$ given by (1.6).
+Then $X_{t}$ is compactly supported with probability one for all $t\geq 0$.
+Besides the cozero set and the support of the solution it is natural to investigate the longtime behavior, in particular the survival probability of the process.
+Theorem 1.4.
+Suppose that $h\neq 0$ is given by
+$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+b_{1}\mathbbm{1}_{x>0},\quad x\in\mathbb{R},$
+with $b_{1}\geq 0$ and let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$ and $f\neq 0$. Then
+$\mathbb{P}(\langle X_{t},1\rangle>0,\quad\forall t>0)>0.$
+The main tool for the proof of Theorems 1.1 and 1.2 is the duality relation derived in Proposition 4.3.
+Our assumptions about the structure of the drift function $h$ may seem peculiar, but they actually allow for significant irregularities of $h$ at zero, precisely the point where the noise coefficient $\sqrt{x}$ becomes non-Lipschitz and degenerate.
+For example, by choosing
+$\nu^{1}=\nu^{2}\equiv 0,\quad b_{0}\neq b_{1},\quad b_{0}\geq 0,$ (1.8)
+one obtains a drift $h(\cdot)$ that is discontinuous at zero.
+More precisely, this choice of parameters $\nu^{i},b_{i},i=1,2$, in (1.8), allows the drift $h$ to take any constant value at points where the solution $X$ of (1.1) is positive, and any constant non-negative value at points where $X$ equals zero.
+Furthermore, with $\nu^{2}\equiv 0$, for any $\alpha\in(0,1)$, by choosing an appropriate measure $\nu^{1}$ and suitable constants $b_{0},b_{1}$, one can obtain $h(x)$ such that it behaves like $cx^{\alpha}$ for small values of $x$, in the sense that
+$\lim_{x\downarrow 0}\frac{h(x)}{x^{\alpha}}=c>0.$
+To the best of our knowledge, weak uniqueness for such irregular drifts at zero has not been previously established.
+In fact, even the existence of a solution to equation (1.1) with a drift discontinuous at zero was not known.
+It is worth noting that standard techniques for handling drifts, such as the Dawson–Girsanov theorem, can no longer be directly applied to establish weak well-posedness when the drift has a Hölder exponent $\alpha<1/2$ at zero or exhibits a discontinuity at zero.
+It is also noteworthy that we establish weak well-posedness of (1.1) for discontinuous drift terms $h$ that lead to non-uniqueness or non-existence of solutions for the one-dimensional counterpart of (1.1) given by
+$dx_{t}=h(x_{t})dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=0,\quad t\geq 0,$ (1.9)
+where $(B_{t})_{t\geq 0}$ is a standard Brownian motion.
+To be more precise, we show that this equation admits non-unique solutions when $h(x)=\mathbbm{1}_{x>0}$ and no solution exists when $h(x)=\mathbbm{1}_{x=0}$.
+Notably, these are exactly the cases for which we prove weak well-posedness of (1.1) in this paper.
+The above examples in SPDE and SDE settings are discussed in more detail in Section 5.
+The question addressed in this paper belongs to a broader area of well-posedness of SPDEs with irregular coefficients. In particular the SPDE
+$d_{t}X_{t}(x)=\frac{1}{2}\Delta X_{t}(x)+h(X_{t}(x))+\sigma(X_{t}(x))\dot{W}(t,x),\quad X_{0}=f,\;x\in\mathbb{R},t\geq 0,$ (1.10)
+where $\dot{W}$ is a space-time white noise, $h$ is a drift and $\sigma$ is a noise coefficient, has been extensively studied.
+Many results on well-posedness for this equation mainly dealt with the question:
+How irregular can the drift be and what are conditions on the noise coefficient so that well-posedness for this equation holds?
+This problem is also known under the name of regularization by noise, as it has been shown for some irregular $h$ that the equation without noise is ill-posed and when the noise is added, the equation becomes well-posed.
+When the noise is additive ($\sigma\equiv 1$), strong well-posedness was derived for some irregular function-valued drifts in [17], [18], and more recently even for some measure-valued and distribution-valued drifts in [1].
+Later the result from [1] was extended in [11] to SPDEs driven by mutiplicative noises with $\sigma$ bounded away from zero (non-degenerate) and sufficiently regular.
+Weak well-posedness was established in [9] for (1.10) with additive noise for more irregular distributions than in [1].
+The so-called path-by-path uniqueness for (1.10) with additive noise and with bounded drifts was derived in [8].
+For SPDEs driven by multiplicative noise with non-Lipschitz and potentially degenerate noise coefficients, results remain pretty scarce.
+Strong well-posedness has been proved in [29] for an SPDE (1.10) where $\sigma$ belongs to the class of Hölder continuous functions with exponent greater than $3/4$ and the drift $h$ is Lipschitz continuous.
+As for the SPDE (1.10) with $\sigma$ being a Hölder function with exponent less than or equal to $3/4$ and zero drift, only weak uniqueness is known for very specific noise coefficients $\sigma$ (such as, for example, $\sigma(u)=u^{\gamma}$ with $\gamma\geq 1/2$, or $\sigma(u)=\sqrt{u(1-u)}$); see [30], [34].
+In [34], weak uniqueness was established for the noise coefficient $\sigma(u)=\sqrt{u(1-u)}$ also in the presence of drifts of the form
+$h(x)=c_{1}(1-x)-c_{2}x+c_{3}x(1-x),\quad x\in[0,1],\;{\rm for}\;c_{1},c_{2}\geq 0,c_{3}\in\mathbb{R}.$
+Later, in [2], [27], [4], for $\sigma(u)=\sqrt{u(1-u)}$, the weak uniqueness results were extended to broader classes of drift terms.
+In [30], [34], [2], [4] the duality technique was used to prove uniqueness, whereas in [27] the Dawson-Girsanov theorem was applied.
+Now we turn to the super-Brownian motion (SBM), corresponding to the setting $\sigma(X)=\sqrt{X}$, that is SPDE (1.1).
+As noted earlier, the uniqueness for drifts of the form $h(X)=cX$ has been known for a long time, and follows from the duality method.
+It is also worth noting, that when SBM is subject to nonzero space-dependent immigration, meaning $h$ is continuous and depends only on the spatial parameter $x$ and not on $X$, pathwise non-uniqueness was shown in [10].
+Proving weak well-posedness for (1.1) requires establishing both weak existence and weak uniqueness.
+Weak existence, under the assumption that $h$ is continuous, follows from [35].
+However, no uniqueness results were provided in [35] for non-Lipschitz coefficients.
+For SBM and other related measure-valued branching processes the duality approach is a widely used method for establishing weak uniqueness.
 We will also employ this approach in our analysis.
-Since this method relies on guessing the correct duality functions it is sensitive to changes in the coefficients of the SPDE and thus one usually needs to assume a certain structure of the coefficients.
-The duality method was for instance used in [1, 21].
-Hereby, the authors of [1] are using moment duality and in [21] the Laplace transform is used.
-In this article we will also use the Laplace transform as a starting point for our duality relation.
-Our approach is essentially an extension of the well known duality relation between the Super Brownian Motion and the log Laplace equation.
-More precisely, if $X$ denotes the solution to (1.1) with $h=0$ and $V(\mu)$ is the solution to
+Our approach is essentially based on an extension of the well-known duality relation between SBM and the corresponding log-Laplace equation.
+More precisely, let $X$ denote the solution to (1.1) with $h=0$ and let $V(\mu)$ be the solution to
 $\begin{cases}&d_{t}V_{t}(\mu)=\frac{1}{2}\Delta V_{t}(\mu)-\frac{1}{2}V_{t}(\mu)^{2},\quad(t,x)\in(0,\infty)\times\mathbb{R}\\ &V_{t}(\mu)\Rightarrow\mu,\text{ as }t\rightarrow 0,\end{cases}$ (1.11)
-for a finite measure $\mu$, then it is well-known (see for instance [8]) that
-$\mathbb{E}[\exp(-\langle X_{t},\mu\rangle)]=\mathbb{E}[\exp(-\langle X_{0},V_{t}(\mu)\rangle)],\quad\forall t\geq 0,$
-for all finite measures $\mu$. Furthermore, if $h\equiv a>0$, we get the following duality relation
+for a finite measure $\mu$, then it is well-known (see for instance [21, Lemma 1.2]) that
+$\mathbb{E}[\exp(-\langle X_{t},\phi\rangle)]=\mathbb{E}[\exp(-\langle X_{0},V_{t}(\phi)\rangle)],\quad\forall t\geq 0,$
+for all non-negative, continuous and compactly supported functions $\phi$. Furthermore, if $h\equiv a>0$, the duality relation becomes:
 $\mathbb{E}[\exp(-\langle X_{t},\mu\rangle)]=\mathbb{E}\Big[\exp\Big(-\langle X_{0},V_{t}(\mu)\rangle-a\int_{0}^{t}\langle V_{s}(\mu),1\rangle ds\Big)\Big],\quad\forall t\geq 0,$ (1.12)
-which was for example demonstrated in [19].
-In this article, inspired by [21, 22] as well as [1], we are able to cover more complex immigration mechanisms $h$ by introducing a space-time jump process to (1.6).
-We outline this procedure in more detail in Section 4.
-More recently, the authors of [3] extended the results of [1] for the one-dimensional Wright–Fisher SPDE by allowing certain irregular drift terms that were previously not covered.
-A key element in their analysis was the phenomenon of coming down from infinity in the dual branching–coalescing particle system, which is discussed in detail in [2].
-Similarly, in this article, we employ the theory of initial trace solutions to the log-Laplace equation (see [6, 17, 20]) to extend the class of admissible drift functions $h$, allowing in particular for certain discontinuities at zero.
-The framework of initial trace solutions enables us to introduce jumps of infinite height in our dual process.
-In Section 4, we give some intuition on how this allows us to obtain a duality relation when $h$ has certain discontinuities at zero.
-We note that the theory of initial trace solutions is closely related to the phenomenon of coming down from infinity used in [3], both conceptually and technically.
-In comparison to [3], we are able to control exponential moments of the number of jumps of the dual process in sufficiently small time intervals.
-This allows us to extend the class of admissible drift terms $h$ to include functions with $b_{0}>0$ and $b_{1}\leq 0$.
-For this purpose, we introduce an artificial branching process that allows us to keep track of the number of jumps of the dual process. Again, this approach was inspired by [3].
+which can be verified by using Itô’s formula.
+In this article, we extend this approach for more complex drift terms $h$ by introducing a space-time jump noise into the log-Laplace equation (1.11).
+The resulting duality relation is presented in Proposition 4.3.
+In fact, Proposition 4.3 is also used in the proof of Theorem 1.2.
+We outline the construction of the dual process in detail in Section 4.
+We would like to highlight again the paper [4] which provided a number of very important insights that helped us in proving our main results.
+In [4] the well-posedness results of [2] were extended for (1.10) with $\sigma(u)=\sqrt{u(1-u)}$ for a large class of irregular drift terms, including some that are Hölder continuous and discontinuos at zero.
+The duality method, based on moment duality, was again used in [4], and the dual process was a branching–coalescing system of Brownian motions.
+A key element used in the analysis in [4] was the property of coming down from infinity of the dual branching–coalescing Brownian particle system.
+This property is discussed in detail in [3] in the case of for coalescing Brownian motions without branching.
+This phenomenon was crucial for establishing well-posedness in [4] for certain Hölder continuous and discontinuous at zero drifts.
+Analogously, in this article, we rely on the property of coming down from infinity, applied not to a particle system, but to solutions of the log-Laplace equation, whereas we use Laplace transform duality instead of moment duality.
+To successfully use this coming down from infinity property, we employ the theory of initial trace solutions to the log-Laplace equation (see [7, 24, 26]) which in turn allows us to extend the class of admissible drift functions $h$, particularly to those exhibiting discontinuities at zero.
+More precisely, this framework of initial trace solutions enables us to introduce jumps of infinite height in the dual process.
+In Section 4, we provide some intuition on how this leads to a duality relation even when $h$ has discontinuities at zero.
+Let us emphasize that the theory of initial trace solutions is conceptually closely related to the coming down from infinity property for particle systems used in [4].
+Compared to [4], we are able to control exponential moments of the number of jumps in the dual process over sufficiently small time intervals.
+This control allows us to extend the class of admissible drift terms $h$ to include cases with $b_{0}>0$ and $b_{1}\leq 0$, whose counterparts were not covered in [4].
+For instance, our results cover the drift $h(x)=b_{0}\mathbbm{1}_{x=0}$, that allows for upward drift only at points where solution equals to zero, an intriguing case for which no well-posedness results were previously known and which is not treated in [4].
+We believe that establishing well-posedness for SBMs with irregular drifts, especially those discontinuous at zero, addresses an important open question in the theory of super-Brownian motions.
+Furthermore, we expect these results can be extended to more general classes of superprocesses, such as $\alpha$-stable superprocesses with branching mechanisms that are not necessarily quadratic.
+It would also be interesting to construct a branching particle system whose scaling limit yields a solution to equation (1.1) with a drift discontinuous at zero.
+The discontinuity of the drift at zero in (1.1) suggests that, in the pre-limiting particle system, there may be different subcritical or supercritical branching mechanisms or mass immigrations, depending on whether the spatial regions have non-negligible or negligible particle "density".
+These questions are left for future work.
 Organization of the Article
-First, in Section 2, we introduce several function spaces and notation that is used throughout this article.
-In Section 3, we give a brief summary of results on certain non-linear parabolic equations, in particular the theory of initial traces and very singular solutions.
+First, in Section 2, we introduce several function spaces and notation used throughout this article.
+In Section 3, we provide a brief summary of results on certain non-linear parabolic equations, focusing in particular on the theory of initial traces and very singular solutions.
 Next, in Section 4, we outline the proof of Theorem 1.1.
-We then proceed to discuss several examples of admissible drift terms.
-In Section 6, we prove the duality relation (4.14).
-To this end, we introduce a branching process that allows us to control the moments of the number of jumps of the dual process in Section 6.1.
-We furthermore approximate the dual process by truncating the size of the jumps and prove the convergence of this approximation in Section 6.2.
-In Sections 7 and 8, we prove the weak existence and uniqueness of solutions by using the duality relation from Section 6.
-Finally, Section 9 is dedicated to the proof of Theorem 4.5, where we use the duality relation to prove properties of the cozero set of solutions to (1.1) under certain additional conditions.
+Then, in Section 5, we discuss several examples of admissible drift terms.
+Section 6 is devoted to the construction of the dual process, its properties, and the duality relation.
+In Section 6.1, we introduce a branching process that in particular allows us to control the moments of the number of jumps in the dual process.
+We also approximate the dual process by truncating the size of the jumps and prove the convergence of this approximation in Section 6.2.
+In Section 6.3, we derive an approximate duality, and finally, Proposition 4.3 (the duality relation) is proved.
+In Sections 7 and 8, we use Proposition 4.3 and other results to prove the weak existence and uniqueness of solutions.
+Section 9 is dedicated to the proof of Theorem 1.2, where we again use the duality relation.
+Finally, in Section 10 we prove Theorem 1.4.
 Convention for Constants
@@
 When necessary, dependence on additional parameters is expressed as $C=C(\lambda)$.
+Acknowledgments
+We want to thank Zhenyao Sun and Ed Perkins for very helpful conversations. The work of the authors was supported in part by the ISF grant No. 1985/22.
 2 Preliminaries
@@
   4. 2.4.
-    $\mathcal{P}(\mathcal{C}_{tem}^{+})$ denotes the set of all probability measures on $\mathcal{C}_{tem}$
+    $\mathcal{P}(\mathcal{C}_{tem}^{+})$ denotes the set of all probability measures on $\mathcal{C}_{tem}$.
   5. 2.5.
@@
   11. 2.11.
+    If $B\subseteq\mathbb{R}$ is an open set $\mathcal{M}_{Rad}^{+}(B)$ denote the set of non-negative Radon measures on $B$.
+  12. 2.12.
     $\mathcal{B}(\mathbb{R})$ denotes the Borel $\sigma$-algebra.
@@
 $\mathcal{W}_{t}(x,\nu)=V_{t}(\nu+\infty\delta_{x}),\quad t>0,x\in\mathbb{R}.$ (3.4)
-Furthermore, for $t<0$ we use the convention that $\mathcal{W}_{t}(A,\nu)=V_{t}(\mu)=0$ and $V_{0}(\mu)=\mu$.
+Furthermore, for $t<0$ we use the convention that $\mathcal{W}_{t}(A,\nu)=V_{t}(\mu)=0$ and $V_{0}(\mu)=\mu$, for $\mu\in\mathcal{M}_{F}^{+}$ and $\nu\in\mathcal{M}_{Rad}^{+}(A^{c})$.
 3.1 Finite Measures
@@
 We summarized their results in this section.
-First, we get get existence and uniqueness by [22, Lemma 2.1], which we restated in the following proposition
+First, we get get existence and uniqueness by [31, Lemma 2.1], which we restated in the following proposition
 Proposition 3.1.
@@
 where $(S_{t})_{t\geq 0}$ is the heat semi-group with transition density $(p_{t})_{t\geq 0}$ given by (4.1).
-The following lemma is a well known fact about (3.2) (see for instance [22, Lemma 2.1]).
+The following lemma is a well known fact about (3.2) (see for instance [31, Lemma 2.1]).
 Lemma 3.2.
@@
 for all $t\geq 0$ and $x\in\mathbb{R}$.
-The following lemma is [22, Lemma 2.6] and provides estimates that we shall use frequently.
+The following lemma is [31, Lemma 2.6] and provides estimates that we shall use frequently.
 Lemma 3.3.
@@
     for all $t\geq 0$.
-Finally, using [14, Lemma 2.3] we can infer the following lemma.
+Finally, using [20, Lemma 2.3] we can infer the following lemma.
 Lemma 3.4.
@@
     $\lim_{n\rightarrow\infty}\int\phi d\eta_{n}=\int\phi d\eta.$
-In the next sections we shall rely heavily on the following result, which is taken from [20, Proposition 3.10] [17, Theorem 4] and [3, Section 2.1].
+In the next sections we shall rely heavily on the following result, which is taken from [26, Proposition 3.10], [24, Theorem 4] and [4, Section 2.1].
 Proposition 3.6.
@@
   2. 3.2.
-    Let $((A_{n},\nu_{n}))_{n\geq 1}\subset\mathcal{T}$ such that $\eta^{A_{n},\nu_{n}}$ converges $m$-weakly to $\eta^{(A,\nu)}$, then it follows that $\mathcal{W}(A_{n},\nu_{n})$ converges to $\mathcal{W}(A,\nu)$ uniformly on compacts.
-In particular, by Lemma 3.3 and Proposition 3.6 it follows that $V_{t}(n\delta_{x})y$ converges monotonically to $\mathcal{W}_{t}(x,0)(y)$ for all $t>0$ and $x,y\in\mathbb{R}$. Therefore,
+[…]
+In particular, by Lemma 3.3 and Proposition 3.6 it follows that $V_{t}(n\delta_{x})(y)$ converges monotonically to $\mathcal{W}_{t}(x,0)(y)$ for all $t>0$ and $x,y\in\mathbb{R}$. Therefore,
 $V_{t}(n\delta_{x})(y)\leq\mathcal{W}_{t}(x,0)(y),\forall t>0,x,y\in\mathbb{R},\quad n\geq 1$ (3.6)
@@
 In this section, we outline the proof of Theorem 1.1 and, in particular, motivate the duality relation, which is the main tool in our analysis.
-Before we outline our arguments in more detail, we define what we understand under a weak solution to (1.1).
+Before we outline our arguments in more detail, we give a definition of weak solution to (1.1).
 Definition 4.1 (Weak Solution).
@@
 $p_{t}(x):=\frac{1}{\sqrt{2\pi t}}\operatorname{e}^{-x^{2}/(2t)}.$ (4.1)
-Both, the proof of weak uniqueness and existence in Theorem 1.1 are based on a duality approach.
+The proofs of weak uniqueness and of particular cases of weak existence in Theorem 1.1 are based on a duality approach.
 We first outline this approach for a simplified setting and then step by step add additional modifications to cover more general drift terms. Let us first consider the case where
-$h_{\infty}\equiv\langle\nu_{1},1\rangle\text{ and }\nu_{2}\equiv 0,$ (4.2)
+$h_{\infty}\equiv\langle\nu^{1},1\rangle\text{ and }\nu^{2}\equiv 0,$ (4.2)
 which means that
@@
 If we apply Itô’s formula to $\exp(-\langle\phi,X_{t}\rangle)$, for $\phi\in\mathcal{C}^{2}(\mathbb{R})$ with $\phi\geq 0$. We obtain that
-$\exp(-\langle\phi,X_{t}\rangle)-\int_{0}^{t}\exp(-\langle\phi,X_{s}\rangle)\Big{(}-\langle\frac{1}{2}\Delta\phi,X_{s}\rangle+\frac{1}{2}\langle\phi^{2},X_{s}\rangle-\langle\phi,h(X_{s})\rangle\Big{)}ds$ (4.3)
+$\exp(-\langle\phi,X_{t}\rangle)-\int_{0}^{t}\exp(-\langle\phi,X_{s}\rangle)\Big(-\langle\frac{1}{2}\Delta\phi,X_{s}\rangle+\frac{1}{2}\langle\phi^{2},X_{s}\rangle-\langle\phi,h(X_{s})\rangle\Big)ds,\quad t\geq 0,$ (4.3)
 is an $\mathcal{F}_{t}^{X}$-martingale. Suppose we can show the existence of a measure-valued process $Y$ such that
-$\exp(-\langle Y_{t},\psi\rangle)-\int_{0}^{t}\exp(-\langle Y_{s},\psi\rangle)\Big{(}-\langle\frac{1}{2}\Delta Y_{s},\psi\rangle+\frac{1}{2}\langle Y_{s}^{2},\psi\rangle-\langle Y_{s},h(\psi)\rangle\Big{)}ds,$ (4.4)
+$\exp(-\langle Y_{t},\psi\rangle)-\int_{0}^{t}\exp(-\langle Y_{s},\psi\rangle)\Big(-\langle\frac{1}{2}\Delta Y_{s},\psi\rangle+\frac{1}{2}\langle Y_{s}^{2},\psi\rangle-\langle Y_{s},h(\psi)\rangle\Big)ds,\quad t\geq 0,$ (4.4)
 is an $\mathcal{F}^{Y}_{t}$-martingale for $\psi\in\mathcal{C}_{tem}^{+}$. Then, if $X,Y$ are independent, under certain technical conditions it follows that
 $\mathbb{E}[\exp(-\langle X_{t},Y_{0}\rangle)]=\mathbb{E}[\exp(-\langle X_{0},Y_{t}\rangle)],\quad\forall t\geq 0,$
-by [11, Theorem 4.4.11]. This would allow us to infer that the one-dimensional laws are unique and thus by standard arguments the finite dimensional laws of $X$ are unique as well. Consider the SPDE
+by [15, Theorem 4.4.11]. This would allow us to infer that the one-dimensional laws are unique and thus by standard arguments the finite dimensional laws of $X$ are unique as well. Consider the SPDE
 $d_{t}Y_{t}=\frac{1}{2}\Delta Y_{t}-\frac{1}{2}Y_{t}^{2}+\int_{0}^{\infty}\int_{0}^{\infty}\int_{0}^{\infty}\lambda\mathbbm{1}_{r\leq Y_{t-}(x)}\mathcal{N}(dr,dx,d\lambda,dt),$ (4.5)
@@
 Now, we turn to the discontinuous part $h_{\infty}$. If we formally consider $\nu^{1}=d_{1}\delta_{\infty}$ and $\nu^{2}=d_{2}\delta_{\infty}$ with $d_{1},d_{2}\geq 0$ and plug it into
-$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+\int_{0}^{\infty}(1+\operatorname{e}^{-\lambda x})\nu^{2}(d\lambda)$ (4.10)
+$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+\int_{0}^{\infty}(1+\operatorname{e}^{-\lambda x})\nu^{2}(d\lambda),$ (4.10)
 we arrive at the function
@@
 Thus, for each fixed $n$, if we set
-$\nu_{n}^{1}=\nu^{1}_{|(0,n]}+d_{1}\delta_{n},\quad\nu_{n}^{2}=\nu^{2}_{|(0,n]}+d_{2}\delta_{n}$
+$\nu_{n}^{1}=\nu^{1}_{|(0,n]}+d_{1}\delta_{n},\quad\nu_{n}^{2}=\nu^{2}_{|(0,n]}+d_{2}\delta_{n},$
 we are again in the setting of (4.6) and (4.10) with $\nu^{i}$ replaced by $\nu^{i}_{n}$, $i=1,2$.
@@
 $\begin{split}h_{1}(x)&=\int_{0}^{\infty}\operatorname{e}^{-\lambda x}(\nu^{2}-\nu^{1})(d\lambda),\\ h_{\infty}(x)&=2d_{2}\mathbbm{1}_{x=0}+(d_{1}+d_{2})\mathbbm{1}_{x>0}+\langle\nu^{1}+\nu^{2},1\rangle.\end{split}$
-Since $d_{1},d_{2}\geq 0$ this does in particular not cover the regime $b_{0}\geq 0$ and $b_{1}\leq 0$. Introducing an additional parameter $a\in\mathbb{R}$ and setting
-$h_{\infty}(x)=2d_{2}\mathbbm{1}_{x=0}+(d_{1}+d_{2})\mathbbm{1}_{x>0}+\langle\nu^{1}+\nu^{2},1\rangle+a$ (4.12)
+Since $d_{1},d_{2}\geq 0$, for the case of $\nu^{1},\nu^{2}\equiv 0$ this, for example, does not cover the regime of $b_{0}>0$ and $b_{1}\leq 0$.
+Introducing an additional parameter $a\in\mathbb{R}$ and setting
+$h_{\infty}(x)=2d_{2}\mathbbm{1}_{x=0}+(d_{1}+d_{2})\mathbbm{1}_{x>0}+\langle\nu^{1}+\nu^{2},1\rangle+a,$ (4.12)
 allows us to cover the complete regime of parameters $b_{0}\geq\langle\nu^{1}-\nu^{2},1\rangle$ and $b_{1}\in\mathbb{R}$ for $h_{\infty}$ given by (1.4). More precisely, we set
 […]
-It is easily verified, that in this case $d_{1},d_{2}\geq 0$ and by plugging these definitions into (4.12) we recover (1.4).
+It is easily verified, that in this case of $d_{1},d_{2}\geq 0$, by plugging these definitions into (4.12) we recover (1.4).
 Remark 4.2.
-There is not necessarily a unique choice for $\nu^{1},\nu^{2}$ and $d_{1}$, $d_{2}$ and $a$ that represent a given function $h$.
-In order to handle this class of drift function we need to modify the duality function again. In this case we will see in Section 6 that the approximate duality relation is given by
+There is not necessarily a unique choice of $\nu^{1},\nu^{2}$ and $d_{1}$, $d_{2}$ and $a$ that allows to represent a given function $h$.
+In order to handle this class of drift functions we need to modify the duality function again. In this case we will see in Section 6 that the approximate duality relation is given by
 $\begin{split}&\mathbb{E}\Big[\exp\Big(-\langle X_{t},Y_{0}\rangle\Big)\Big]\\ &\quad=\lim_{n\rightarrow\infty}\mathbb{E}\Big[(-1)^{J_{t}^{n}}\exp\Big(-\langle X_{0},Y_{t}^{n}\rangle\Big)\exp\Big(-a\int_{0}^{t}\langle Y_{s}^{n},1\rangle ds\Big)\Big],\quad\forall t\geq 0,\end{split}$ (4.13)
@@
 As before, we are able to construct the limits of $Y^{n}$ and $J^{n}$ and thus do away with the limit in (4.13).
-These considerations lead us to the following proposition, which we prove in Section 6
+These considerations lead us to the following proposition, which we will prove in Section 6
 Proposition 4.3.
@@
 Let $Y_{0}=\mu$ for $\mu\in\mathcal{M}_{F}^{+}$.
-Then, there exists $T_{0}>0$ such that there exist processes $(J_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$, taking values in $\mathbb{N}$ and $\mathcal{M}_{reg}^{+}$, respectively, that are independent of $X$ and such that
+Then, there exists $T_{0}>0$ such that there exist processes $(J_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$, taking values in $\mathbb{N}\cup\{0\}$ and $\mathcal{M}_{reg}^{+}$, respectively, that are independent of $X$ and such that
 $\mathbb{E}\Big[\exp\Big(-\langle X_{t},Y_{0}\rangle\Big)\Big]=\mathbb{E}\Big[(-1)^{J_{t}}\exp\Big(-\langle X_{0},Y_{t}\rangle\Big)\exp\Big(-a\int_{0}^{t}\langle Y_{s},1\rangle ds\Big)\Big],$ (4.14)
@@
 Note that $(Y,J)$ are the limits of $(J^{n},Y^{n})$ in (4.13) that we discussed above with $d_{1},d_{2}$ as in (A).
-Finally, when $\nu_{2}\equiv 0$ and $h_{\infty}\equiv\langle\nu_{1},1\rangle$ we can use Proposition 4.3 to show that the cozero set of solutions to (1.1) has finite Lebesgue measure.
-That is, we obtain the following result.
-Theorem 4.5.
-Suppose that almost surely, $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$.
-And let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $\nu_{n}^{2}\equiv 0$ and $b_{0}=0$.
-Then it follows that there exists $T_{0}>0$ such that for all $t\in[0,T_{0}]$ we have
-$\mathbb{P}(\mathrm{Leb}(\{x\in\mathbb{R}\colon X_{t}(x)>0\})<\infty)=1.$
+Finally, when $\nu^{2}\equiv 0$ and $h_{\infty}\equiv\langle\nu^{1},1\rangle$ we can use Proposition 4.3 to prove Theorem 1.2.
 5 Examples
@@
 Let us now consider the case, where $b_{0}=b_{1}=0$. That means, we have
-$h(x)=h_{1}(x)=\int_{0}^{\infty}\operatorname{e}^{-\lambda x}(\nu_{2}-\nu_{1})(d\lambda)$ (5.1)
-where $\langle\nu_{2}-\nu_{1},1\rangle\geq 0$.
+$h(x)=h_{1}(x)=\int_{0}^{\infty}\operatorname{e}^{-\lambda x}(\nu^{2}-\nu^{1})(d\lambda)$ (5.1)
+where $\langle\nu^{2}-\nu^{1},1\rangle\geq 0$.
 By Bernsteins theorem on completely monotone functions (see [5]) all functions in the span of completely monotone functions that are non-negative at zero take the form (5.1).
@@
 $\int_{0}^{t}\mathbbm{1}_{y_{s}=0}ds=0.$
-However, this can be proven along the same lines as the proof of [3, Lemma 1.3]. That is,by [24, Theorem VI.1.7] we get that
+However, this can be proven along the same lines as the proof of [4, Lemma 1.3]. That is, by [33, Theorem VI.1.7] we get that
 $L_{t}^{0}(y)=2\int_{0}^{t}\mathbbm{1}_{y_{s}=0}ds,$
-where $L_{t}^{0}(y)$ is the local time of $y$ at 0. By the same arguments as in the proof of [24, Proposition XI.1.5] we get that $L_{t}^{0}(y)\equiv 0$. This finishes the proof of $(i)$.
+where $L_{t}^{0}(y)$ is the local time of $y$ at 0. By the same arguments as in the proof of [33, Proposition XI.1.5] we get that $L_{t}^{0}(y)\equiv 0$. This finishes the proof of $(i)$.
 Next we prove $(ii)$. In order to prove the claim, we show that any solution to (5.4) with $c\in(0,1/2)$ solves
@@
 Finally, we prove $(iii)$.
-Assume that there exists a weak solution $x$ to (5.5) and note that this solution has to be non-negative by a slight modification of [16, Theorem 1.4 ].
+Assume that there exists a weak solution $x$ to (5.5) and note that this solution has to be non-negative by a slight modification of [23, Theorem 1.4].
 Then, similar to above we get that
@@
 Proof.
-The proof is similar to the proof of [3, Lemma 1.2].
+The proof is similar to the proof of [4, Lemma 1.2].
 Fix $c_{1},c_{2}\in(0,1)$ with $c_{1}\neq c_{2}$ and denote by $X^{1},X^{2}$ the solution to (5.7) with drift $h_{c_{1},1}$ and $h_{c_{2},1}$, respectively.
@@
 To this end assume that $X^{1}\overset{d}{=}X^{2}$.
-Then it follows along the same lines as in the prove of [3, Lemma 1.2] that
+Then it follows along the same lines as in the prove of [4, Lemma 1.2] that
 $\begin{split}c_{1}\mathbb{E}\Big[\int_{\mathbb{R}}\int_{0}^{t}\phi(x)\mathbbm{1}_{X_{s}^{1}(x)=0}dsdx\Big]=c_{2}\mathbb{E}\Big[\int_{\mathbb{R}}\int_{0}^{t}\phi(x)\mathbbm{1}_{X_{s}^{2}(x)=0}dsdx\Big]\end{split}$
@@
 $\mathbb{E}[\mathbbm{1}_{X_{t}^{i}(x)=0}]>0,\quad\forall t>0,x\in\mathbb{R},i=1,2,$
-by a comparison argument, which follows along similar lines to the proof of [3, Lemma 1.2]. Let $X$ be the unique weak solution to
+by a comparison argument, which follows along similar lines to the proof of [4, Lemma 1.2]. Let $X$ be the unique weak solution to
 $dX_{t}=\frac{1}{2}\Delta X_{t}+1dt+\sqrt{X_{t}}\dot{W},\quad x_{0}=f,\quad t\geq 0,$
@@
 And that $X^{i}$ as well as $X^{1}$ can be constructed as weak limits of an SPDE with Lipschitz continuous coefficients.
-Thus, by [25, Corollary 2.4] and weak uniqueness we get that $X$ stochastically dominates $X^{i}$, which implies
+Thus, by [35, Corollary 2.4] and weak uniqueness we get that $X$ stochastically dominates $X^{i}$, which implies
 $\mathbb{E}[\mathbbm{1}_{X_{t}^{i}(x)=0}]\geq\mathbb{E}[\mathbbm{1}_{X_{t}(x)=0}],\quad\forall t>0,x\in\mathbb{R},i=1,2.$
-However, using the duality formula for the Super Brownian motion with immigration given by (1.7) we get for $t>0$ that
+However, using the duality formula for the Super-Brownian motion with immigration given by (1.12) we get for $t>0$ that
 $\begin{split}&\mathbb{E}[\mathbbm{1}_{X_{t}(x)=0}]=\lim_{m\rightarrow\infty}\mathbb{E}[\exp(-mX_{t}(x))]\\ &=\lim_{m\rightarrow\infty}\exp\Big(-\langle V_{t}(m\delta_{x}),X_{0}\rangle-\int_{0}^{t}\langle V_{s}(m\delta_{x}),1\rangle ds\Big).\end{split}$
@@
 where $d_{1},d_{2}$ are defined as in (A) and
-$\nu([b,\infty])=(\nu^{1}+\nu^{2})([b,\infty))+d_{1}+d_{2},\quad b\in\overline{\mathbb{R}}_{+}.$
+$\nu([b,\infty])=(\nu^{1}+\nu^{2})([b,\infty))+d_{1}+d_{2},\quad b\in\mathbb{R}_{+}.$ (6.1)
 Furthermore, assume that $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu$ is a non-negative finite measure, $m\in\mathbb{R}_{+}\cup\{\infty\}$, and $x_{0}\in\mathbb{R}$.
@@
 Furthermore, let $M_{1}$ be an independent random variable with values in $\{1,2\}$ given by
-$\mathbb{P}(M_{k}=i)=\frac{\nu^{i}((0,\infty))+d_{i}}{\nu((0,\infty])},\quad i\in\{1,2\},$ (6.1)
+$\mathbb{P}(M_{1}=i)=\frac{\nu^{i}((0,\infty))+d_{i}}{\nu((0,\infty])},\quad i\in\{1,2\},$ (6.2)
 and let $Z_{1}$ be a random variable such that
@@
 These bounds are crucial in order to prove that $Y$ only has finitely many jumps in each finite time interval and in order to show that all the terms in (4.14) have finite moments.
-To this end, we introduce an artificial branching mechanism that will allow us to keep track of the number of jumps of $Y$.
+To this end, we introduce an artificial branching mechanism that will allow us to control the number of jumps of $Y$.
 This approach is inspired by the ideas from [4].
 Let
-$I_{t}(Y_{0})\subset\mathcal{U}=\cup_{k=1}^{\infty}\mathbbm{N}^{k}$
-be the collections of labels of particles that are alive at time $t$. And denote by
+$I^{\nu}_{t}(Y_{0})\subset\mathcal{U}=\cup_{k=1}^{\infty}\mathbbm{N}^{k}$
+be the collections of labels of particles that are alive at time $t$. When there is no ambiguity we may suppress the superscript $\nu$ and just write $I_{t}(Y_{0})$ . And denote by
 $|I_{t}(Y_{0})|,$
@@
 Below we give a description of the branching mechanism.
-  1. 6.1.
+  1. 6(i)
     We denote the first particle by $\mathcal{Z}^{(1)}=(Y^{(1)},\mathfrak{M}^{(1)})$, where $Y^{(1)}_{t}=V_{t}(Y_{0})$ for $t\geq 0$ and we let $\mathfrak{M}^{(1)}(t)$ for $t\geq 0$ be a Poisson point process on $(0,\infty]\times(0,\infty)^{2}\times\mathbb{R}$ with intensity
@@
     We furthermore denote by $\mathfrak{N}^{(1)}$ the random measure defined by
-    $\mathfrak{N}^{(1)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{Y_{t}^{(1)}(x)\leq r}dxdrdt,$
+    $\mathfrak{N}^{(1)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{r\leq Y_{t}^{(1)}(x)}dxdrdt,$
     for $G_{1}\subseteq(0,\infty],G_{2},G_{3}\subseteq(0,\infty)$and $G_{4}\subseteq\mathbb{R}$ are measurable sets.
-  2. 6.2.
-    To every particle $\mathcal{Z}^{(\alpha)}$ that was born at time $\tau_{\alpha}$ we associate a tuple $(Y^{\alpha},\mathfrak{M}^{\alpha})$, where $Y^{(\alpha)}_{t}=\mathcal{W}_{t-\tau_{\alpha}-}(0,0)\mathbbm{1}_{t>\tau_{\alpha}}$.
+  2. 6(ii)
+    To every particle $\mathcal{Z}^{(\alpha)}$ that was born at time $\tau_{\alpha}$ we associate a tuple $(Y^{(\alpha)},\mathfrak{M}^{(\alpha)})$, where $Y^{(\alpha)}_{t}=\mathcal{W}_{t-\tau_{\alpha}-}(0,0)\mathbbm{1}_{t>\tau_{\alpha}}$.
 Furthermore, $\mathfrak{M}^{(\alpha)}$ is a Poisson point process on $(0,\infty]\times(0,\infty)^{2}\times\mathbb{R}$ with intensity given by (6.7).
@@
 Moreover, we define the random measure $\mathfrak{N}^{(\alpha)}$ given by
-    $\mathfrak{N}^{(\alpha)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{Y_{t}^{(\alpha)}(x)\leq r}dxdrdt,$
+    $\mathfrak{N}^{(\alpha)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{r\leq Y_{t}^{(\alpha)}(x)}dxdrdt,$
     for $G_{1}\subseteq(0,\infty],G_{2},G_{3}\subseteq(0,\infty)$and $G_{4}\subseteq\mathbb{R}$ are measurable sets.
-  3. 6.3.
-    A particle $\mathcal{Z}^{(\alpha)}$, $\alpha\in\mathcal{U}$ gives birth to a new particles at time $\tau$, which is an arrival time of $\mathfrak{N}^{(\alpha)}$.
-If $\tau$ is the time of the $i$-th branching of $\mathcal{Z}^{(\alpha)}$ then the new particle will be labeled $\beta=(\alpha,i)$ and the parent keeps the label $\alpha$.
-Recall that we denote by $I_{t}(Y_{0})$ the set of particles that are alive at time $t$. That is we have
-$I_{t}(Y_{0})=\{\alpha\in\mathcal{U}\colon\langle Y_{t}^{(\alpha)},1\rangle>0\}.$
+  3. 6(iii)
+    A particle $\mathcal{Z}^{(\alpha)}$, $\alpha\in\mathcal{U}$ gives births to new particles at arrival times of $\mathfrak{M}^{(\alpha)}$.
+For the $i$-th branching of $\mathcal{Z}^{(\alpha)}$, the new particle will be labeled $\beta=(\alpha,i)$ and the parent keeps the label $\alpha$.
+Recall that we denote by $I_{t}(Y_{0})$ the set of particles that are alive at time $t$. That is, we have
+$I_{t}(Y_{0})=\{\alpha\in\mathcal{U}\colon\langle Y_{t}^{(\alpha)},1\rangle>0\}.$ (6.8)
 The next lemmas allow us to control the moments of the number of jumps of the dual process.
@@
 Proof.
-First we prove the statement for $T_{0}>0$ sufficiently small. First, we define
-$I_{t}^{0}=\{(1)\},\quad I_{t}^{i}=\{\alpha\in\mathcal{U}||\alpha|=i+1,\langle Y_{t}^{(\alpha)},1\rangle>0\},\quad i\geq 1$
+Define
+$I_{t}^{0}=\{(1)\},\quad I_{t}^{i}=\{\alpha\in\mathcal{U}||\alpha|=i+1,\langle Y_{t}^{(\alpha)},1\rangle>0\},\quad i\geq 1,$
 where $|\alpha|$ denotes the length of the vector $\alpha$. Let $T_{0}>0$ be chosen sufficiently small and $t\leq T_{0}$. The precise value of $T_{0}$ will be determined later. We will show that
@@
 First, we prove (6.9) for the case $i=1$.
-The expected number of children of the particle $Z^{(1)}$ until time $t\in[0,T_{0}]$ is given by the expectation of $\mathfrak{N}^{(1)}_{t}=\mathfrak{N}^{(1)}((0,\infty],(0,t],(0,\infty),\mathbb{R})$, that is
+The expected number of children of the particle $Z^{(1)}$ until time $t\in[0,T_{0}]$ is given by the expectation of $\mathfrak{N}^{(1)}_{t}=\mathfrak{N}^{(1)}((0,\infty]\times(0,t]\times(0,\infty)\times\mathbb{R})$, that is
 $\begin{split}&\mathbb{E}{|I_{t}^{1}|}=\mathbb{E}[\mathfrak{N}^{(1)}_{t}]\\ &=\nu((0,\infty])\mathbb{E}\Big[\int_{0}^{t}\int_{\mathbb{R}}Y_{s}^{(1)}(x)dxds\Big]=\nu((0,\infty])\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds.\end{split}$
-Since $(t,x)\mapsto V_{t}(Y_{0})(x)$ is integrable on $0,T]\times\mathbb{R}$ for all $T>0$, we can find $T_{0}>0$ sufficiently small so that
+Since, by Lemma 3.2, $(t,x)\mapsto V_{t}(Y_{0})(x)$ is integrable on $0,T]\times\mathbb{R}$ for all $T>0$, we can find $T_{0}>0$ sufficiently small so that
 $\mathbb{E}{|I_{t}^{1}|}\leq 2^{-1}.$
@@
 Then it follows that
-$\mathbb{E}[|I_{t}^{{i+1}}|]=\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}^{i}}\mathfrak{N}^{(\alpha)}_{t}\Big{]}\leq\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}^{i}}\tilde{\mathfrak{N}}^{(\alpha)}_{t}\Big{]}\leq\mathbb{E}[|I_{t}^{i}|]\nu((0,\infty])\int_{0}^{t}\int_{\mathbb{R}}\mathcal{W}_{s}(0,0)(x)dxds\leq 2^{-i}2^{-1},$
+$\begin{split}\mathbb{E}[|I_{t}^{{i+1}}|]&=\mathbb{E}\Big[\sum_{\alpha\in I_{t}^{i}}\mathfrak{N}^{(\alpha)}((0,\infty]\times(0,t]\times(0,\infty)\times\mathbb{R})\Big]\\ &\leq\mathbb{E}\Big[\sum_{\alpha\in I_{t}^{i}}\tilde{\mathfrak{N}}^{(\alpha)}_{t}\Big]\leq\mathbb{E}[|I_{t}^{i}|]\nu((0,\infty])\int_{0}^{t}\int_{\mathbb{R}}\mathcal{W}_{s}(0,0)(x)dxds\leq 2^{-i}2^{-1},\end{split}$
 if we choose $T_{0}$ sufficiently small, so that
 $\nu((0,\infty])\int_{0}^{T_{0}}\int_{\mathbb{R}}\mathcal{W}_{s}(0,0)(x)dxds\leq 2^{-1}.$
-Hereby, we used that the intensity of $\tilde{\mathfrak{N}}^{(\alpha)}$ is larger than the intensity of $\mathfrak{N}^{(\alpha)}((0,\infty],\cdot,(0,\infty),\mathbb{R})$ for each $\alpha$.
+Hereby, we used that the intensity of $\tilde{\mathfrak{N}}^{(\alpha)}$ bounds the intensity of $\mathfrak{N}^{(\alpha)}((0,\infty],dt,(0,\infty),\mathbb{R})$ from above for each $\alpha$.
 Thus, we arrive at
-$\mathbb{E}[|I_{t}|]\leq\sum_{i=0}^{\infty}\mathbb{E}[|I_{t}^{i}|]\leq 1+\sum_{i=1}^{\infty}2^{-i},$
+$\mathbb{E}[|I_{t}(Y_{0})|]=\sum_{i=0}^{\infty}\mathbb{E}[|I_{t}^{i}|]\leq 1+\sum_{i=1}^{\infty}2^{-i},$
 for all $0\leq t\leq T_{0}$.
@@
 $\mathbb{E}[|I^{\alpha}_{T_{0},MT_{0}}|]\leq\mathbb{E}[|I_{(M-1)T_{0}}(\mathcal{W})|],$
-where $I_{t}(\mathcal{W})$ denotes the particles that are alive at time $t$ in the branching process described in 6.1.-6.3.
-where $Y^{(1)}$ is replaced by $\mathcal{W}(0,0)$.
+where $I_{t}(\mathcal{W})$ denotes the particles that are alive at time $t$ in the branching process described in 6(i)-6(iii), where $Y^{(1)}$ is replaced by $\mathcal{W}(0,0)$.
 This simply follows from the fact that
@@
 which follows from Proposition 3.7. Thus, we iteratively get that
-$\begin{split}&\mathbb{E}[|I_{T}(Y_{0})|]\leq\mathbb{E}[|I_{MT_{0}}(Y_{0})|]\\ &\leq\mathbb{E}[|I_{T_{0}}(Y_{0})|-1]\mathbb{E}[|I_{(M-1)T_{0}}(\mathcal{W})|]+\mathbb{E}[|I_{(M-1)T_{0}}(V_{T_{0}}(Y_{0}))|]\leq\ldots\leq C(T)\end{split}$
+$\begin{split}&\mathbb{E}[|I_{T}(Y_{0})|]\leq\mathbb{E}[|I_{MT_{0}}(Y_{0})|]\\ &\leq\mathbb{E}[|I_{T_{0}}(Y_{0})|-1]\mathbb{E}[|I_{(M-1)T_{0}}(\mathcal{W})|]+\mathbb{E}[|I_{(M-1)T_{0}}(V_{T_{0}}(Y_{0}))|]\leq\ldots\leq C(T).\end{split}$
 This finishes the proof. ∎
 Lemma 6.2.
-Let $Y_{0}=\mu+m\delta_{x_{0}}$ with $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then for every $\gamma>0$, there exists $T_{0}>0$ such that
-$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}(Y_{0})||}]<\infty,$ (6.8)
-for all $t\leq T_{0}$ . Hereby, $T_{0}>0$ only depends on $\nu((0,\infty])$, $\langle\mu,1\rangle$ and $\gamma$.
+Let $Y_{0}=\mu+m\delta_{x_{0}}$ with $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$.
+Furthermore, let $I_{t}(Y_{0})$ be defined as in (6.8).
+Then for every $\gamma>0$, $K>0$, there exists $T_{0}>0$ such that
+$\sup_{\nu:\nu((0,\infty])\leq K}\mathbb{E}[\operatorname{e}^{\gamma|I_{t}(Y_{0})||}]<\infty,$ (6.10)
+for all $t\leq T_{0}$ . Hereby, $T_{0}>0$ only depends on $K$, $\langle\mu,1\rangle$ and $\gamma$.
 Proof.
-Fix $t\leq T_{0}$, where $T_{0}$ is sufficiently small, the precise value will be chosen later.
+Fix arbitrary $\gamma>0$, $K>0$ and $t\leq T_{0}$, where $T_{0}$ is sufficiently small, the precise value will be chosen later.
 Define $\eta_{t}^{i,k}$ for $i,k\in\mathbb{N}$ to be i.i.d.
 Poisson distributed random variables with parameter
-$\lambda=\nu((0,\infty])\Big{(}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds\Big{)}.$
+$\lambda=K\Big(\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds\Big).$ (6.11)
 Furthermore, set $Z_{0}=1$ and
 $Z_{n}^{t}=\sum_{k=1}^{Z_{n-1}^{t}}\eta_{t}^{n,k},\quad n\geq 1.$
-Then it easily follows that $|I_{t}^{i}|$ is stochastically dominated by $Z_{i}^{t}$ for each $i\geq 1$. Thus, we can also infer that
-$|I_{t}|=\sum_{i=0}^{\infty}|I_{t}^{i}|$
+Let $\nu$ be a measure such that $\nu((0,\infty])\leq K$. Then it easily follows that $|I_{t}^{i}|$ is stochastically dominated by $Z_{i}^{t}$ for each $i\geq 1$. Thus, we can also infer that
+$|I_{t}(Y_{0})|=\sum_{i=0}^{\infty}|I_{t}^{i}|$
 is stochastically dominated by
@@
 The random variable $Z^{t}$ is nothing else, but the total progeny of $(Z^{t}_{n})_{n\geq 0}$ and its law can be made explicit by using the Otter-Dwass formula (see [13]), whenever $\lambda\leq 1$.
-Thus $T_{0}$ must be small enough such that $\lambda\leq 1$.
-In our setting, we get
+By (6.11) we can choose $T_{0}$ small enough such that
+$\lambda\leq 1,\quad\log(\lambda)<\gamma-1,\quad\forall t\leq T_{0}.$ (6.12)
+Then we get
 $\mathbb{P}(Z^{t}=k)=\operatorname{e}^{-\lambda k}\frac{(\lambda k)^{k-1}}{k!},\quad k\geq 1.$
-which is the Borel-Tanner distribution. Using Stirling’s approximation, we thus get that
+that is $Z^{t}$ has the Borel-Tanner distribution. Using Stirling’s approximation, we thus get that
 $\mathbb{P}(Z^{t}=k)\asymp\sqrt{2\pi}\operatorname{e}^{(1-\lambda)k}\lambda^{k-1}k^{-3/2}\asymp\sqrt{2\pi}\operatorname{e}^{(1-\lambda+\log(\lambda))k-\log(\lambda)}k^{-3/2}$
@@
 $\sum_{k=1}^{\infty}k^{-3/2}\operatorname{e}^{(\gamma+1-\lambda+\log(\lambda))k}<\infty,$
-whenever
-$\gamma\leq\lambda-1-\log(\lambda).$ (6.9)
-When $\lambda\rightarrow 0$, we get that $\lambda-1-\log(\lambda)\rightarrow\infty$.
-If we choose $T_{0}>0$ small enough, we thus get that (6.9) is satisfied and we obtain that
+by (6.12). Therefore, we obtain that
 $\mathbb{E}[\operatorname{e}^{\gamma Z^{t}}]<\infty.$
-Note that the function $x\mapsto\operatorname{e}^{\gamma x}$ is non-decreasing. Thus, the fact that $|I_{t}|$ is stochastically dominated by $Z^{t}$ implies that
-$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}|}]\leq\mathbb{E}[\operatorname{e}^{\gamma Z^{t}}]<\infty.$
-This completes the proof. ∎
+Note that the function $x\mapsto\operatorname{e}^{\gamma x}$ is non-decreasing. Thus, the fact that $|I_{t}(Y_{0})|$ is stochastically dominated by $Z^{t}$ implies that
+$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}(Y_{0})|}]\leq\mathbb{E}[\operatorname{e}^{\gamma Z^{t}}]<\infty.$
+Since $\nu$ such that $\nu((0,\infty])\leq K$ was arbitrary, this completes the proof. ∎
 Now, we need to relate the branching process to the process $Y$.
-Since the birth times of new particles happen with higher frequency than the jumps of $t\mapsto\langle Y_{t},1\rangle$, we aim to control its time integral, which is sufficient for our arguments.
-Recall that $(S_{i})_{i=1}^{\infty}$ is the sequence of independent $\exp(1)$-distributed random variables from the construction of $Y$ as defined in (6.2).
+Since the birth times of new particles in the branching process happen with higher rate than the jumps of $t\mapsto\langle Y_{t},1\rangle$, we aim to control $\int_{0}^{t}\langle Y_{s},1\rangle ds$, which is sufficient for our arguments.
+Recall that $(S_{i})_{i=1}^{\infty}$ is the sequence of independent $\exp(1)$-distributed random variables used for the construction of $Y$ as defined in the beginning of Section 6.
 We now define random variables $(\bar{R}_{i})_{i=1}^{\infty}$ as follows
@@
 Then, by construction we get the following identity
-$\hat{Y}_{t}=V_{t}(Y_{0})+\sum_{k=1}^{|I_{t}|}\mathcal{W}_{t-\bar{R}_{k}}(0,0).$ (6.10)
+$\hat{Y}_{t}=V_{t}(Y_{0})+\sum_{k=1}^{|I_{t}(Y_{0})|}\mathcal{W}_{t-\bar{R}_{k}}(0,0).$ (6.13)
 Using this construction we can prove our first estimates.
@@
 $\bar{T}_{i}-\bar{T}_{k}=\sum_{\ell=k+1}^{i}T_{\ell}\geq\bar{R}_{i}-\bar{R}_{k},$
-for all $0\leq k\leq i$. By the monotonicity relations
+for all $0\leq k\leq i$. We have the monotonicity relations
 $\langle\mathcal{W}_{s}(0,0),1\rangle\leq\langle\mathcal{W}_{t}(0,0),1\rangle\quad 0<t\leq s,$ (6.17)
-$\langle V_{s}(Y_{0}),1\rangle\leq\langle V_{t}(Y_{0}),1\rangle\quad 0<t\leq s,$ (6.15)
-we therefore arrive at
-$\begin{split}&\langle Y_{\bar{T}_{i}+s},1\rangle\leq\langle V_{s+\bar{R}_{i}}(Y_{0}),1\rangle+\sum_{k=1}^{i}\langle\mathcal{W}_{s+\bar{R}_{i}-\bar{R}_{k}}(0,0),1\rangle\\ &\leq\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle\quad a.s.,\end{split}$
-for all $0<s<\bar{T}_{i+1}$. Hereby, (6.14) follows from Proposition 3.7 and (6.15) follows from from integrating (3.5).
-Now, we show
+$\langle V_{s}(Y_{0}),1\rangle\leq\langle V_{t}(Y_{0}),1\rangle\quad 0<t\leq s.$ (6.18)
+Hereby, (6.17) follows from Proposition 3.7 and (6.18) follows from integrating (3.5). We arrive at
+$\begin{split}&\langle Y_{\bar{T}_{i}+s},1\rangle\leq\langle V_{s+\bar{R}_{i}}(Y_{0}),1\rangle+\sum_{k=1}^{i}\langle\mathcal{W}_{s+\bar{R}_{i}-\bar{R}_{k}}(0,0),1\rangle\\ &\leq\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle\quad a.s.,\end{split}$ (6.19)
+for all $0<s<\bar{T}_{i+1}$.
+Now, we will show
 $R_{i+1}\leq T_{i+1}$ (6.20)
@@
 $\int_{0}^{t}\langle Y_{s},1\rangle ds\leq\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds,\quad\forall t\in(\bar{T}_{i},\bar{T}_{i+1}).$
-Let $t\in(\bar{T}_{i},\bar{T}_{i+1})$ be fixed but arbitrary and define
+Fix arbitrary $t\in(\bar{T}_{i},\bar{T}_{i+1})$ and define
 $j^{\ast}=\sup\{k\geq 1\colon\bar{R}_{k}\leq t\}.$
@@
 $\int_{\bar{T}_{i}}^{t}\langle Y_{s},1\rangle ds\leq\int_{\bar{R}_{i}}^{t}\langle\hat{Y}_{s},1\rangle ds.$
-Using (6.12) for $i+1$ and using that $\bar{R}_{i}\leq\bar{T}_{i}$ we get
+Using (6.15) for $i+1$, which follows from (6.19), and using that $\bar{R}_{i}\leq\bar{T}_{i}$, we get
 $\begin{split}&\int_{\bar{T}_{i}}^{t}\langle Y_{s},1\rangle ds=\int_{0}^{t-\bar{T}_{i}}\langle Y_{\bar{T}_{i}+s},1\rangle ds\leq\int_{0}^{t-\bar{T}_{i}}\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle ds\\ &\quad\leq\int_{0}^{t-\bar{R}_{i}}\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle ds=\int_{\bar{R}_{i}}^{t}\langle\hat{Y}_{s},1\rangle ds.\end{split}$
@@
 Proposition 6.4.
-Let $Y$ be the process constructed in (6.5) with $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then it follows that
-$\sup_{n}\mathbb{E}\Big{[}\int_{0}^{T}\langle Y_{t},1\rangle dt\Big{]}\leq C(T),$ (6.17)
+Let $Y$ be the process constructed in (6.6) with $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$.
+Then it follows for all $T>0$ that
+$\sup_{m}\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{T}\langle Y_{t},1\rangle dt\Big]\leq C(T),$ (6.21)
 where $C(T)$ only depends on $\nu((0,\infty])$ and $\langle\mu,1\rangle$.
@@
 Since $\bar{R}_{i}\rightarrow\infty$ as $i\rightarrow\infty$ we get by the monotone convergence theorem and using (6.16) that
-$\begin{split}&\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{]}=\lim_{i\rightarrow\infty}\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\mathbbm{1}_{t\leq\bar{T}_{i}}\Big{]}\\ &\leq\mathbb{E}[\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds].\end{split}.$
+$\begin{split}&\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s},1\rangle ds\Big]=\lim_{i\rightarrow\infty}\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s},1\rangle ds\mathbbm{1}_{t\leq\bar{T}_{i}}\Big]\\ &\leq\mathbb{E}_{Y_{0}^{m}}[\int_{0}^{t}\langle\hat{Y}_{s},1\rangle ds].\end{split}.$
 By using Lemma 6.1, this immediately yields the following inequality
-$\begin{split}&\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{]}\leq\mathbb{E}\Big{[}\int_{0}^{t}\sum_{\alpha\in I_{s}}\langle Y_{s}^{(\alpha)},1\rangle ds\Big{]}\\ &\leq\mathbb{E}\Big{[}\int_{0}^{t}\langle Y_{s}^{(1)},1\rangle ds\Big{]}+\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds\Big{]}\\ &\leq\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds+C\mathbb{E}[I_{t}]\leq C(t),\end{split}$
+$\begin{split}&\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s},1\rangle ds\Big]\leq\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\sum_{\alpha\in I_{s}}\langle Y_{s}^{(\alpha)},1\rangle ds\Big]\\ &\leq\mathbb{E}_{Y_{0}^{m}}\Big[\int_{0}^{t}\langle Y_{s}^{(1)},1\rangle ds\Big]+\mathbb{E}\Big[\sum_{\alpha\in I_{t}}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds\Big]\\ &\leq\int_{0}^{t}\langle V_{s}(Y_{0}^{m}),1\rangle ds+C(t)\mathbb{E}[I_{t}]\leq C(t),\end{split}$
 where $C(t)>0$ only depends on $\nu((0,\infty])$ and $\langle\mu,1\rangle$. ∎
 Lemma 6.5.
-Let $Y_{0}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$.
-Then, for every $\gamma\in\mathbb{R}$, there exists $T_{0}>0$ such that for all $t\leq T_{0}$ we have
-$\mathbb{E}\Big{[}\exp\Big{(}-\gamma\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{)}\Big{]}<C,$
-Hereby, $T_{0}>0$ and $C>0$ only depend on $\langle\mu,1\rangle$ and $\nu((0,\infty])$.
+Let $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$.
+Let $Y(\nu)$ be the process constructed above with measure $\nu$.
+Then, for every $\gamma\in\mathbb{R}$ and $K>0$, there exists $T_{0}>0$ such that for all $t\leq T_{0}$ we have
+$\sup_{\nu:\nu((0,\infty])\leq K}\sup_{m}\mathbb{E}_{Y_{0}^{m}}\Big[\exp\Big(-\gamma\int_{0}^{t}\langle Y_{s}(\nu),1\rangle ds\Big)\Big]<C,$ (6.22)
+Hereby, $T_{0}>0$ and $C>0$ only depend on $\langle\mu,1\rangle$ and $K$.
 Proof.
 If $\gamma\geq 0$, the statement follows immediately. Thus, we only need to consider the case $\gamma<0$. By Lemma 6.2 and (6.16) and using that
-$\int_{0}^{t}\langle Y^{\alpha}_{s},1\rangle ds\leq\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds,\quad\forall\alpha\in I_{t},t\geq 0,$
+$\int_{0}^{t}\langle Y^{\alpha}_{s}(\nu),1\rangle ds\leq\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds,\quad\forall\alpha\in I_{t},t\geq 0,$
 we get that there exists $T_{0}>0$ such that
-$\mathbb{E}\Big{[}\exp\Big{(}-\gamma\int_{0}^{t}\langle Y_{s},1\rangle ds\Big{)}\Big{]}\leq\mathbb{E}\Big{[}\exp\Big{(}-\gamma|I_{t}|\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds-\gamma\int_{0}^{t}\langle V_{s}(\mu),1\rangle ds\Big{)}\Big{]}<C,$
-for all $t\leq T_{0}$. This finishes the proof. ∎
+$\mathbb{E}_{Y_{0}^{m}}\Big[\exp\Big(-\gamma\int_{0}^{t}\langle Y_{s}(\nu),1\rangle ds\Big)\Big]\leq\mathbb{E}\Big[\exp\Big(-\gamma|I_{t}^{\nu}|\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds-\gamma\int_{0}^{t}\langle V_{s}(\mu),1\rangle ds\Big)\Big]<C,$
+for all $t\leq T_{0}$, $m\geq 1$ and $\nu$ with $\nu((0,\infty])\leq K$. This finishes the proof. ∎
 6.2 Convergence of the Dual Process
-Denote by $(Y_{t}^{n,m}(Y_{0}^{m}))_{t\geq 0}$ the process as constructed in (6.5) with initial condition $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$ and $n\in\mathbb{N}\cup\{\infty\}$, $m\in\mathbb{R}_{+}\cup\{\infty\}$.
-We assume that $Y_{0}^{n,m}=Y_{0}^{m}$, $\langle Y_{0}^{m},1\rangle>0$ and that, for $n\in\mathbb{N}$, $\nu$ is replaced by
-$\nu_{n}=\nu^{1}_{|[0,n]}+\nu^{2}_{|[0,n]}+d_{1}\delta_{n}+d_{2}\delta_{n}=\nu^{1}_{n}+\nu^{2}_{n}+d_{1}\delta_{n}+d_{2}\delta_{n}.$
-Here, $d_{1}$, $d_{2}$ and $a$ are defined as in (A).
+Let $\nu$ be as in (6.1).
+For $n\in\mathbbm{N}\cup\{\infty\}$ and $m\in\mathbb{R}_{+}\cup\{\infty\}$ denote by $(Y_{t}^{n,m}(Y_{0}^{m}))_{t\geq 0}$ the process as constructed in (6.6) with initial condition $Y_{0}^{m}=\mu+m\delta_{x_{0}}$, where $\mu\in\mathcal{M}_{F}^{+}$ and the measure $\nu$ is replaced by
+$\nu_{n}=\nu^{1}_{|[0,n]}+\nu^{2}_{|[0,n]}+d_{1}\delta_{n}+d_{2}\delta_{n}=\nu^{1}_{n}+\nu^{2}_{n}+d_{1}\delta_{n}+d_{2}\delta_{n},$
+for $n\in\mathbb{N}$.
+Here, $d_{1}$, $d_{2}$ and $a$ are defined as in (A) and for $n=\infty$ we set $\nu_{\infty}\equiv\nu$.
+We also assume that $\langle Y_{0}^{m},1\rangle>0$.
 For $i\geq 1$, we denote by $\bar{T}_{n,i}^{m}$ the jump times, by $T_{n,i}=\bar{T}_{n,i}^{m}-\bar{T}_{n,i-1}^{m}$ the time in between jumps, by $Z_{n,i}^{m}$ the jump sizes and by $U_{n,i}^{m}$ the jump positions defined as in (6.3)-(6.5).
@@
 Finally, we shall denote by $(S_{n,k}^{m})_{k\geq 1}$ the sequence independent of $\exp(1)$-distributed random variables from the construction of the process $Y^{n,m}$ in the beginning of Section 6.
+Let $X_{0}$ be a bounded, continuous and non-neagtive function on $\mathbb{R}$.
 In the remainder of the section we prove that
@@
 $\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big[(-1)^{J_{t}^{\infty,m}}\exp\Big(-\langle Y_{t}^{\infty,m},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,m},1\rangle ds\Big)\Big]\\ &\quad=\mathbb{E}\Big[(-1)^{J_{t}^{\infty,\infty}}\exp\Big(-\langle Y_{t}^{\infty,\infty},X_{0}\rangle-a\int_{0}^{t}\langle Y_{s}^{\infty,\infty},1\rangle ds\Big)\Big],\end{split}$
-for $t\in[0,T_{0}]$, where $T_{0}$ is chosen so that (6.8) holds with $\gamma=2a$ and where $X_{0}$ is a bounded and continuous function on $\mathbb{R}$.
-Note, that $T_{0}$ can be chosen independent of $n,m$ since
+for $t\in[0,T_{0}]$. Here $T_{0}$ is chosen as in Lemma 6.5 with $\gamma=2a$ and $K=\nu((0,\infty])$.
+Recall that by Lemma 6.5, $T_{0}$ can be chosen independent of $n,m$ since
 $\sup_{n}\nu_{n}((0,\infty))\leq\nu((0,\infty]).$
@@
 and thus
-$|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{x}},s)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|\leq|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s+\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|+|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s-\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|$
+$|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{x}},s)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|\leq|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s+\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|+|\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s-\eta)-\Theta^{k}(\overset{\leftarrow}{\boldsymbol{y}},s)|.$
 It thus remains to prove that $\eta(\delta)\rightarrow 0$ as $\delta\rightarrow 0$ and
@@
 as $\tilde{\eta}\rightarrow 0$. First, denote
-$D_{R}=\bigcup_{i=1}^{k+1}B_{R}(u_{i}).$
+$D_{R}=\bigcup_{i=1}^{k+1}B_{R}(u_{i}),\quad R>0.$
 We get for $\bar{t}\in(0,T)$ and $R>0$ that
-$\begin{split}&\eta(\delta)\leq\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,1\rangle dt\\ &\leq 2\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt\\ &\qquad+\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{\bar{t}}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,\mathbbm{1}_{D_{R}}\rangle dt\\ &\qquad+2k\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,\infty),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt.\end{split}$ (6.25)
+$\begin{split}&\eta(\delta)\leq\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,1\rangle dt\\ &\leq 2\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt\\ &\qquad+\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{\bar{t}}^{T}\langle|V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}})-V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{u}},\overset{\leftarrow}{\boldsymbol{z}},\boldsymbol{t})+z_{k+1}\delta_{u_{k+1}})|,\mathbbm{1}_{D_{R}}\rangle dt\\ &\qquad+2(k+1)\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,\infty),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt.\end{split}$ (6.30)
 The second term on the right hand side of (6.30) converges to zero as $\delta\rightarrow 0$ by Proposition 3.6 and for $\tilde{\epsilon}>0$ arbitrary we can choose $\bar{t}\in(0,T)$ small enough and $R>0$ large enough such that
-$\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt+k\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,0),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt\quad<\tilde{\epsilon}$.
+$\sup_{\boldsymbol{y}\in B_{\delta}(\boldsymbol{x})}\int_{0}^{\bar{t}}\langle V_{t}(\mathfrak{V}^{k}(\overset{\leftarrow}{\boldsymbol{v}},\overset{\leftarrow}{\boldsymbol{w}},\boldsymbol{\tau})+w_{k+1}\delta_{v_{k+1}}),1\rangle dt+(k+1)\int_{\bar{t}}^{T}\langle\mathcal{W}_{t}(0,0),\mathbbm{1}_{B_{R^{c}}(0)}\rangle dt\quad<\tilde{\epsilon}$.
 Thus, it remains to prove (6.29). To this end it is enough to prove
@@
   3. (iii)
-    $\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,m}}\exp\Big{(}-\langle Y^{\infty,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{\infty,m}_{s},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}\Big{[}(-1)^{J_{t}^{\infty,\infty}}\exp\Big{(}-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big{)}\Big{]}\end{split}$
+    $\begin{split}&\lim_{m\rightarrow\infty}\mathbb{E}\Big[(-1)^{J_{t}^{\infty,m}}\exp\Big(-\langle Y^{\infty,m}_{t}(x),X_{0}\rangle-a\int_{0}^{t}\langle Y^{\infty,m}_{s},1\rangle ds\Big)\Big]\\ &=\mathbb{E}\Big[(-1)^{J_{t}^{\infty,\infty}}\exp\Big(-\langle Y^{\infty,\infty}_{t}(x),X_{0})\rangle-a\int_{0}^{t}\langle Y^{\infty,\infty}_{s},1\rangle ds\Big)\Big],\end{split}$
     for all $t\in[0,T_{0}]$.
@@
 We only show (i), since (ii) and (iii) follow along similar lines.
-Let $T_{0}>0$ be such that (6.8) is satisfied with $\gamma=2a$.
+Let $T_{0}>0$ be such that (6.10) is satisfied with $\gamma=-2a$.
 Fix arbitrary $t>0$ and let $\epsilon>0$.
@@
 Lemma 6.10.
-Let $T_{0}>0$ be such that (6.8) holds for $\gamma=2a$ and let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with initial condition $X_{0}$ satisfying (1.5).
+Let $T_{0}>0$ be such that (6.22) in Lemma 6.5 holds for $\gamma=2a$ and $K=\nu((0,\infty])$.
+Furthermore, let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with initial condition $X_{0}$ satisfying (1.5).
 Furthermore, let $(g_{n})_{n\geq 1}$ be a sequence of measurable bounded functions on $\mathbb{R}$, that converge boundedly pointwise to a bounded, measurable function $g$ on $\mathbb{R}$.
@@
 Proof.
-Fix $m\in\mathbb{N}$. To lighten notation, we suppress the subscript $m$ for the rest of the proof. Note that by Lemma 6.2 and (6.13) we have
-$\sup_{n}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{4}\Big{]}\leq C\sup_{n}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},1\rangle ds\Big{)}^{4}\Big{]}\leq C\mathbb{E}[I_{T}^{4}]<\infty$
-Furthermore, by Lemma 6.8 it follows that
-$\begin{split}&(U_{n,1},U_{n,2},\ldots,T_{n,1},T_{n_{2}},\ldots,Z_{n,1},Z_{n,2},\ldots)\\ &\qquad\Rightarrow(U_{\infty,1},U_{\infty,2},\ldots,T_{\infty,1},T_{\infty,2},\ldots,Z_{\infty,1},Z_{\infty,2},\ldots)\end{split}$ (6.47)
-and thus by the Skorohod representation theorem there exists a probability space such that the convergence (6.47) holds almost surely. We now show that almost surely
-$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds=0.$
-First, we can use a similar argument to the proof of Lemma 6.8 to see that for almost all $\omega$ there exists $M(\omega)\in\mathbb{N}$ such that
-$\sup_{n}\bar{T}(\omega)_{n,M(\omega)+1}>T\quad,M(\omega)<\infty.$
-Furthermore, note that there exists $D(\omega)>0$ such that almost surely
+Fix $m\in\mathbb{N}$. To lighten notation, we suppress the superscript $m$ for the rest of the proof. Moreover, fix arbitrary $T\leq T_{0}$. Note that by Lemma 6.5 with $K=\nu((0,\infty])$ we have
+$\sup_{n}\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big)^{4}\Big]\leq c\sup_{n}\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},1\rangle ds\Big)^{4}\Big]<\infty.$ (6.52)
+Furthermore, we can use a similar argument to the proof of Lemma 6.9 to see that there exists $M\in\mathbb{N}$ such that
+$\mathbb{P}(\inf_{n}\bar{T}_{n,M}\leq T)<\epsilon.$
+and by Lemma 6.8 it follows that
+$\begin{split}&(U_{n,1},\ldots,U_{n,M},T_{n,1},\ldots,T_{n,M},Z_{n,1},\ldots,Z_{n,M})\\ &\qquad\Rightarrow(U_{\infty,1},\ldots,U_{\infty,M},T_{\infty,1},\ldots,T_{\infty,M},Z_{\infty,1},\ldots,Z_{\infty,M})\end{split}$ (6.53)
+and thus by the Skorohod representation theorem there exists a probability space such that the convergence (6.53) holds almost surely. We now show that
+$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds=0$
+for almost all $\omega$ in the event
+$A_{\epsilon}=\{\inf_{n}\bar{T}_{n,M}>T\}.$
+To this end, note that there exists $D(\omega)>0$ for $\omega\in A_{\epsilon}$ such that almost surely on $A_{\epsilon}$,
 $\sup_{n}|U_{n,k}(\omega)|<D(\omega),\quad D(\omega)<\infty,$
 for all $k\leq M$, since $(U_{n,k})_{n\geq 1}$ converges almost surely for each fixed $k$.
-Let $\epsilon>0$ be arbitrary.
+Let $\tilde{\epsilon}>0$ be arbitrary.
 Then, by using Lemma 3.2 and Proposition 3.6 we get the following.
-There exists $\bar{D}(\omega)>D(\omega)$ such that
-$\int_{0}^{T}\int_{B(0,\bar{D})^{c}}Y^{n}_{s-}(x)dxds\leq M\int_{0}^{T}\int_{B(0,\bar{D})^{c}}\big{(}\mathcal{W}(D,0)\big{)}(x)dxds\leq\epsilon,$
-with $\bar{D}(\omega)<\infty$ almost surely. Furthermore, there exists $\delta_{1}(\omega)>0$ such that
-$M\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds<\epsilon.$
+There exists $\bar{D}(\omega)>D(\omega)$ such that for $\omega\in A_{\epsilon}$ we have
+$\int_{0}^{T}\int_{B(0,\bar{D})^{c}}Y^{n}_{s-}(x)dxds\leq M\int_{0}^{T}\int_{B(0,\bar{D})^{c}}\big(\mathcal{W}(D,0)\big)(x)dxds\leq\tilde{\epsilon}$,
+with $\bar{D}(\omega)<\infty$ a.e. on $A_{\epsilon}$. We continue to work with $\omega\in A_{\epsilon}$. There exists $\delta_{1}>0$ such that
+$M\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds<\tilde{\epsilon}$.
 Finally, once again by Proposition 3.7, there exists $K(\omega)>0$ such that
 $M\sup_{s\in(\delta_{1},T),x\in B(0,\bar{D}+D)}\mathcal{W}_{s}(0,0)(x)\leq K<\infty,$
-almost surely. Thus, we get
-$\begin{split}&\int_{0}^{T}\langle Y_{s-}^{n},|h_{n}(X_{T-s})-h(X_{T-s})|\rangle ds\\ &\leq C\epsilon+\sum_{k=1}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\epsilon+\int_{0}^{T}\int_{B(0,\bar{D})}\big{(}V_{s}(\phi+m\delta_{x_{0}})\big{)}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\quad+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds,\end{split}$
+almost surely on $A_{\epsilon}$. Thus, we get
+$\begin{split}&\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\\ &\leq C\tilde{\epsilon}+\sum_{k=1}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\tilde{\epsilon}+\int_{0}^{T}\int_{B(0,\bar{D})}\big(V_{s}(\phi+m\delta_{x_{0}})\big)(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\quad+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds,\end{split}$
 where $C=\sup_{x\in\mathbb{R}}|g(x)|+\sup_{n\geq 1}\sup_{x\in\mathbb{R}}|g_{n}(x)|<\infty$ is independent of $n$. Furthermore, for $\delta_{1}>0$ sufficiently small we get
-$\begin{split}&\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq CM\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n+\delta_{1}}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|h_{n}(X_{T-s}(x))-h(X_{T-s}(x))|dxds\\ &\leq C\epsilon+M\big{(}\sup_{s\in(\delta_{1},T),x\in B(0,\bar{D}+D)}\mathcal{W}_{s}(0,0)(x)\big{)}\int_{0}^{T}\int_{B(0,\bar{D})}|h_{n}(X_{T-s}(x))-h(X_{T-s}(x))|dxds\\ &\leq C\epsilon+K\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\end{split}$
+$\begin{split}&\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq CM\int_{0}^{\delta_{1}}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\sum_{k=2}^{M}\int_{\bar{T}_{k-1,n+\delta_{1}}}^{\bar{T}_{k,n}\wedge T}\int_{B(0,\bar{D})}Y_{s-}^{n}(x)|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\tilde{\epsilon}+M\big(\sup_{s\in(\delta_{1},T),x\in B(0,\bar{D}+D)}\mathcal{W}_{s}(0,0)(x)\big)\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\\ &\leq C\tilde{\epsilon}+K\int_{0}^{T}\int_{B(0,\bar{D})}|g_{n}(X_{T-s}(x))-g(X_{T-s}(x))|dxds\end{split}$
 By the dominated convergence theorem, we get that
@@
 as $n\rightarrow\infty$. This demonstrates that
-$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\leq 2C\epsilon$
-almost surely. Since $\epsilon>0$ was arbitrary, this proves that
+$\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\leq 2C\tilde{\epsilon}$
+almost surely on $A_{\epsilon}$. Since $\tilde{\epsilon}>0$ was arbitrary, this proves that
 $\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds=0,$
-almost surely. Thus, by uniform integrability, we get
-$\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}\Big{[}\Big{(}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{2}]\\ &=\mathbb{E}\Big{[}\Big{(}\lim_{n\rightarrow\infty}\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big{)}^{2}\Big{]}=0.\end{split}$
+for almost all $\omega\in A_{\epsilon}$. Thus, by uniform integrability (see (6.52)), we get
+$\begin{split}\lim_{n\rightarrow\infty}\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big)^{2}\mathbbm{1}_{A_{\epsilon}}]=0.\end{split}$
+Finally, by Hölder’s inequality and (6.52), we get
+$\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s-}^{n},|g_{n}(X_{T-s})-g(X_{T-s})|\rangle ds\Big)^{2}\mathbbm{1}_{(A_{\epsilon})^{c}}]\leq c\epsilon\sqrt{\mathbb{E}\Big[\Big(\int_{0}^{T}\langle Y_{s}^{n},1\rangle ds\Big)^{4}\Big]}\leq C\epsilon,$
+and since $\epsilon>0$ was arbitrary, this finishes the proof. ∎
 6.3 Proof of Approximate Duality
 In this section, we assume that there exists a weak solution $(X_{t})_{t\geq 0}$ to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) is satisfied and prove an (approximate) duality relation.
-When $h$ is continuous, the existence of a solution was already resolved in [25, Theorem 1.1].
+When $h$ is continuous, the existence of a solution was already resolved in [35, Theorem 1.1].
 The existence of solutions when $h$ is discontinuous is proven in Section 7 and relies on the duality relation for the continuous case.
@@
 Proof.
-Note that
-$\begin{split}&\mathbb{E}[X_{t}(x)]=\mathbb{E}[\big{(}S_{t}X_{0}\big{)}(x)]+\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}\mathbb{E}[h(X_{s}(y))]dyds\\ &\leq\mathbb{E}[\big{(}S_{t}X_{0}\big{)}(x)]+C\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}(x-y)dyds,\end{split}$
+Note that, since $h$ is bounded,
+$\begin{split}&\mathbb{E}[X_{t}(x)]=\mathbb{E}[\big(S_{t}X_{0}\big)(x)]+\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}\mathbb{E}[h(X_{s}(y))]dyds\\ &\leq\mathbb{E}[\big(S_{t}X_{0}\big)(x)]+C\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}(x-y)dyds\\ &=\mathbb{E}[\big(S_{t}X_{0}\big)(x)]+Ct,\end{split}$
 for all $x\in\mathbb{R}$ and $t\geq 0$. This implies the result by standard estimates. ∎
@@
 Proof.
-The proof follows along similar lines as the proof of [21, Lemma 2.6], but we include it for the sake of completeness.
+The proof follows along similar lines as the proof of [30, Lemma 2.6], but we include it for the sake of completeness.
 Let $(\mu_{k})_{k\geq 1}$ be a sequence in $\mathcal{C}_{c}^{\infty}$ that converges weakly to $\mu$.
@@
 since $V_{t}(\mu_{k})$ solves (3.2) and
-$\mathbb{E}_{X}\Big{[}\int_{t}^{T}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-\cdots}(\mu_{k}))\Big{]}=0.$ (6.52)
+$\mathbb{E}_{X}\Big[\int_{t}^{T}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-\cdot}(\mu_{k}))\Big]=0.$ (6.58)
 Here, $M$ is an orthogonal martingale measure such that for all $\psi\in\mathcal{C}(\mathbb{R}_{+}\times\mathbb{R})\cap L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$, $M_{t}(\psi)$ is a square integrable martingale with quadratic variation given by
@@
 It remains to show the convergence as $k\rightarrow\infty$.
-Using [22, Lemma 2.1], the dominated convergence theorem and setting $\tilde{t}=T-t$ in (6.51) yields
+Using [31, Lemma 2.1], the dominated convergence theorem and setting $\tilde{t}=T-t$ in (6.57) yields
 $\begin{split}&\mathbb{E}_{X}\Big[\exp(-\langle V_{t}(\mu)),X_{T-t}\rangle+v_{t})\Big]=\mathbb{E}_{X}\Big[\exp(-\langle V_{0}(\mu),X_{T}\rangle)\Big]\\ &\qquad+\mathbb{E}_{X}\Big[\int_{0}^{t}\exp(-\langle V_{s}(\mu),X_{T-s}\rangle+v_{s})\big(\langle V_{s}(\mu),h(X_{T-s})\rangle+v_{s}^{\prime}\big)ds\Big],\quad\forall t\in[0,T].\end{split}$
@@
 Lemma 6.14.
-Let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) holds and let $a$ be defined as in (A).
+Let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$, $p>2$, such that (1.5) holds and let $a$ be defined as in (A).
 Let $Y^{n}$ be as above, independent of $X$.
@@
 Proof.
-First, choose $T_{0}>0$ small enough such that Lemma 6.5 is satisfied with $\gamma=2a$ and let $T\leq T_{0}$.
+First, choose $T_{0}>0$ small enough such that Lemma 6.5 is satisfied with $\gamma=2a$ and $K=\nu((0,\infty])$ and let $T\leq T_{0}$.
 Furthermore, set $v_{t}=-a\int_{0}^{t}\langle Y^{n}_{s},1\rangle ds$ for $t\geq 0$.
-By Lemma 6.12 we have
+Let $k\in\mathbb{N}$ be such that $\bar{T}_{n,k+1}\leq T$.
+Then, by Lemma 6.12 we have
 $\begin{split}&\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\Big(-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle-a\int_{0}^{\bar{T}_{n,k+1-}}\langle Y_{s}^{n},1\rangle ds\Big)\Big]\\ &=\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k}}}\exp\Big(-\langle Y_{\bar{T}_{n,k}},X_{T-\bar{T}_{n,k}}\rangle-a\int_{0}^{\bar{T}_{n,k}}\langle Y_{s}^{n},1\rangle ds\Big)\Big]\\ &\quad+\mathbb{E}_{X}\Big[\int_{\bar{T}_{n,k}}^{\bar{T}_{n,k+1}}(-1)^{J^{n}_{s-}}\exp\Big(-\langle Y_{s-}^{n},X_{T-s}\rangle-a\int_{0}^{s}\langle Y_{r}^{n},1\rangle dr\Big)\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big],\end{split}$ (6.61)
-since $t\mapsto X_{t}$ is continuous and $J_{\bar{T}_{n,k+1}-}=J_{s}$ for all $s\in(\bar{T}_{n,k},\bar{T}_{n,k+1})$. Let $\boldsymbol{\Delta}Y_{s}=Y_{s}-Y_{s-}$ for all $s\geq 0$, then we get
-$\begin{split}&\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle-a\int_{0}^{\bar{T}_{n,k+1-}}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle+v_{\bar{T}_{n,k+1}}\Big{)}\Big{]}\\ &+\exp\big{(}v_{\bar{T}_{n,k+1}}\big{)}\mathbb{E}_{X}\Big{[}(-1)^{J_{\bar{T}_{n,k+1}}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\\ &\hskip 199.16928pt-(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\Big{]}\\ &=\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k}}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k}},X_{T-\bar{T}_{n,k}}\rangle+v_{\bar{T}_{n,k}}\Big{)}\Big{]}\\ &+\mathbb{E}_{X}\Big{[}\int_{\bar{T}_{n,k}}^{\bar{T}_{n,k+1}}(-1)^{J^{n}_{s-}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle+v_{s}\Big{)}\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big{]}\\ &-\exp\big{(}v_{\bar{T}_{n,k+1}}\big{)}\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\\ &\times\Big{(}(1-\exp(-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}))\mathbbm{1}_{M_{n,k}=1}+\big{(}1+\exp\big{(}-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\big{)}\mathbbm{1}_{M_{n,k}=2}\Big{)}\Big{]},\end{split}$
+since $t\mapsto X_{t}$ is continuous and $J_{\bar{T}_{n,k+1}-}=J_{s}$ for all $s\in[\bar{T}_{n,k},\bar{T}_{n,k+1})$. Let $\boldsymbol{\Delta}Y_{s}=Y_{s}-Y_{s-}$ for all $s\geq 0$, then we get
+$\begin{split}&\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k+1}}}\exp\Big(-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle-a\int_{0}^{\bar{T}_{n,k+1}}\langle Y_{s}^{n},1\rangle ds\Big)\Big]\\ &=\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\Big(-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle+v_{\bar{T}_{n,k+1}}\Big)\Big]\\ &+\exp\big(v_{\bar{T}_{n,k+1}}\big)\mathbb{E}_{X}\Big[(-1)^{J_{\bar{T}_{n,k+1}}}\exp\big(-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big)\\ &\hskip 199.16928pt-(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big(-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big)\Big]\\ &=\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k}}}\exp\Big(-\langle Y_{\bar{T}_{n,k}},X_{T-\bar{T}_{n,k}}\rangle+v_{\bar{T}_{n,k}}\Big)\Big]\\ &+\mathbb{E}_{X}\Big[\int_{\bar{T}_{n,k}}^{\bar{T}_{n,k+1}}(-1)^{J^{n}_{s-}}\exp\Big(-\langle Y_{s-}^{n},X_{T-s}\rangle+v_{s}\Big)\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big]\\ &-\exp\big(v_{\bar{T}_{n,k+1}}\big)\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big(-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big)\\ &\times\Big((1-\exp(-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}))\mathbbm{1}_{M_{n,k}=1}+\big(1+\exp\big(-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}\rangle\big)\big)\mathbbm{1}_{M_{n,k}=2}\Big)\Big],\end{split}$
 where in the second equality we used (6.61) and the definition of $J^{n}$. This leads to
@@
 Proof.
-By Lemma 6.14 we have
+By Lemma 6.14 there exists $T_{0}>0$ such that for all $T\leq T_{0}$ we have
 $\begin{split}&\lim_{n\rightarrow\infty}\mathbb{E}\Big[(-1)^{J^{n}_{T}}\exp\Big(-\langle Y_{T}^{n},X_{0}\rangle-a\int_{0}^{T}\langle Y_{s}^{n},1\rangle ds\Big)\Big]\\ &=\mathbb{E}[\exp\big(-\langle Y_{0},X_{T}\rangle\big)]\\ &+\lim_{n\rightarrow\infty}\mathbb{E}\Big[\int_{0}^{t}(-1)^{J^{n}_{s-}}\exp\Big(-\langle Y_{s-}^{n},X_{T-s}\rangle-a\int_{0}^{s}\langle Y_{r-}^{n},1\rangle dr\Big)\Big(\langle Y_{s-}^{n},h(X_{T-s})\rangle-\langle Y_{s-}^{n},h_{n}(X_{T-s})\rangle\Big)ds\Big].\end{split}$
@@
 In this section we prove the weak existence part of Theorem 1.1. Again, our arguments were inspired by [4].
-Note, that we only need to proof weak existence in the case where $h$ is discontinuous due to [25, Theorem 1.1].
+Note, that we only need to proof weak existence in the case where $h$ is discontinuous due to [35, Theorem 1.1].
 That means we need to consider the case where $b_{0}\neq b_{1}\neq 0$.
@@
 The sequence of functions $(h_{n})_{n\geq 1}$ converges boundedly pointwise to $h$.
-As mentioned above, we get the existence of a weak solution $X^{n}$ by applying [25, Theorem 1.1] for each $n\in N$.
+As mentioned above, we get the existence of a weak solution $X^{n}$ by applying [35, Theorem 1.1] for each $n\in N$.
 When $h$ is replaced by $h_{n}$, we also get the weak uniqueness of solutions by applying the results from Section 8.
-Furthermore, we can apply the duality relation from Corollary 6.15 for $X^{n}$, which is conditional on existence.
+Furthermore, we can apply the duality relation from Corollary 6.15 for $X^{n}$, which is conditional on its existence.
 In a first step, we prove tightness of the sequence $(X^{n})_{n\geq 1}$, which can be achieved in a similar manner to the proves from [35].
@@
 Proof.
-The proof follows from standard estimates as in the proof of [25, Theorem 2.6]. ∎
-Lemma 7.1 demonstrates by [25, Lemma 6.3] that the sequence $(X^{n})_{n\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables is tight.
+The proof follows from standard estimates as in the proof of [35, Theorem 2.6]. ∎
+Lemma 7.1 demonstrates by [35, Lemma 6.3] that the sequence $(X^{n})_{n\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables is tight.
 Thus, by Prohorov’s theorem and Skorohod’s representation theorem, there exists a sub-sequence $(X^{n_{k}})_{k\geq 1}$ and another sequence $(\tilde{X}^{n_{k}})_{k\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables such that the following holds.
@@
 Proof.
-The proof is similar to the proof of [3, Lemma 5.3].
+The proof is similar to the proof of [4, Lemma 5.3].
 First, we prove the $L^{1}$ convergence.
 Let $T_{0}>0$ be such that Proposition 4.3 holds for $X^{n}$, for all $n\geq 0$.
-Note that it is possible to choose $T_{0}>0$ uniformly in $n$, since it only depends on $\sup_{n}\nu_{n}^{1}(0,\infty)+\nu_{n}^{2}(0,\infty)$.
-For simplicity denote $X_{t}(x)$ by $Z$ and $X^{n}_{t}(x)$ by $Z_{n}$.
-Below we use the fact that almost surely $Z_{n},Z\geq 0$.
-Thus we get
+Note that it is possible to choose $T_{0}>0$ uniformly in $n$, since it only depends on $\sup_{n}\nu_{n}^{1}(0,\infty)+\nu_{n}^{2}(0,\infty)$, where $\nu_{n}^{1},\nu_{n}^{2}$ are given by
+$\nu_{n}^{i}=\nu^{i}_{|[0,n]}+d_{i}\delta_{n},\quad i\in\{1,2\},n\geq 1.$
+For simplicity denote $X_{t}(x)$ by $Z$ and $X^{n}_{t}(x)$ by $Z_{n}$. Below we use the fact that almost surely $Z_{n},Z\geq 0$. Thus we get
 $\begin{split}&\mathbb{E}[|\exp(-nZ_{n})-\mathbbm{1}_{Z=0}|]\\ &=\mathbb{E}[|\exp(-nZ_{n})-\mathbbm{1}_{Z=0}|\mathbbm{1}_{Z=0}]+\mathbb{E}[|\exp(-nZ_{n})-\mathbbm{1}_{Z=0}|\mathbbm{1}_{Z>0}]\\ &=\mathbb{E}[|\exp(-nZ_{n})-1|\mathbbm{1}_{Z=0}]+\mathbb{E}[|\exp(-nZ_{n})|\mathbbm{1}_{Z>0}]\\ &=\mathbb{E}[\mathbbm{1}_{Z=0}]-\mathbb{E}[\exp(-nZ_{n})]+2\mathbb{E}[|\exp(-nZ_{n})|\mathbbm{1}_{Z>0}].\end{split}$ (7.3)
@@
 $\mathbb{E}[\exp(-\lambda\langle f,X_{t}^{1}\rangle)]=\mathbb{E}[\exp(-\lambda\langle f,X_{t}^{2}\rangle)]\quad\forall\lambda\geq 0.$
-Since $f$ was arbitrary it follows by [13, Corollary 2.3] that $X^{1}_{t}\overset{d}{=}X_{t}^{2}$. ∎
+Since $f$ was arbitrary it follows by [19, Corollary 2.3] that $X^{1}_{t}\overset{d}{=}X_{t}^{2}$. ∎
 Theorem 8.2.
@@
 Thus, it remains to prove uniqueness on $0,T_{0}]$.
-However, by using Lemma 8.1, this follows along the same lines as the proof of [11, Theorem 4.2].
+However, by using Lemma 8.1, this follows along the same lines as the proof of [15, Theorem 4.2].
 It is only necessary to show that t $\mathbb{P}\circ(X_{t})^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and that $X_{t}$ satisfies (1.5) for all $t\in[0,T_{0}]$.
@@
 Proof of Theorem 1.2.
-Let $T_{0}>0$ be small enough such that
-$\mathbb{E}[\operatorname{e}^{|I_{T_{0}}(Y_{0})||}]<\infty$ (9.1)
-with $Y_{0}=\infty\delta_{0}$. Such $T_{0}>0$ exists due to Lemma 6.2. For $s=0$ the assertion follows by assumption. Thus, let $s\in(0,T_{0}]$. Using Proposition 4.3 and Remark 4.4 it follows that
+We only prove the statement for the case $b_{1}\geq 0$.
+When $b_{1}<0$ the statement follows by a comparison with the solution when $b_{1}=0$, by using weak uniqueness, [35, Corollary 2.4] and standard approximation and tightness arguments.
+Thus, we assume without loss of generality that $b_{1}\geq 0$ from now on.
+The proof relies on the duality relation.
+To this end, let $Y$ be the process constructed in the beginning of Section 6 with $Y_{0}=\infty\delta_{0}$ and let $I_{t}(Y_{0})$ for $t\geq 0$ be defined as in (6.8).
+By Lemma 6.2, there exists $T_{0}>0$ such that
+$\mathbb{E}[\operatorname{e}^{|I_{T_{0}}(Y_{0})||}]<\infty.$ (9.1)
+For $t=0$ the assertion (1.7) follows by the assumptions of the theorem. Thus, let $s\in(0,T_{0}]$. Using Proposition 4.3 and Remark 4.4 it follows that
 $\begin{split}&\mathbb{E}\Big[\int_{\mathbb{R}}\mathbbm{1}_{X_{s}(x)>0}dx\Big]=\lim_{n\rightarrow\infty}\mathbb{E}\Big[\int_{\mathbb{R}}\big(1-\operatorname{e}^{-nX_{s}(x)}\big)dx\Big]\\ &=\lim_{n\rightarrow\infty}\int_{\mathbb{R}}\big(1-\mathbb{E}[\operatorname{e}^{-nX_{s}(x)}]\big)dx\\ &=\lim_{n\rightarrow\infty}\int_{\mathbb{R}}\big(1-\mathbb{E}[\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})\rangle}]\big)dx.\end{split}$ (9.2)
@@
 where $\bar{T}_{n,k}^{x}$ is time of the $k$-th jump of the process $Y(n\delta_{x})$.
-That is, $K^{n}_{t}$ is the number of jumps of $Y(n\delta)$ on the time interval $(0,t]$.
+That is, $K^{n}_{t}$ is the number of jumps of $Y(n\delta_{x})$ on the time interval $(0,t]$.
 We furthermore denote $T_{n,k}^{x}=\bar{T}_{n,k}^{x}-\bar{T}_{n,k-1}^{x}$, for $k\geq 1$.
@@
 $I_{3}^{n}(s,x)\leq C(|x|^{1/4}+1)\min(s,R_{x}^{2})^{-1/2}\operatorname{e}^{-\frac{(R_{x})^{2}}{2\min(R_{x}^{2},s)}}.$ (9.7)
-Finally, we estimate $I_{2}^{n}(s,x)$. To this end, we brake up the event into several parts that we can handle more easily.
+Finally, we estimate $I_{2}^{n}(s,x)$. To this end, we break up the event into several parts that we can handle more easily.
 $\begin{split}&I_{2}^{n}(s,x)\leq\mathbb{P}(\tau_{R_{x},n}\leq K^{n}_{s},K^{n}_{s}\leq|x|^{1/4})\\ &\leq\mathbb{P}(U_{k,x}^{n}\leq R_{x}+R\text{ for some }k\in\{1,\ldots,K^{n}_{s}\},K^{n}_{s}\leq|x|^{1/4})\\ &\leq\sum_{m=1}^{\lceil|x|^{1/4}\rceil}\sum_{k=1}^{m}\mathbb{P}(U_{n,j}^{x}>R_{x}+R,j\in\{1,\ldots k-1\},U_{n,k}^{x}\leq R_{x}+R,K_{s}^{n}=m).\end{split}$ (9.8)
 Now, we denote $\eta_{x}=R_{x}/(|x|^{1/4}+1)$. Then, for $1\leq k\leq m$ we get
-$\begin{split}&\mathbb{P}(U_{n,j}^{x}>R_{x}+R,j\in\{1,\ldots k-1\},U_{n,k}^{x}\leq R_{x}+R,K_{s}^{n}=m)\\ &\leq\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\\ &\qquad+\mathbb{P}(U_{n,j}^{x}\leq x-j\eta_{x}\text{ for some }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\end{split}$ (9.9)
+$\begin{split}&\mathbb{P}(U_{n,j}^{x}>R_{x}+R,j\in\{1,\ldots k-1\},U_{n,k}^{x}\leq R_{x}+R,K_{s}^{n}=m)\\ &\leq\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\\ &\qquad+\mathbb{P}(U_{n,j}^{x}\leq x-j\eta_{x}\text{ for some }j\in\{1,\ldots k-1\},K_{s}^{n}=m).\end{split}$ (9.9)
 We first estimate the first term on the right hand side of (9.9). For $1\leq k\leq m\leq|x|^{1/4}+1$ we get
@@
 where $\mathcal{G}_{n,x,k}$ is the $\sigma$-algebra generated by $(U_{n,1}^{x},\ldots U_{n,k-1}^{x},T_{n,1}^{x},\ldots T_{n,k}^{x})$ .
-For $y\in\mathbb{R}$ we get using Lemma 3.3, Proposition 3.6 and the construction of the process $Y(n\delta_{x})$ that
+For $y\in\mathbb{R}$, using Lemma 3.3, Proposition 3.6 and the construction of the process $Y(n\delta_{x})$ we get that
 $\begin{split}&\mathbbm{1}_{\bar{T}_{n,k}^{x}\leq s}\mathbb{E}[\mathbbm{1}_{U_{n,k}^{x}\leq y}|\mathcal{G}_{n,x,k}]\\ &=\mathbbm{1}_{\bar{T}_{n,k}^{x}\leq s}\frac{\int_{-\infty}^{y}Y_{\bar{T}_{n,k}^{x}-}(n\delta_{x})(y^{\prime})dy^{\prime}}{\int_{\mathbb{R}}Y_{\bar{T}_{n,k}^{x}-}(n\delta_{x})(y^{\prime})dy^{\prime}}\leq C\sum_{j=0}^{k-1}\int_{-\infty}^{y}\mathcal{W}_{\bar{T}_{n,k}^{x}-\bar{T}_{n,j}^{x}}(U_{n,j}^{x},0)(y^{\prime})dy^{\prime}),\end{split}$ (9.10)
@@
 $\begin{split}&\lim_{n\rightarrow\infty}\int_{\mathbb{R}}\big(1-\mathbb{E}[\operatorname{e}^{-\langle X_{0},Y_{s}(n\delta_{x})\rangle}]\big)dx\\ &\leq C\Big(R^{\ast}+s^{-1/2}\int_{R^{\ast}}^{\infty}x\operatorname{e}^{-\frac{(x-R)^{2}}{8s|x|^{1/2}}}dx\Big),<\infty.\end{split}$
-where we used that $\eta_{x}\leq R_{x}$ for $x\geq R^{\ast}$. This proves the assertion. ∎
+where we used that $\eta_{x}\leq R_{x}$ for $x\geq R^{\ast}$. This proves (1.7). ∎
+10 Proof of Theorem 1.4
+The proof of Theorem 1.4 relies on a rescaling of (1.1) as well as a comparison with solutions to
+$\begin{split}d_{t}\bar{X}^{\theta}_{t}(x)&=\frac{1}{6}\Delta\bar{X}^{\theta}_{t}(x)+\theta\bar{X}^{\theta}_{t}(x)-\big(\bar{X}^{\theta}_{t}(x)\big)^{2}+\sqrt{\bar{X}^{\theta}_{t}(x)}\dot{W}(t,x),\quad t>0,\quad x\in\mathbb{R},\\ \bar{X}^{\theta}_{0}(x)&=f(x),\quad x\in\mathbb{R},\end{split}$ (10.1)
+where $\theta>0$ and $f\in\mathcal{C}_{c}^{+}$.
+Lemma 10.1.
+Let $a,b,c>0$ and let $X$ be a solution to (1.1). Then
+$\tilde{X}_{t}(x)=cX_{at}(bx),\quad t\geq 0,\quad x\in\mathbb{R},$
+is a solution to
+$\begin{split}d_{t}\tilde{X}_{t}(x)&=\frac{a}{2b^{2}}\Delta\tilde{X}_{t}(x)+ach_{c}(\tilde{X}_{t}(x))+\frac{c^{1/2}a^{1/2}}{b^{1/2}}\sqrt{\tilde{X}_{t}(x)}\dot{W}(t,x),\quad x\in\mathbb{R},t\geq 0,\\ \tilde{X}_{0}(x)&=cX_{0}(bx),\quad x\in\mathbb{R},\end{split}$
+where $h_{c}(x)=h(c^{-1}x)$ for $x\in\mathbb{R}$.
+Proof.
+This follows along the same lines as the proof of [28, Lemma 2.1.2]. ∎
+Lemma 10.2.
+Let
+$h(x)=\theta(1-e^{-\lambda x})$
+for some $\lambda,\theta>0$ and let $X$ be a solution to (1.1) with $X_{0}=f\in\mathcal{C}_{c}^{+}$ such that $f\neq 0$. Then it follows that
+$\mathbb{P}(\langle X_{t},1\rangle>0,\forall t>0)>0.$
+Proof.
+By [28, Theorem 1] there exists $\theta^{\ast}>0$ such that
+$\mathbb{P}(\langle\bar{X}^{\theta^{\ast}}_{t},1\rangle>0,\forall t>0)>0.$
+Now, we set $a=1/3b^{2}$, and $c=3b^{-1}$ for $b>0$. Then, by Lemma 10.1, $\tilde{X}_{t}(x)=cX_{at}(bx)$ solves
+$d_{t}\tilde{X}_{t}=\frac{1}{6}\Delta\tilde{X}_{t}(x)+b\theta(1-\operatorname{e}^{-\frac{b}{3}\lambda\tilde{X}_{t}(x)})+\sqrt{\tilde{X}_{t}(x)}\dot{W}(t,x),\quad x\in\mathbb{R},t\geq 0.$ (10.2)
+Suppose, that there exists $b>0$ such that
+$b\theta(1-\operatorname{e}^{-\frac{b}{3}\lambda x})\geq\theta^{\ast}x-x^{2},\quad\forall x\geq 0.$ (10.3)
+By [22, Theorem 3.3] and standard approximation and tightness arguments it follows by weak uniqueness that $\tilde{X}$ and $\bar{X}$ can be defined on a common probability space such that for all $t\geq 0$ we have
+$\tilde{X}_{t}(x)\geq\bar{X}^{\theta^{\ast}}_{t}(x),\quad x\in\mathbb{R},$
+almost surely. Again by weak uniqueness and by continuity of $\tilde{X}$ and $\bar{X}$ we get
+$\mathbb{P}(\tilde{X}_{t}(x)\geq\bar{X}^{\theta^{\ast}}_{t}(x),\forall t\geq 0,x\in\mathbb{R})=1.$
+This in particular implies that $\mathbb{P}\langle(\tilde{X}_{t},1\rangle>0,\forall t>0)>0$ and thus $\mathbb{P}(\langle X_{t},1\rangle>0,\forall t>0)>0$.
+Thus, it only remains to establish (10.3), which is equivalent to proving
+$\frac{b^{3}\lambda^{2}}{9}\theta(1-\operatorname{e}^{-x})\geq\theta^{\ast}\frac{b\lambda}{3}x-x^{2},\quad\forall x\geq 0,$
+for some $b>0$. For $x\in[0,1]$ we may use that $(1-\operatorname{e}^{-x})\geq\frac{1}{2}x$ and thus in this regime it is sufficient to choose $b>0$ sufficiently large so that
+$\frac{b^{3}\lambda^{2}}{9}\theta\frac{1}{2}\geq\theta^{\ast}\frac{b\lambda}{3}.$
+Furthermore, for note that
+$\Big(\theta^{\ast}\frac{b\lambda}{3}\Big)^{2}\frac{1}{4}\geq\theta^{\ast}\frac{b\lambda}{3}x-x^{2},$
+for $x\in\mathbb{R}$. Thus, for $x>1$, we need to choose $b>0$ sufficiently large such that also
+$\frac{b^{3}\lambda^{2}}{9}\theta(1-\operatorname{e}^{-1})\geq\Big(\theta^{\ast}\frac{b\lambda}{3}\Big)^{2}\frac{1}{4}.$
+That means, (10.3) is satisfied for
+$b>\max\Big(\sqrt{\frac{6\theta^{\ast}}{\lambda\theta}},\frac{(\theta^{\ast})^{2}}{4\theta(1-e^{-1})}\Big),$
+which finishes the proof. ∎
+Proof of Theorem 1.4.
+Suppose that there exists $\lambda,\theta>0$ such that
+$h(x)\geq\theta(1-\operatorname{e}^{-\lambda x}),\quad\forall x\geq 0.$ (10.4)
+Then, by weak uniqueness, [35, Corollary 2.4] and standard approximation and tightness arguments we have that
+$\mathbb{P}(\langle X_{t},1\rangle>0,\forall t>0)>\mathbb{P}(\langle\hat{X}_{t},1\rangle>0,\forall t>0)$
+where $\hat{X}$ is the solution to (1.1) with $h(x)=\theta(1-\operatorname{e}^{-\lambda x})$ for all $x\in\mathbb{R}$.
+Thus, by Lemma 10.2 it is sufficient to verify (10.4).
+If $b_{1}>0$, then (10.4) is clearly satisfied for all $\lambda>0$ and $\theta\leq b_{1}$.
+When $b_{1}=0$ we have the estimate
+$\int_{0}^{\infty}(1-\operatorname{e}^{-\tilde{\lambda}x})\nu^{1}(d\tilde{\lambda})\geq(1-\operatorname{e}^{-\lambda x})\nu^{1}([\lambda,\infty)),$
+where $\lambda>0$ is chosen in such a way that $\nu^{1}([\lambda,\infty))>0$. Then, (10.4) is clearly satisfied for $\theta\leq\nu^{1}([\lambda,\infty))$. ∎